### v1

# Signature matrices of membranes

Abstract
We prove that, unlike in the case of paths, the signature matrix of a membrane does not satisfy any algebraic relations. We derive novel closed-form expressions for the signatures of polynomial membranes and piecewise bilinear interpolations for arbitrary $2$-parameter data in $d$-dimensional space. We show that these two families of membranes admit the same set of signature matrices and scrutinize the corresponding affine variety.
Keywords: two-parameter signatures, affine algebraic varieties, matrix congruence, geometry of paths and membranes, iterated integrals
MSC codes: 60L10, 14Q15, 13P25
1 Introduction
The signature of a path.
A path is a continuous map $X:[0,1]\to\mathbb{R}^{d}$ with piecewise continuously differentiable coordinate functions. Its $k$-th level signature $\sigma^{(k)}(X)$, as introduced by Chen in the seminal work [8], is a $k$-tensor of format $d\times\dots\times d$, defined by forming certain iterated integrals. In particular, its second level signature can be viewed as a signature matrix. It is defined as follows:
$\sigma^{(2)}(X)_{i,j}$ $:=\int_{0}^{1}\int_{0}^{t_{2}}\partial X_{i}(t_{1})\,\partial X_{j}(t_{2})\;\mathrm{d}t_{1}\mathrm{d}t_{2}$ (1)
These matrices were studied in [2]. They satisfy some interesting properties: for example, their symmetric part has always rank $1$, due to the shuffle relations satisfied by path signatures.
Moreover, the set of signature matrices of polynomial paths with degree $\leq m$ and the set of signature matrices of piecewise linear paths with at most $m$ segments agree. The equations that vanish on these sets, that is, the varieties of signature matrices, are well-understood: they can be described as certain determinantal varieties [2, Theorem 3.4].
The signature of a membrane.
A membrane is a continuous map $X:[0,1]^{2}\rightarrow{\mathbb{R}}^{d}$ where the coordinate functions $X_{1},X_{2},\dots,X_{d}$ are (piecewise) continuously differentiable functions in the two arguments. In particular the differentials $\partial_{12}X_{j}(s,t):=\frac{\partial^{2}}{\partial s\partial t}X_{j}(s,t)$ obey the usual rules of calculus. Generalizing the path setting, one can associate a collection of tensors to $X$ via iterated integrals (see Definition 2.2). In particular, the signature matrix of a membrane $X$ is defined as an $\mathbb{R}$-valued $d\times d$-matrix with entries
$\sigma^{(2)}(X)_{i,j}$ $:=\int_{0}^{1}\int_{0}^{1}\int_{0}^{t_{2}}\int_{0}^{s_{2}}\partial_{12}X_{i}(s_{1},t_{1})\,\partial_{12}X_{j}(s_{2},t_{2})\;\mathrm{d}s_{1}\mathrm{d}t_{1}\mathrm{d}s_{2}\mathrm{d}t_{2}$ (2)
for every $1\leq i,j\leq d$. This matrix has recently been introduced as the second level of the two-parameter $\mathsf{id}$-signature [26, 48, 14].
While it is easy to see that, contrary to path signatures, this signature does not satisfy shuffle relations, little is known about the algebraic relations satisfied by signature tensors in general. In this paper, inspired by [2], we address this question for the case of matrices. We write $\sigma(X):=\sigma^{(2)}(X)$ in the following. We focus on polynomial membranes of order $(m,n)$, that is, membranes $X$ whose component functions are polynomials of multidegree $\leq(m,n)$, and piecewise bilinear membranes of order $(m,n)$, see Definition 3.14. Their signature matrices (2) form semialgebraic subsets of $\mathbb{R}^{d^{2}}$. To study the polynomials that vanish on these sets, we move to the algebraically closed field $\mathbb{C}$ and consider the Zariski closures
$\mathcal{P}_{d,m,n}^{\mathsf{pol}}\quad\text{ and }\quad\mathcal{P}_{d,m,n}^{\mathsf{bil}}.$ (3)
Main question.
What can be said about the varieties of signature matrices of polynomial membranes and piecewise bilinear membranes?
Figure 1: A polynomial membrane in $\mathbb{R}^{2}$ and a piecewise bilinear membrane in $\mathbb{R}^{3}$. The colors correspond to images of rectangles in a uniform partitioning of $[0,1]^{2}$. See Example 3.7 and Figure 3 for a more detailed description.
An implicitization problem.
Our strategy is to view the varieties (3) as solutions to an implicitization problem. To this end, we show that both classes of membranes admit a dictionary membrane (see [41]), i.e. a membrane which yields all other membranes in the class via linear transforms, up to certain terms that vanish under the signature map.
Our first candidate of a dictionary is the $mn$-dimensional moment membrane,
$\mathsf{Mom}^{m,n}:[0,1]^{2}$ $\to\mathbb{R}^{mn},(s,t)\mapsto(s,s^{2},\dots,s^{m})\otimes(t,t^{2},\dots,t^{n}),$ (4)
where $\otimes$ is the Kronecker product. If $m=n=2$, then $\mathsf{Mom}^{2,2}(s,t)=(st,st^{2},s^{2}t,s^{2}t^{2})$ has the signature matrix
$\sigma(\mathsf{Mom}^{2,2})=\begin{pmatrix}\frac{1}{4}&\frac{1}{3}&\frac{1}{3}&\frac{4}{9}\\[4.0pt] \frac{1}{6}&\frac{1}{4}&\frac{2}{9}&\frac{1}{3}\\[4.0pt] \frac{1}{6}&\frac{2}{9}&\frac{1}{4}&\frac{1}{3}\\[4.0pt] \frac{1}{9}&\frac{1}{6}&\frac{1}{6}&\frac{1}{4}\end{pmatrix}=\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}\otimes\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}$ (5)
with a Kronecker factorization into path signatures (Corollary 3.6). Note that indeed every $d$-dimensional polynomial membrane $X$ with terms of multidegree $\geq(1,1)$ is of the form $X=A\,\mathsf{Mom}^{m,n}$. With equivariance of the signature (Lemma 2.6) we obtain the closed formula,
$\sigma(X)=A\sigma(\mathsf{Mom}^{m,n})A^{\top},$ (6)
as a matrix congruence. Hence, continuing with our example (5) and $d=2$, we have
$X=A\,\mathsf{Mom}^{2,2}=\begin{pmatrix}a_{1,1}st+a_{1,2}st^{2}+a_{1,3}s^{2}t+a_{1,4}s^{2}t^{2}\\ a_{2,1}st+a_{2,2}st^{2}+a_{2,3}s^{2}t+a_{2,4}s^{2}t^{2}\end{pmatrix}$ (7)
Thus, we obtain the first entry of our signature matrix from (5) and (6): It is the homogeneous polynomial
$\sigma(X)_{1,1}=\frac{1}{4}a_{1,1}^{2}$ $+\frac{1}{2}a_{1,1}a_{1,2}+\frac{1}{2}a_{1,1}a_{1,3}+\frac{5}{9}a_{1,1}a_{1,4}+\frac{1}{4}a_{1,2}^{2}$
$+\frac{4}{9}a_{1,2}a_{1,3}+\frac{1}{2}a_{1,2}a_{1,4}+\frac{1}{4}a_{1,3}^{2}+\frac{1}{2}a_{1,3}a_{1,4}+\frac{1}{4}a_{1,4}^{2}$
of degree two. The remaining three polynomial entries of $\sigma(X)$ are provided in Example 3.7. We illustrate the polynomial membrane (7) for explicit $A$ in Figure 3.
The axis membrane.
Interpolation is a necessary step to fit discrete data into the framework of iterated integrals. Our second important dictionary, the axis membrane $\mathsf{Axis}^{m,n}$, is inspired by the dictionary for piecewise linear paths [2, 41] (see Definition 3.15). This $mn$-dimensional membrane allows us to interpolate arbitrary $d$-dimensional discrete data on an $m\times n$ grid via a piecewise bilinear membrane. For an illustration of such an interpolation see $Y$ from Figure 3. We provide novel closed-form expressions and computational complexity statements for its signatures (Theorem 3.19).
Main results.
Using our description via dictionaries, we obtain an analogue of [2, Theorem 3.3] for membranes:
Theorem (4.1).
For all $d,m$ and $n$,
$\mathcal{P}^{\mathsf{pol}}_{d,m,n}=\mathcal{P}^{\mathsf{bil}}_{d,m,n}=:\mathcal{P}_{d,m,n}.$
In fact, already the defining semi-algebraic sets agree.
These varieties are irreducible (Proposition 4.3) and we obtain a grid
[…]
where we can choose $M,N\in\mathbb{N}$ such that
[…]
$\mathcal{P}_{d,1,n}\subseteq\dots\subseteq\mathcal{P}_{d,M-1,n}\subseteq\mathcal{P}_{d,M,n}=\mathcal{P}_{d,M+1,n}$ =$\dots$= …
for all $m,n\in\mathbb{N}$. As $\mathcal{P}_{d,1,m}=\mathcal{P}_{d,m,1}$ agrees with the path varieties studied in [2] (Lemma 4.5), it follows that $\mathcal{P}_{d,1,N}=\mathcal{P}_{d,M,1}$ is the universal path variety. However, in strong contrast to the path case, in general there are no algebraic relations in our two-parameter signature matrices (2) at all:
Theorem (4.7).
The limit variety $\mathcal{P}_{d,M,N}$ in (8) attains the full ambient dimension, i.e. $\mathcal{P}_{d,M,N}=\mathbb{C}^{d^{2}}$.
The varieties become more interesting if $m$ and $n$ are small compared to $d$. Employing the theory of orbits under congruence [45, 30] we were able to find explicit formulas for the dimensions of $\mathcal{P}_{d,m,n}$ whenever $mn\leq d$, generalizing [2, Theorem 3.4]. Note that there are $dmn\leq d^{2}$ parameters involved, so the dimension can be at most $dmn$.
Theorem (4.9).
For $mn\leq d$ the dimension of $\mathcal{P}_{d,m,n}$ is
  • $dmn-\frac{1}{2}m^{2}n^{2}+m^{2}(n-1)+(m-1)n^{2}-\frac{7}{2}mn+4(m+n)-4$ for even $m,n$.
  • $dmn-\frac{1}{2}m^{2}n^{2}+m^{2}(n-1)+(m-1)n^{2}-\frac{3}{2}mn+m+n$ for even $m$, odd $n$.
  • $dmn-\frac{1}{2}m^{2}n^{2}+m^{2}(n-1)+(m-1)n^{2}-\frac{7}{2}mn+3(m+n)-2$ for odd $m,n$.
These results are complemented by some explicit computations using the open-source computer algebra systems Macaulay 2 [28, 19], and OSCAR [40, 13], see e.g. Examples 4.12 and 4.2, Remark 4.13 and Example 4.14.
Related work
The collection of iterated integrals of a path has been used as a tool for classification in various settings since Chen’s seminal 1954 work [8]. After pioneering work by Lyons and collaborators on the use of signatures as features for data streams in the early 2010s [29, 32], there has been a surge in the number of applications, which range from motion recognition [47], finance [4], deep learning [33, 34, 46, 35], oceanography [44], and lately data science for images [48, 31]. We refer the interested reader to the surveys [10, 38].
Since [2, 41, 24] established the initial connection between stochastic analysis and algebraic geometry, the signature has also been applied in various related fields, including the theory of tensor decompositions [3, 22, 25], clustering tasks [11, 39], tropical time series analysis [15, 16] and the geometry of paths [17, 1, 42].
Recently, the signature has been extended to two parameters, generalizing from paths to membranes: the mapping space signature [26] provides a principled two-parameter extension of Chen’s iterated integrals, based on its original use in topology, whereas the two-parameter sums signature [18] characterizes equivalence up to warping in two directions. These two approaches have recently been combined in [14] using mixed partial derivatives. Due to the lack of a two-parameter version of Chen’s identity in all the above references, recent works [9, 37, 36] use higher category theory to find an appropriate generalization of a group that is rich enough to represent the horizontal and vertical concatenation of membranes.
2 Iterated integrals
The iterated-integrals signature $\sigma=\sigma(X)$ of a $d$-dimensional path $X:[0,1]\to\mathbb{R}^{d}$ is an element of the tensor series algebra $T((\mathbb{R}^{d}))\cong\operatorname{Hom}(T(\mathbb{R}^{d}),\mathbb{R})$, defined by
$\sigma:T(\mathbb{R}^{d})$ $\to\mathbb{R},$
$e_{i_{1}}\otimes\dots\otimes e_{i_{k}}$ $\mapsto\int_{\Delta}\mathrm{d}X_{i_{1}}(t_{1})\dots\mathrm{d}X_{i_{k}}(t_{k}),$ (9)
where $\Delta$ is the simplex $10<t_{1}<\dots<t_{k}<$ 1. It is common to write the tensor $e_{i_{1}}\otimes\dots\otimes e_{i_{k}}$ as a word $i_{1}\dots i_{k}$, identifying the tensor algebra with the free associative algebra over the letters 1 to d. However, the tensor algebra viewpoint allows us to consider the signature for paths (and later membranes) in any $\mathbb{R}$-vector space $V$, without choosing a basis, which will be useful in the following. Let us first recall some examples in the (classical, one-parameter) path setting.
Example 2.1.
  1. 1.
    For fixed $u\in\mathbb{R}^{d}$, the $k$-th signature tensor of the linear path $[0,1]\rightarrow\mathbb{R}^{d},t\mapsto ut$ is given by
    $\sigma^{(k)}_{i_{1},\dots,i_{k}}=\frac{u_{i_{1}}\dots u_{i_{k}}}{k!}$ (10)
    where $1\leq i_{1},\dots,i_{k}\leq d$.
  2. 2.
    Let $\mathsf{Axis}^{d}:[0,1]\to\mathbb{R}^{d}$ be the canonical axis path from [2, Example 2.1] or [41, eq. (8)] of order $d$, i.e., the unique piecewise linear path with $d$ segments whose graph interpolates the $d+1$ equidistant support points $(\frac{i}{d},\sum_{1\leq j\leq i}e_{j})$ with $0\leq i\leq d$. This path is illustrated in Figure 2 for $d=3$. For all $d$, its signature matrix is the upper triangular matrix
    $\sigma^{(2)}_{i_{1},i_{2}}=\begin{cases}1&\text{if }i_{1}<i_{2}\\ \frac{1}{2}&\text{if }i_{1}=i_{2}\\ 0&\text{if }i_{1}>i_{2}.\end{cases}$ (11)
  3. 3.
    We denote by $\mathsf{Mom}^{d}:[0,1]\to\mathbb{R}^{d},t\mapsto(t,\dots,t^{d})$ the moment path of degree $d$. We have the closed formula
    $\sigma^{(k)}_{i_{1},i_{2},\dots,i_{k}}=\frac{i_{2}i_{3}\dots i_{k}}{(i_{1}+i_{2})(i_{1}+i_{2}+i_{3})\dots(i_{1}+\dots+i_{k})}$ (12)
    for all $i_{1},\dots,i_{k}$ due to [2, Example 2.2].
$\mathsf{graph}(\mathsf{Axis}^{3}_{1})$
$\mathsf{graph}(\mathsf{Axis}^{3}_{2})$
$\mathsf{graph}(\mathsf{Axis}^{3}_{3})$
$\mathsf{Im}(\mathsf{Axis}^{3}$)
Figure 2: The $3$ graphs of the coordinate functions $\mathsf{Axis}^{3}_{i}$ with the non-constant part marked in red, and the entire image of the $3$-dimensional path $\mathsf{Axis}^{3}$ on the right. Compare also Figure 4 for the correspondending axis membrane.
In the following we recall the ‘nearest’ extension of the path signature (2) for multiple parameters and comment on its various modifications in Remark 2.3.
Definition 2.2.
Let $V$ be a finite dimensional vector space and $X:[0,1]^{p}\to V$ a Lipschitz continuous map. Let $\partial_{1\dots p}:=\frac{\partial^{p}}{\partial x_{1}\dots\partial x_{p}}$ be the linear partial differential operator. Then, since both the integral and $\partial_{1\dots p}$ are linear operators, the following linear form $\sigma=\sigma(X)\in T(V^{*})^{*}$ is well-defined:
$\sigma:T(V^{*})$ $\to\mathbb{R},$
$\alpha_{1}\otimes\dots\otimes\alpha_{k}$ $\mapsto\int_{\Delta^{p}}\partial_{1\dots p}(\alpha_{1}\circ X)(t_{11},\dots,t_{1p})\dots\partial_{1\dots p}(\alpha_{k}\circ X)(t_{k1},\dots,t_{kp})\,\mathrm{d}\mathbf{t}$
where $\mathrm{d}\mathbf{t}=\mathrm{d}t_{11}\dots\mathrm{d}t_{1p}\dots\mathrm{d}t_{k1}\dots\mathrm{d}t_{kp}$ and $\Delta^{p}=\Delta\times\dots\times\Delta$ is the Cartesian product of $p$ simplices. We call $\sigma$ the iterated-integrals signature of $X$. The $k$-th level signature $\sigma^{(k)}(X)$ is the restriction of $\sigma(X)$ along $(V^{*})^{\otimes k}\to T(V^{*})$, in particular it is naturally an element of $V^{\otimes k}$.
Remark 2.3.
  1. 1.
    For the one-parameter framework $p=1$ in $d$-dimensional space $V=\mathbb{R}^{d}$ equipped with standard basis $e_{1},\dots,e_{d}$ we obtain the classical formulation (2) for paths. We refer the interested reader to the text books [23, 21].
  2. 2.
    When ‘replacing’ our square increments $\partial_{1\dots p}$ with Jacobian minors we obtain the $\mathsf{id}$-signature introduced in [26].
  3. 3.
    For membranes, that is $p=2$ parameters and standard basis $e_{1},\dots,e_{d}$ of $V=\mathbb{R}^{d}$, we obtain the $\mathsf{id}$-signature according to [14]. Its level $k$ component can be stored in the $d\times\dots\times d$ array,
    $\sigma^{(k)}_{i_{1},\dots,i_{k}}$ $:=\sigma(e_{i_{1}}\otimes\dots\otimes e_{i_{k}})$
    $=\int\limits_{\begin{subarray}{c}0<s_{1}<\dots<s_{k}<1\\ 0<t_{1}<\dots<t_{k}<1\end{subarray}}\partial_{12}X_{i_{1}}(s_{1},t_{1})\dots\partial_{12}X_{i_{k}}(s_{k},t_{k})\,\mathrm{d}s_{1}\mathrm{d}t_{1}\dots\mathrm{d}s_{k}\mathrm{d}t_{k}.$
    In particular, for level $k=2$, we obtain our signature matrix $\sigma=\sigma^{(2)}$ from the introduction, (2).
  4. 4.
    When ‘replacing’ integrals by sums, and $\partial_{12}$ by its discrete counterpart, we obtain the two-parameter sums signature [18], evaluated at all diagonal matrix compositions. For this compare also [14, Example 4.16].
Remark 2.4.
It is clear from the definition that the signatures of any two Lipschitz continuous maps $X,Y:[0,1]^{p}\to V$ agree if $\partial_{1...p}(\alpha\circ(X-Y))=0$ for all $\alpha\in V^{*}$. For $p=1$ this is just the translation invariance of the signature, for membranes, i.e. $p=2$, this is invariance under adding terms only depending on one of the variables. In particular, if $X:[0,1]^{2}\to V$ is a membrane, we can always find a membrane $X_{\mathrm{red}}$ which vanishes on $[0,1]\times\{0\}\cup\{0\}\times[0,1]$ and has the same signature as $X$, by letting $X_{\mathrm{red}}(s,t):=X(s,t)-X(0,t)-X(s,0)+X(0,0)$. See also Lemma 3.18 and Figure 4 for an illustration of piecewise bilinear membranes.
We provide some first examples for membranes ($p=2$).
Example 2.5.
  1. 1.
    For fixed $u\in\mathbb{R}^{d}$, the $k$-th signature tensor of the $d$-dimensional bilinear membrane $[0,1]^{2}\rightarrow\mathbb{R}^{d},(s,t)\mapsto ust$ is given by
    $\sigma^{(k)}_{i_{1},\dots,i_{k}}=\frac{u_{i_{1}}\dots u_{i_{k}}}{(k!)^{2}}$ (13)
    where $1\leq i_{1},\dots,i_{k}\leq d$. This formula follows from Corollary 3.5 or can be verified by an immediate computation.
  2. 2.
    We recall the $2$-dimensional polynomial membrane $X=A\,\mathsf{Mom}^{2,2}$ from the introduction, (7). With integration over the differentials of the two-parameter moment coordinates,
    $\int_{0}^{1}\int_{0}^{1}\,\mathrm{d}s\mathrm{d}t=\int_{0}^{1}\int_{0}^{1}2t\,\mathrm{d}s\mathrm{d}t=\int_{0}^{1}\int_{0}^{1}2s\,\mathrm{d}s\mathrm{d}t=\int_{0}^{1}\int_{0}^{1}4st\,\mathrm{d}s\mathrm{d}t=1,$
    we obtain, together with linearity, the level $1$ signature,
    $\sigma^{(1)}(X)=\begin{pmatrix}a_{1,1}+a_{1,2}+a_{1,3}+a_{1,4}\\ a_{2,1}+a_{2,2}+a_{2,3}+a_{2,4}\end{pmatrix}\in\mathbb{R}^{2}.$
    In Example 3.7 and 3.11 we provide closed formulas for its level $2$ and $3$ signature tensors, respectively.
As in the previous example, we can use linearity of the integral to compute level $k$ signatures of transformed functions by corresponding actions on tensors.
Lemma 2.6 (Equivariance).
Let $V,W$ be finite dimensional $\mathbb{R}$-vector spaces, and $X:[0,1]^{p}\to V$ be a Lipschitz continuous map. If $A:V\to W$ is linear, then $\sigma(AX)=\sigma(X)\circ T(A^{*})$.
Proof.
This follows immediately from the definition, since for $w=w_{1}\otimes\dots\otimes w_{k}$ (where $w_{i}\in W^{*}$) we have $w_{i}\circ(A\circ X)=(w_{i}\circ A)\circ X=A^{*}(w_{i})\circ X$. ∎
In particular, Lemma 2.6 implies that $\sigma^{(k)}(AX)=A^{\otimes k}(\sigma^{(k)}(X))$ for all $k$. If $V=\mathbb{R}^{d}$ we can write this in the Tucker format $\sigma^{(k)}(AX)=[\![\sigma^{(k)}(X);A,\dots,A]\!]$. Explicitly, if $k=2$ we have
$\sigma^{(k)}(AX)=A\sigma^{(k)}(X)A^{\top},$
as was already mentioned in the introduction (6). The equivariance of signatures is a key property used in [41], allowing the computation of signature tensors for various families of paths from a small collection of dictionaries, also called core tensors. In the following section we introduce such dictionaries for membranes.
3 Product membranes
Some membranes arise from combining paths: given a path $[0,1]\to V$ and another path $[0,1]\to W$ we can consider the Cartesian product $[0,1]^{2}\to V\times W$. Composing this with any suitable function $\phi:V\times W\to Z$ into another vector space $Z$ yields a membrane. The first goal of this section is to show that if $\phi$ is bilinear, then we can express the signature of the resulting membrane in terms of the signatures of the defining paths. This is particularly important for us, as we will see that both polynomial and piecewise bilinear membranes can be obtained in this way. Recalling equivariance of the signature (Lemma 2.6) and that any bilinear map $V\times W\to Z$ factors over a linear map $V\otimes W\to Z$, it suffices to understand the signature of the following type of membrane:
Definition 3.1.
Let $V$ and $W$ be finite dimensional vector spaces. The product membrane of two paths $X:[0,1]\to V$ and $Y:[0,1]\to W$ is the membrane in the tensor product $V\otimes W$ induced by the Cartesian product:
$X\boxtimes Y:[0,1]^{2}\to V\times W\to V\otimes W.$
In other words, $X\boxtimes Y(s,t):=X(s)\otimes Y(t)$ for all $(s,t)\in[0,1]^{2}$.
Example 3.2.
  1. 1.
    Consider the path $X$ mapping $s\mapsto(s,s^{2})$ and the path $Y$ mapping $t\mapsto(e^{t},\log(1+t))$. Writing elements of $\mathbb{R}^{2}\otimes\mathbb{R}^{2}$ as matrices, the product membrane $X\boxtimes Y$ is given by
    $(s,t)\mapsto\begin{pmatrix}se^{t}&s\log(1+t)\\ s^{2}e^{t}&s^{2}\log(1+t)\end{pmatrix}\in\mathbb{R}^{2\times 2}.$
  2. 2.
    The moment membrane of multidegree $(m,n)$ in the introduction (4) is the product
    $\mathsf{Mom}^{m,n}=\mathsf{Mom}^{m}\boxtimes\mathsf{Mom}^{n}:[0,1]^{2}$ $\to\mathbb{R}^{m}\otimes\mathbb{R}^{n},$ (14)
    where the moment path $\mathsf{Mom}^{m}$ is according to Example 2.1. Note that we sometimes omit the last vectorization, i.e., we identity $\mathbb{R}^{m\times n}$ with $\mathbb{R}^{mn}$ whenever the context allows this. For this compare also Corollary 3.6. Note that every $d$-dimensional polynomial membrane $X$ with multidegree $(m,n)$ without terms of degree $\leq(1,1)$ is of the form $X=A\circ\mathsf{Mom}^{m,n}$, where $A$ is a linear map, e.g. our running Example 2.5 with $m=n=d=2$.
3.1 Factorizations into path signatures
We state our main result of this section, Proposition 3.3, showing that the signature is multiplicative with respect to products of paths (Definition 3.1). In Corollary 3.6, we rephrase this for the matrix case $k=2$, with explicit vectorizations and the use of the Kronecker matrix product. We illustrate these results in Examples 3.11 and 3.7, respectively.
Proposition 3.3.
Let $X\boxtimes Y:[0,1]^{2}\to V\otimes W$ be a product membrane. Then, as linear maps,
$\sigma(X\boxtimes Y)=\sigma(X)\otimes\sigma(Y)$
in the following sense: the map $\sigma(X\boxtimes Y)$ factors as
[…]
where the second map is induced by the universal property of the tensor algebra, and the last map is given by $\sigma(X)\otimes\sigma(Y)$. More explicitly, for all $k\geq 1$, $\alpha_{i}\in V^{*}$, $\beta_{i}\in W^{*}$ we have
$\sigma(X\boxtimes Y),(\alpha_{1}\otimes\beta_{1})\otimes\dots\otimes(\alpha_{k}\otimes\beta_{k})\right>$ $=$
$\sigma(X),\alpha_{1}\otimes\dots\otimes\alpha_{k}\right>$ $\sigma(Y),\beta_{1}\otimes\dots\otimes\beta_{k}\right>$
Caveat 3.4.
Note that the tensor product of linear forms $\sigma(X):T(V^{*})\to\mathbb{R}$, $\sigma(Y):T(V^{*})\to\mathbb{R}$ is not the tensor product in the tensor series space $T(V^{*})$; the formula above should not be confused with Chen’s formula for the signature of a concatenation of paths.
Before proving Proposition 3.3, we provide examples and discuss several implications. As a first application, we compute the signature of product membranes where one path is $1$-dimensional, explaining for instance our example (13) using path signatures.
Corollary 3.5.
Let $X:[0,1]\to\mathbb{R}$ be a $1$-dimensional path, and $Y:[0,1]\to W$ be arbitrary. Then $\sigma(X\boxtimes Y)$ is a map $T(\mathbb{R}\otimes W^{*})\to\mathbb{R}$, so we can view it as a map $T(W^{*})\to\mathbb{R}$. Then it sends a simple tensor $w=w_{1}\otimes\dots\otimes w_{\ell}$ to
$\sigma(X)(e_{1}^{\otimes\ell})\cdot\sigma(Y)(w)=\frac{(X(1)-X(0))^{\ell}}{\ell!}\cdot\sigma(Y)(w).$
In particular, $\sigma^{(k)}(X\boxtimes Y)=\frac{1}{k!}(X(1)-X(0))^{k}\cdot\sigma^{(k)}(Y)$, so $\sigma^{(k)}(X\boxtimes Y)$ and $\sigma^{(k)}(Y)$ differ only by some scalar for any $k$.
Proof.
Unravel the factorization of $\sigma(X\boxtimes Y)$ in Proposition 3.3. ∎
If $X:[0,1]^{2}\to V\otimes W$ is a product membrane, it can be convenient to use a vectorization $\nu:V\otimes W\cong\mathbb{R}^{\dim V}$, for example to obtain an actual signature matrix. Using Lemma 2.6, we see that $\sigma(\nu X)$ then factors as
$T(\mathbb{R}^{d})\to T((V\otimes W)^{*})\to T(V^{*})\otimes T(W^{*})\to\mathbb{R}$
where the first map is $T(\nu^{*})$.
As a special case, we obtain compatibility of the signature with the Kronecker product of paths:
Corollary 3.6.
Let $X:[0,1]\to\mathbb{R}^{m}$ and $Y:[0,1]\to\mathbb{R}^{n}$ be two paths and let $\mathsf{kron}(X,Y):[0,1]^{2}\to\mathbb{R}^{mn}$ denote the membrane that maps $(s,t)$ to the Kronecker product of the column vectors $X(s)$ and $Y(t)$. Then
$\sigma^{(2)}(\mathsf{kron}(X,Y))$ $=\sigma^{(2)}(X)\otimes\sigma^{(2)}(Y)$
where $\otimes$ denotes the Kronecker product of matrices.
Example 3.7.
We return to our running Example 2.5 with $X=A\,\mathsf{Mom}^{2,2}$ and matrix $A=(a_{i,j})_{1\leq i\leq 2,1\leq j\leq 4}$. With Corollary 3.6 and (12) we get the core matrix of the moment membrane,
$C^{(2)}:=\sigma^{(2)}(\mathsf{vec}\circ\mathsf{Mom}^{2,2})=\sigma^{(2)}(\mathsf{Mom}^{2})\otimes\sigma^{(2)}(\mathsf{Mom}^{2})=\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}\otimes\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}$
as it is presented in the introduction, (5). With $S^{(2)}:=\sigma^{(2)}(X)=AC^{(2)}A^{\top}$ we provide the three missing homogeneous entries,
$S^{(2)}_{2,1}$ $=\frac{1}{4}a_{1,1}a_{2,1}+\frac{1}{6}a_{1,1}a_{2,2}+\frac{1}{6}a_{1,1}a_{2,3}+\frac{1}{9}a_{1,1}a_{2,4}+\frac{1}{3}a_{2,1}a_{1,2}+\frac{1}{3}a_{2,1}a_{1,3}$
$\quad+\frac{4}{9}a_{2,1}a_{1,4}+\frac{1}{4}a_{1,2}a_{2,2}+\frac{2}{9}a_{1,2}a_{2,3}+\frac{1}{6}a_{1,2}a_{2,4}+\frac{2}{9}a_{2,2}a_{1,3}+\frac{1}{3}a_{2,2}a_{1,4}$
$\quad+\frac{1}{4}a_{1,3}a_{2,3}+\frac{1}{6}a_{1,3}a_{2,4}+\frac{1}{3}a_{2,3}a_{1,4}+\frac{1}{4}a_{1,4}a_{2,4}$
$S^{(2)}_{1,2}$ $=\frac{1}{4}a_{1,1}a_{2,1}+\frac{1}{3}a_{1,1}a_{2,2}+\frac{1}{3}a_{1,1}a_{2,3}+\frac{4}{9}a_{1,1}a_{2,4}+\frac{1}{6}a_{2,1}a_{1,2}+\frac{1}{6}a_{2,1}a_{1,3}$
$\quad+\frac{1}{9}a_{2,1}a_{1,4}+\frac{1}{4}a_{1,2}a_{2,2}+\frac{2}{9}a_{1,2}a_{2,3}+\frac{1}{3}a_{1,2}a_{2,4}+\frac{2}{9}a_{2,2}a_{1,3}+\frac{1}{6}a_{2,2}a_{1,4}$
$\quad+\frac{1}{4}a_{1,3}a_{2,3}+\frac{1}{3}a_{1,3}a_{2,4}+\frac{1}{6}a_{2,3}a_{1,4}+\frac{1}{4}a_{1,4}a_{2,4}$
and
$S^{(2)}_{2,2}$ $=\frac{1}{4}a_{2,1}^{2}+\frac{1}{2}a_{2,1}a_{2,2}+\frac{1}{2}a_{2,1}a_{2,3}+\frac{5}{9}a_{2,1}a_{2,4}+\frac{1}{4}a_{2,2}^{2}+\frac{4}{9}a_{2,2}a_{2,3}+\frac{1}{2}a_{2,2}a_{2,4}$
$\quad+\frac{1}{4}a_{2,3}^{2}+\frac{1}{2}a_{2,3}a_{2,4}+\frac{1}{4}a_{2,4}^{2}$
using equivariance, Lemma 2.6. The polynomial $S_{1,1}^{(2)}$ is already provided in the introduction.
We illustrate the membrane $X$ in the first row of Figure 3, for $A=\begin{pmatrix}1&-1&1&1\\ 1&1&0&-1\end{pmatrix}$. In particular, we provide the graphs of the coordinate functions $\mathsf{graph}(X_{1})$ and $\mathsf{graph}(X_{2})$, and a plot of the image $\mathsf{Im}(X)$ in $\mathbb{R}^{2}$. Here, the colors correspond to the image of rectangles in a uniform partition of $[0,1]^{2}$, e.g. $X([0,\frac{1}{3}]\times[0,\frac{1}{2}])$ in pink, $X([\frac{1}{3},\frac{2}{3}]\times[0,\frac{1}{2}])$ in orange, or $X([\frac{2}{3},1]\times[\frac{1}{2},1])$ in violet. Note that $X$ maps three sides of the square $[0,1]^{2}$ to the origin. The right-most picture shows an ”untangled” version of the membrane $X$ which has the same signature, cf. Remark 2.4.
$\mathsf{graph}(X_{1})$
$\mathsf{graph}(X_{2})$
$\mathsf{Im}(X)$
$\mathsf{Im}(X+\begin{pmatrix}s\\ 2t\end{pmatrix})$
$\mathsf{graph}(Y_{1})$
$\mathsf{graph}(Y_{2})$
$\mathsf{graph}(Y_{3})$
$\mathsf{Im}(Y)$
Figure 3: Plots of the polynomial membrane $X$ in $\mathbb{R}^{2}$ from Example 3.7 and a piecewise bilinear membrane $Y$ interpolating $12$ points in $\mathbb{R}^{3}$, represented by black dots. Let $X_{1},X_{2},Y_{1},Y_{2}$ and $Y_{3}$ denote the coordinate functions of $X$ and $Y$, respectively.
Proof of Corollary 3.6.
Choosing the vectorization $\nu:\mathbb{R}^{m}\otimes\mathbb{R}^{n}\to\mathbb{R}^{mn},e_{i}\otimes e_{j}\mapsto e_{n(i-1)+j}$, we see that $\mathsf{kron}=\nu\circ X\boxtimes Y$. Then we have
$\sigma(\nu(X\boxtimes Y)),e_{i}\otimes e_{j}\right>=\left<\sigma(X\boxtimes Y),\nu^{*}(e_{i})\otimes\nu^{*}(e_{j})\right>$
[…]
where $i_{1}=\lceil\frac{i}{n}\rceil$, $j_{1}=\lceil\frac{j}{n}\rceil$, $i_{2}=(i-1\text{ mod }n)+1$, $j_{2}=(j-1\text{ mod }n)+1$. In terms of signature matrices, this amounts to the identity
$\sigma^{(2)}(\nu(X\boxtimes Y))=\sigma^{(2)}(X)\otimes\sigma^{(2)}(Y).$
As another consequence of Proposition 3.3 and equivariance we obtain the formula from [14, Example 4.2] for the Hadamard product of two paths:
Corollary 3.8.
If $V=W=\mathbb{R}^{d}$ then
$\sigma(\mu(X\boxtimes Y))=\sigma(X)\cdot\sigma(Y)$ (15)
as morphisms $T(\mathbb{R}^{d})\to\mathbb{R}$, and where $\mu:\mathbb{R}^{d}\otimes\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the componentwise multiplication map.
Proof.
The dual map $\mu^{*}$ sends $e_{i}$ to $e_{i}\otimes e_{i}$. ∎
We close this section with our omitted proof, followed by an example for higher tensor levels, and an implication for the computational complexity of the signature of a polynomial membrane.
Proof of Proposition 3.3.
The isomorphism $T(V^{*}\otimes W^{*})\to T((V\otimes W)^{*})$ displayed in the statement is induced by the map that sends $\alpha\otimes\beta$ to the form $\alpha\cdot\beta$ on $V\otimes W$. Now, by definition, $\sigma(X)$ maps a simple tensor $\alpha_{1}\cdot\beta_{1}\otimes\dots\otimes\alpha_{k}\cdot\beta_{k}$ to
$\int_{\Delta}\int_{\Delta}\prod_{i}\partial_{12}z_{i}(s_{i},t_{i})\,\mathrm{d}s_{1}\dots\mathrm{d}s_{k}\mathrm{d}t_{1}\dots\mathrm{d}t_{k}$
where $z_{i}(s,t):=(\alpha_{i}\circ X)(s)\cdot(\beta_{i}\circ Y)(t)$. Here we reordered the differential forms which is allowed because we are integrating over the product simplex $\Delta^{2}$. In particular, we have $(\partial_{12}z_{i})(s,t)=\partial(\alpha_{i}\circ X)(s)\cdot\partial(\beta_{i}\circ Y)(t)$. By definition,
$\langle\sigma(X^{(1)}),\alpha_{1}\otimes\dots\otimes\alpha_{k}\rangle=\int_{\Delta}\prod_{i}\partial(\alpha_{i}\circ X)(s_{i})\,\mathrm{d}s_{1}\dots\mathrm{d}s_{k}$
and similarly for $Y$; thus reordering the integral reveals the claimed equality. ∎
Remark 3.9.
Proposition 3.3 and its proof immediately generalize to higher $p$-parameter patches $[0,1]^{p}\rightarrow\mathbb{R}^{d}$.
In order to provide an explicit computation for higher tensor levels $k\geq 3$, we introduce some further notation.
Remark and Notation 3.10.
If $X:[0,1]\to\mathbb{R}^{m}$ and $Y:[0,1]\to\mathbb{R}^{n}$ are two paths, then by Proposition 3.3 the signature of $X\boxtimes Y$ is a linear form on $T((\mathbb{R}^{m}\otimes\mathbb{R}^{n})^{*})$ which can be identified with the free associative algebra $\mathbb{R}\langle(i,j)\rangle$ over tuples of letters $i\in[m]$ and $j\in[n]$. Thus, in the following we also write $(i_{1},j_{1})\dots(i_{k},j_{k})$ for the tensor $(e_{i_{1}}\otimes e_{j_{1}})^{*}\otimes\dots\otimes(e_{i_{k}}\otimes e_{j_{k}})^{*}$.
Example 3.11.
With Proposition 3.3, our identification in 3.10, and the closed formula for signatures of polynomial paths in Example 2.1, we obtain a closed formula for the level $k$ tensor of the moment membrane $\mathsf{Mom}^{m,n}$ from (14),
$\langle\sigma(\mathsf{Mom}^{m,n}),(i_{1},j_{1})\dots(i_{k},j_{k})\rangle=\frac{i_{2}\dots i_{k}}{(i_{1}+i_{2})\dots(i_{1}+\dots+i_{k})}\frac{j_{2}\dots j_{k}}{(j_{1}+j_{2})\dots(j_{1}+\dots+j_{k})}$
where $i_{1},\dots,i_{k}\in[m]$ and $j_{1},\dots,j_{k}\in[n]$. For instance for the $2$-dimensional polynomial membrane from our running example Example 2.5, we obtain the level $k=3$ core tensor
$C^{(3)}:=\scriptsize\left(\begin{array}[]{cccc|cccc|cccc|cccc}\frac{1}{36}&\frac{1}{36}&\frac{1}{36}&\frac{1}{36}&\frac{1}{24}&\frac{2}{45}&\frac{1}{24}&\frac{2}{45}&\frac{1}{24}&\frac{1}{24}&\frac{2}{45}&\frac{2}{45}&\frac{1}{16}&\frac{1}{15}&\frac{1}{15}&\frac{16}{225}\\[4.0pt] \frac{1}{72}&\frac{1}{60}&\frac{1}{72}&\frac{1}{60}&\frac{1}{45}&\frac{1}{36}&\frac{1}{45}&\frac{1}{36}&\frac{1}{48}&\frac{1}{40}&\frac{1}{45}&\frac{2}{75}&\frac{1}{30}&\frac{1}{24}&\frac{8}{225}&\frac{2}{45}\\[4.0pt] \frac{1}{72}&\frac{1}{72}&\frac{1}{60}&\frac{1}{60}&\frac{1}{48}&\frac{1}{45}&\frac{1}{40}&\frac{2}{75}&\frac{1}{45}&\frac{1}{45}&\frac{1}{36}&\frac{1}{36}&\frac{1}{30}&\frac{8}{225}&\frac{1}{24}&\frac{2}{45}\\[4.0pt] \frac{1}{144}&\frac{1}{120}&\frac{1}{120}&\frac{1}{100}&\frac{1}{90}&\frac{1}{72}&\frac{1}{75}&\frac{1}{60}&\frac{1}{90}&\frac{1}{75}&\frac{1}{72}&\frac{1}{60}&\frac{4}{225}&\frac{1}{45}&\frac{1}{45}&\frac{1}{36}\end{array}\right)$
of shape $4\times 4\times 4$, and with equivariance (Lemma 2.6), the level $k=3$ signature tensor $\sigma^{(3)}(X)=[\![C^{(3)};A,A,A]\!]$ of shape $2\times 2\times 2$. We provide its first homogeneous entry,
${\sigma^{(3)}(X)}_{1,1,1}$ $=\frac{1}{36}a_{1,1}^{3}+\frac{1}{12}a_{1,1}^{2}a_{1,2}+\frac{1}{12}a_{1,1}^{2}a_{1,3}+\frac{7}{72}a_{1,1}^{2}a_{1,4}+\frac{1}{12}a_{1,1}a_{1,2}^{2}$
$\quad+\frac{11}{72}a_{1,1}a_{1,2}a_{1,3}+\frac{13}{72}a_{1,1}a_{1,2}a_{1,4}+\frac{1}{12}a_{1,1}a_{1,3}^{2}+\frac{13}{72}a_{1,1}a_{1,3}a_{1,4}$
$\quad+\frac{89}{900}a_{1,1}a_{1,4}^{2}+\frac{1}{36}a_{1,2}^{3}+\frac{5}{72}a_{1,2}^{2}a_{1,3}+\frac{1}{12}a_{1,2}^{2}a_{1,4}$
$\quad+\frac{5}{72}a_{1,2}a_{1,3}^{2}+\frac{34}{225}a_{1,2}a_{1,3}a_{1,4}+\frac{1}{12}a_{1,2}a_{1,4}^{2}+\frac{1}{36}a_{1,3}^{3}$
$\quad+\frac{1}{12}a_{1,3}^{2}a_{1,4}+\frac{1}{12}a_{1,3}a_{1,4}^{2}+\frac{1}{36}a_{1,4}^{3}.$
The other (polynomial) entries of $\sigma^{(3)}(X)$ are similar and are therefore omitted here.
Inspired by [18, Theorem 4.5], we obtain the following computational implication.
Corollary 3.12.
For every $d$-dimensional polynomial membrane $X$ of multi-degree $(m,n)$ we can compute $\sigma^{(k)}(X)$ in $\mathcal{O}(m^{k}n^{k}+kd^{k+1}mn)$ elementary operations.
Proof.
We use the closed formula in Example 3.11 and $\mathsf{vec}^{\otimes k}$ to obtain the core tensor $C^{(k)}$ in $\mathcal{O}(m^{k}n^{k})$. The remaining follows with equivariance. ∎
Remark 3.13.
We can generalize Corollary 3.12 for any class of membranes which are induced via linear transforms of product membranes, e.g., piecewise bilinear membranes from the next section. Note that for the latter we can improve the analogous computational complexity bound by Theorem 3.19.
3.2 Dictionaries for bilinear interpolation
Definition 3.14.
A piecewise bilinear membrane is a membrane $X$ such that there are some $m,n\in\mathbb{N}$ with the property that $X$ is biaffine on all squares $[\frac{i-1}{m},\frac{i}{m}]\times[\frac{j-1}{n},\frac{j}{n}]$ with $i\in[m]$ and $j\in[n]$. We say that $X$ has order $(m,n)$.
Our interest in this class of membranes is threefold: from the path case, examined in [2], we expect a certain relation of this class of membranes to the polynomial membranes in terms of their signatures, e.g. Theorem 4.1. We also know that the recovery problem is solvable for piecewise linear paths of low order from [41, Corollary 6.3], and we would like to obtain a similar result for membranes (5.4). Finally, they form a rather natural class that can be used for the interpolation of discrete data to fit into the frame of iterated integrals, e.g. image data used in [48].
Bilinear interpolation of $d$-dimensional discrete data, arranged on an $(m+1)\times(n+1)$ grid, is a well-known procedure, e.g. in image processing [6, Chapter 3.7.2] or in the field of finite elements [49, Chapters 5.1.3.2 and 6.2.2.1], [20, Chapter 1.3.2]. The approach involves considering the four (unique) nodal bilinear basis functions on the reference rectangle, where each function is $1$ at one corner and 00 at all others. The nodal basis on any arbitrary rectangle is then obtained through a linear transformation to the reference rectangle. Bilinear interpolation is simply the linear combination of the nodal basis functions, where the coefficients correspond exactly to the function values of the function being interpolated at the four grid points.
As in the polynomial case (14) we provide a product dictionary for piecewise bilinear membranes, so that its signature can be computed via Proposition 3.3. For this, we have already recalled in Example 2.1 the canonical axis path $\mathsf{Axis}^{m}$, which is illustrated in Figure 2 for $m=3$. The resulting axis membrane takes the role of the above mentioned bases for interpolation. It is illustrated for order $(3,2)$ in Figure 4.
Definition 3.15.
The (canonical) axis membrane of order $(m,n)$ is the product
$\mathsf{Axis}^{m,n}:=\mathsf{Axis}^{m}\boxtimes\mathsf{Axis}^{n}:[0,1]^{2}$ $\to\mathbb{R}^{m}\otimes\mathbb{R}^{n},$ (16)
according to Definition 3.1.
Any piecewise linear path with $m$ segments starting at the origin can be realized through a suitable linear transformation of $\mathsf{Axis}^{m}$. We want to show that the analogue is true for membranes, i.e., (16) is a suitable dictionary for all piecewise bilinear membranes of order $(m,n)$ which map the coordinate axes $[0,1]\times 0\cup 0\times[0,1]$ to the origin. Let us first describe the axis membrane a bit more explicitly.
$\mathsf{Axis}^{3,2}_{1,1}$
$\mathsf{Axis}^{3,2}_{2,1}$
$\mathsf{Axis}^{3,2}_{3,1}$
$A\circ\mathsf{Axis}_{3,2}$
$\mathsf{Axis}^{3,2}_{1,2}$
$\mathsf{Axis}^{3,2}_{2,2}$
$\mathsf{Axis}^{3,2}_{3,2}$
$R$
Figure 4: On the left we have the $6$ graphs of the coordinate functions from the axis membrane $\mathsf{Axis}^{3,2}$, with the non-linear rectangle marked in red, and the equidistant support points from (3.2) in black. On the right we illustrate the decomposition of the $1$-dimensional piecewise bilinear membrane $Y_{1}$ from Figure 3 into the linear transform of the axis membrane $A\circ\mathsf{Axis}^{3,2}$ and the linear part $R$ according to Lemma 3.18.
Lemma 3.16.
We consider the codomain from the axis membrane (16) as a space of matrices, that is $\mathsf{Axis}^{m,n}:[0,1]^{2}\rightarrow\mathbb{R}^{m\times n}$. Then each of its coordinate functions is piecewise bilinear,
$\mathsf{Axis}^{m,n}_{i,j}(s,t):=\scriptsize\begin{cases}0&s\leq\frac{i-1}{m}\text{ or }t\leq\frac{j-1}{n}\\ mnst+m(1-j)s+n(1-i)t+(i-1)(j-1)&\frac{i-1}{m}<s\leq\frac{i}{m}\text{ and }\frac{j-1}{n}<t\leq\frac{j}{n}\\ nt+1-j&\frac{i}{m}<s\text{ and }\frac{j-1}{n}<t\leq\frac{j}{n}\\ ms+1-i&\frac{i-1}{m}<s\leq\frac{i}{m}\text{ and }\frac{j}{n}<t\\ 1&\frac{i}{m}<s\text{ and }\frac{j}{n}<t,\end{cases}$
where $i\in[m]$, $j\in[n]$ and $s,t\in[0,1]$.
Proof.
Despite this being well-known we provide a bilinear interpolation in each of the $mn$ dimensions, i.e., we search for coefficients $c\in\mathbb{R}^{4}$ such that
$\psi_{i,j}:\left[\frac{i-1}{m},\frac{i}{m}\right]\times\left[\frac{j-1}{n},\frac{j}{n}\right]$ $\rightarrow\mathbb{R},(s,t)\mapsto c_{1}st+c_{2}s+c_{3}t+c_{4}$
satisfies the system
$\psi_{i,j}\left(\frac{i-1}{m},\frac{j-1}{n}\right)$ $=0,\quad\psi_{i,j}\left(\frac{i}{m},\frac{j-1}{n}\right)=0,$
$\psi_{i,j}\left(\frac{i-1}{m},\frac{j}{n}\right)$ $=0,\quad\psi_{i,j}\left(\frac{i}{m},\frac{j}{n}\right)=1.$ (17)
By solving (3.2) in $c$ we obtain $c=\left(mn,m(1-j),n(1-i),(i-1)(j-1)\right)$ and hence $\psi=\mathsf{Axis}^{m,n}$ when restricted on the square $\left[\frac{i-1}{m},\frac{i}{m}\right]\times\left[\frac{j-1}{n},\frac{j}{n}\right]$. The claim follows with a similar (linear) system for the remaining squares. ∎
With the product structure in (16) we can use Proposition 3.3 to compute its core signature tensor with the help of the core signature tensors of axis paths.
Example 3.17.
From [2, Example 2.1] we know that $\sigma(\mathsf{Axis}^{m})_{i_{1},\dots,i_{k}}$ is zero unless $i_{1}\leq\dots\leq i_{k}$, and $\frac{1}{k!}$ times the number of distinct permutations of the string $i_{1}\dots i_{k}$ otherwise. Explicitly, with (11) the signature matrix is
$\langle\sigma(\mathsf{Axis}^{m,n}),(i_{1},j_{1})(i_{2},j_{2})\rangle:=\begin{cases}1&i_{1}<i_{2}\text{ and }j_{1}<j_{2}\\ \frac{1}{2}&i_{1}=i_{2}\text{ and }j_{1}<j_{2}\\ \frac{1}{2}&i_{1}<i_{2}\text{ and }j_{1}=j_{2}\\ \frac{1}{4}&i_{1}=i_{2}\text{ and }j_{1}=j_{2}\\ 0&\text{otherwise.}\end{cases}$
Recall from Remark 2.4 that we can always restrict our computations to membranes which are 00 on the coordinate axes.
Lemma 3.18.
Every piecewise bilinear membrane $X$ has a (unique) decomposition
$X=A\circ\mathsf{Axis}^{m,n}+R$ (18)
where $\mathsf{Axis}^{m,n}$ is according to (16), $A$ is a linear transform, and $R$ is a piecewise linear (not bilinear) membrane in the two arguments $s$ and $t$. Then,
$\sigma(X)=\sigma(A\circ\mathsf{Axis}^{m,n}).$
Proof.
Clearly, $\{\mathsf{Axis}^{m,n}_{i,j}\mid i,j\}$ is an $\mathbb{R}$-basis of the space of $1$-dimensional piecewise bilinear membranes which are 00 on $[0,1]\times 0\cup 0\times[0,1]$. We can extend this basis via $m+n-1$ functions which are piecewise linear in $s$ and constant in $t$, or vice versa. Applying a basis decomposition in each dimension, we obtain $A$ and $R$ in (18) for all $d$ and conclude by Remark 2.4. ∎
The decomposition is illustrated in Figure 4. See also [18, Lemma 2.22] for a similar result applied on discrete data, using two-parameter sums signatures.
Theorem 3.19.
Let $X$ be a piecewise bilinear membrane in $\mathbb{R}^{d}$ of order $(m,n)$ and $w\in T(\mathbb{R}^{d})$ a word of length $k$. Then we can compute $\langle\sigma(X),w\rangle$ in $\mathcal{O}(k^{2}mn^{k})$.
Proof.
With Lemma 3.18 we assume that $X$ is 00 on $\{0\}\times[0,1]\cup[0,1]\times\{0\}$. Let $\Lambda$ denote the piecewise linear path in $\mathbb{R}^{n}$ of order $n$ with control points $(0,e_{1},\dots,e_{n})$. Then we can write
$X(s,t)=\sum_{i=0}^{n}X(s,\frac{i}{n})\Lambda_{i}(t)$
Indeed, the right hand side is clearly piecewise bilinear of order $(m,n)$ and agrees with $X$ on all control points. We can write this as a linear transformation $A$ applied to the product membrane $Y\boxtimes\Lambda$ in $\mathbb{R}^{nd}\otimes\mathbb{R}^{n}$ where
$Y(s):=\begin{pmatrix}X(s,\frac{1}{n})\\ \vdots\\ X(s,1)\end{pmatrix}$
is a piecewise linear path of order $m$. The dual of $A$ is given via
$e_{i}^{*}\mapsto\sum_{j=1}^{n}e_{i+(j-1)d}^{*}\otimes e_{j}^{*}.$
In particular, we obtain a decomposition of $(A^{*})^{\otimes k}w$ into $n^{k}$ biwords of length $k$. By Proposition 3.3 and Chen’s identity (for paths) we can compute each of the corresponding signatures in $\mathcal{O}(k^{2}m)$. Thus, by Lemma 2.6, we obtain $\langle\sigma(X),w\rangle$ in $\mathcal{O}(k^{2}mn^{k})$. ∎
4 Varieties of two-parameter signature matrices
Our primary interest lies in the image of polynomial resp. piecewise bilinear membranes under the signature map at a specific level $k$. Here, we focus on the special case $k=2$. Following the approach in [2] we restrict to families of $d$-dimensional membranes,
$\left\{\sigma^{(2)}(X)\;\begin{array}[]{|l}X:[0,1]^{2}\rightarrow{\mathbb{R}}^{d}\text{ polynomial}\\ \text{with multi-degree}\leq(m,n)\end{array}\right\}$ (19)
and
$\left\{\sigma^{(2)}(X)\;\begin{array}[]{|l}X:[0,1]^{2}\rightarrow{\mathbb{R}}^{d}\text{ piecewise bilinear}\\ \text{of order }\leq(m,n)\end{array}\right\}$ (20)
with fixed $m,n\in\mathbb{N}$. By Lemma 2.6 and Lemma 3.18, both (19) and (20) can be considered as the image of a polynomial map of degree two,
$\sigma^{(2)}:\mathbb{R}^{d\times mn}\rightarrow\mathbb{R}^{d^{2}},A\mapsto ACA^{\top},$ (21)
where $C\in\mathbb{R}^{mn\times mn}$ is either the signature core matrix $\sigma^{(2)}(\nu\mathsf{Mom}^{m,n})$ (Example 3.11) for polynomial membranes or the axis core matrix $\sigma^{(2)}(\nu\mathsf{Axis}^{m,n})$ for the piecewise bilinear family (Example 3.17) where we choose $\nu:\mathbb{R}^{m}\otimes\mathbb{R}^{n}\to\mathbb{R}^{mn}$ as in the proof of Corollary 3.6. In particular, these signature images are semialgebraic subsets of $\mathbb{R}^{d\times d}$ by the Tarski-Seidenberg theorem ([5, Theorem 2.2.1]). For convenience, we omit $\nu$ from the notation in the following. Using our result on the signature of product membranes, we see that [2, Theorem 3.3] generalizes immediately:
Theorem 4.1.
For all $d,m$ and $n$, the semialgebraic sets (19) and (20) agree.
Proof.
By definition of the two sets, it suffices to find an invertible matrix $A$ such that $\sigma^{(2)}(\mathsf{Mom}^{m,n})=A\sigma^{(2)}(\mathsf{Axis}^{m,n})A^{\top}$. Now by Corollary 3.6, we have $\sigma^{(2)}(\mathsf{Mom}^{m,n})=\sigma^{(2)}(\mathsf{Mom}^{m})\otimes\sigma^{(2)}(\mathsf{Mom}^{n})$ and $\sigma^{(2)}(\mathsf{Axis}^{m,n})=\sigma^{(2)}(\mathsf{Axis}^{m})\otimes\sigma^{(2)}(\mathsf{Axis}^{n})$. Moreover, in the proof of [2, Theorem 3.3] it is shown that for arbitrary $\ell$ there is an invertible $\ell\times\ell$ matrix $H_{\ell}$ such that $\sigma^{(2)}(\mathsf{Mom}_{\ell})(w)=H_{\ell}\sigma^{(2)}(\mathsf{Axis}_{\ell})H_{\ell}^{\top}$. Thus, we can choose $A:=H_{m}\otimes H_{n}$. ∎
We want to study the polynomials that vanish on this set, recorded in the homogeneous prime ideals $P_{d,m,n}$. Equivalently, we examine the tightest outer approximation by an algebraic variety, i.e., we move to $\mathbb{C}$ and consider the Zariski closure $\mathcal{P}_{d,m,n}$ of (19) $=$ (20) with arbitrary $d,m,n\in\mathbb{N}$. For its vanishing ideal we consider the graph of (21), which is an algebraic variety with vanishing ideal
$I_{d,m,n}:=\langle(ACA^{\top})_{i,j}-s_{i,j}\mid 1\leq i,j\leq d\rangle.$
The projection to the target of (21) is then cut out by the prime ideal
$P_{d,m,n}:=I_{d,m,n}\cap\mathbb{C}[s_{i,j}\mid 1\leq i,j\leq d],$
and agrees with $\mathcal{P}_{d,m,n}$ by construction. Compare Example 4.2 for an step-by-step construction with explicit choices of $d,m$ and $n$. As in the path framework, the level $k=2$ signature matrices of polynomial and piecewise bilinear membranes coincide.
Example 4.2.
For explicit $d,m$ and $n$ we compute $P_{d,m,n}$ with the moment core matrix $C:=\sigma^{(2)}(\mathsf{Mom}^{m})\otimes\sigma^{(2)}(\mathsf{Mom}^{n})$ and with Gröbner bases, applying an elimination of variables procedure. We refer to [12] for an introduction into the latter. If $m=d=2$ and $n=1$, then $ACA^{\top}$ is
$\begin{bmatrix}\frac{1}{4}a_{1,1}^{2}+\frac{1}{2}a_{1,1}a_{1,2}+\frac{1}{4}a_{1,2}^{2}&\!\!\!\!\frac{1}{4}a_{1,1}a_{2,1}+\frac{1}{3}a_{1,1}a_{2,2}+\frac{1}{6}a_{2,1}a_{1,2}+\frac{1}{4}a_{1,2}a_{2,2}\\ \frac{1}{4}a_{1,1}a_{2,1}+\frac{1}{6}a_{1,1}a_{2,2}+\frac{1}{3}a_{2,1}a_{1,2}+\frac{1}{4}a_{1,2}a_{2,2}&\frac{1}{4}a_{2,1}^{2}+\frac{1}{2}a_{2,1}a_{2,2}+\frac{1}{4}a_{2,2}^{2}\end{bmatrix}$
and with elimination of variables,
$P_{2,2,1}=\langle 4s_{1,1}s_{2,2}-s_{2,1}^{2}-2s_{2,1}s_{1,2}-s_{1,2}^{2}\rangle$ (22)
is generated by the relation [2, eq. (5)] in $\mathbb{C}^{2\times 2}$ (the determinant of the symmetric part of the $2\times 2$-matrix $S$), according to Corollary 3.5. If, on the other hand $m=d=n=2$, then
$I_{2,2,1}=\langle S^{(2)}_{1,1}-s_{1,1},S^{(2)}_{2,1}-s_{2,1},S^{(2)}_{1,2}-s_{1,2},S^{(2)}_{2,2}-s_{2,2}\rangle$ (23)
with polynomials $S^{(2)}_{i,j}$ according to Example 3.7. With Gröbner bases we can show that $P_{d,m,n}$ is the zero ideal, thus $\dim(\mathcal{P}_{2,2,2})=4$.
Even for $m=n=2$, the last example illustrates that membranes generate larger varieties than in the path framework [2]. In fact, the dimension of $\mathcal{P}_{2,2,2}$ is maximal, providing an initial illustration of our main result, Theorem 4.7. We now proceed with some technical preparations.
Proposition 4.3.
The variety $\mathcal{P}_{d,m,n}$ is irreducible.
Proof.
By Lemma 2.6, $\mathcal{L}_{d,m,n}$ is the schematic image of $A\mapsto A^{\otimes 2}\sigma^{({2})}(\mathsf{Axis}^{m,n})$ where
$A\in\mathsf{Lin}(\mathbb{C}^{m}\otimes\mathbb{C}^{n},\mathbb{C}^{d})\cong\mathbb{C}^{mnd}$
is the variety of linear maps $\mathbb{C}^{m}\otimes\mathbb{C}^{n}\to\mathbb{C}^{d}$. Thus, as $\mathsf{Lin}(\mathbb{C}^{m}\otimes\mathbb{C}^{n},\mathbb{C}^{d})$ is irreducible, so is $\mathcal{P}_{d,m,n}$. ∎
Analogously to the stabilizing chain of varieties from [2, Theorem 5.7], we obtain a stabilizing grid:
Corollary 4.4.
For every $d\in\mathbb{N}$ there is a grid (8) where we can choose $M,N\in\mathbb{N}$ such that all row and column chains stabilize at (or before) $M$ resp. $N$, that is, for all $m,n\in\mathbb{N}$:
[…]
$\mathcal{P}_{d,1,n}\subseteq\dots\subseteq\mathcal{P}_{d,M-1,n}\subseteq\mathcal{P}_{d,M,n}=\mathcal{P}_{d,M+1,n}$ $=\dots$
Here $M$ and $N$ are constants and depend only on $k$ and $d$. We denote the limit in (8) by $\mathcal{U}_{d}:=\mathcal{P}_{d,M,N}\subseteq\mathbb{C}^{d^{2}}$.
Proof.
Follows with Proposition 4.3 and $\dim(\mathcal{P}_{d,m,n})\leq d^{2}$. ∎
Note that $\mathcal{P}_{d,m,n}=\mathcal{P}_{d,n,m}$ for all $m$ and $n$, since $\mathsf{Axis}^{m,n}$ is transformed into $\mathsf{Axis}^{n,m}$ via the natural isomorphism $\mathbb{R}^{m}\otimes\mathbb{R}^{n}\cong\mathbb{R}^{n}\otimes\mathbb{R}^{m}$.
Taking a product of an $m$-dimensional path with a $1$-dimensional path results in (scaled) path signature tensors. This immediately implies the following:
Lemma 4.5.
$\mathcal{P}_{d,m,1}=\mathcal{P}_{d,m}^{\mathsf{path}}$ where $\mathcal{P}^{\mathsf{path}}_{d,m}$ denotes (the affine cone of) the signature matrix variety considered in [2, Section 3.1].
Proof.
This follows from Corollary 3.5 as we can always identify a map $A:\mathbb{R}\otimes\mathbb{R}^{m}\to\mathbb{R}^{d}$ with a map $A^{\prime}:\mathbb{R}^{m}\to\mathbb{R}^{d}$ such that $A=\mathsf{id}\otimes A^{\prime}$ and thus $A\circ\mathsf{Axis}^{m,1}=\left(A^{\prime}\circ\mathsf{Axis}^{m}\right)\boxtimes\mathsf{Axis}^{1}$. ∎
Corollary 4.6.
For each $d$ there exists an $M\in\mathbb{N}$ such that
$\mathcal{P}_{d,m,1}=\mathcal{P}_{d,1,m}=\mathcal{U}^{\mathsf{path}}_{d,2}$
for all $m\geq M$, and where $\mathcal{U}^{\mathsf{path}}_{d,2}$ is the universal path variety at level $2$.
Proof.
This is now immediate from Lemma 4.5 and [2, Theorem 5.6]. ∎
This gives a lower bound on the dimension of $\mathcal{U}_{d}$ because of the grid structure (8); in the next section we show that $\mathcal{U}_{d}$ is actually the whole space.
4.1 The case $m,n\gg d$
Theorem 4.7.
The limit variety $\mathcal{U}_{d}$ attains the full ambient dimension, that is, $\mathcal{U}_{d}=\mathbb{C}^{d^{2}}$.
Proof.
We claim that already $\mathcal{P}_{d,d,2}=\mathbb{C}^{d^{2}}$. For this, write $A=\sigma^{(2)}(\mathsf{Axis}^{d})$ and note that by Lemma 2.6, Proposition 3.3 and Example 3.17, $\mathcal{P}_{d,d,2}$ is the Zariski closure of the image of the map $\phi:\mathbb{C}^{d\times d}\times\mathbb{C}^{d\times d}\to\mathbb{C}^{d\times d}$ sending $(S,T)$ to
$\begin{pmatrix}S^{\top}&T^{\top}\end{pmatrix}\begin{pmatrix}A&A\\ -A&0\end{pmatrix}\begin{pmatrix}S\\ T\end{pmatrix}=S^{\top}AS+S^{\top}AT-T^{\top}A.$
Here we used that $\begin{pmatrix}1&1\\ -1&0\end{pmatrix}$ is congruent to $\sigma^{(2)}(\mathsf{Axis}^{2})$. The Jacobian of $\phi$ at $\begin{pmatrix}I_{d}\\ I_{d}\end{pmatrix}$ is the linear map
$\mathbb{C}^{d\times d}\times\mathbb{C}^{d\times d}$ $\to\mathbb{C}^{d\times d},$
$(S,T)$ $\mapsto 2S^{\top}A+AT-T^{\top}A.$
As $A$ is invertible ($\det(A)=\frac{1}{2^{d}}$ by (11)), we see that the kernel of this map has dimension $d^{2}$ since, if $(S,T)$ lies in the kernel, we can solve for $S$. Thus the Jacobian has rank $d^{2}$ at this point, proving that $\phi$ is smooth in a neighbourhood $U$ of $\begin{pmatrix}I_{d}\\ I_{d}\end{pmatrix}$, [27, Definition 6.14, Proposition 6.15], and thus $\phi|_{U}$ is open by [43, Tag 056G]. This proves the claim since $\mathbb{C}^{d\times d}$ is irreducible. ∎
Remark 4.8.
The same proof idea yields a more general bound: for $e\leq d$, we have $\dim(\mathcal{P}_{d,e,2})\geq 2de-e^{2}$. To see this, one only needs to adapt the matrix formats in the proof above and replace $\begin{pmatrix}I_{d}\\ I_{d}\end{pmatrix}$ by the matrix $\begin{pmatrix}X\\ X\end{pmatrix}$ where $X=\begin{pmatrix}I_{e}&0\end{pmatrix}$. Unfortunately, this bound is not tight in any of our examples, cf. Figure 5 below.
4.2 The case $mn\leq d$
Theorem 4.9.
For $mn\leq d$ the dimension of $\mathcal{P}_{d,m,n}$ is
  • $dmn-\frac{1}{2}m^{2}n^{2}+m^{2}(n-1)+(m-1)n^{2}-\frac{7}{2}mn+4(m+n)-4$ for even $m,n$.
  • $dmn-\frac{1}{2}m^{2}n^{2}+m^{2}(n-1)+(m-1)n^{2}-\frac{3}{2}mn+m+n$ for even $m$, odd $n$.
  • $dmn-\frac{1}{2}m^{2}n^{2}+m^{2}(n-1)+(m-1)n^{2}-\frac{7}{2}mn+3(m+n)-2$ for odd $m,n$.
We omit the case $m$ odd, $n$ even due to symmetry.
Proof.
The variety $\mathcal{P}_{d,m,n}$ is the Zariski closure of the image of the map $\phi:\mathbb{C}^{d\times mn}\to\mathbb{C}^{d\times d},\ S\mapsto SA_{m,n}S^{\top}$ where $A_{m,n}=\sigma^{(2)}(\text{Axis}^{m,n})$. Equivalently, it is the Zariski closure of the $\operatorname{GL}_{d}$-congruence orbit of the $d\times d$-matrix $A^{\prime}_{m,n}$ which is obtained from $A_{m,n}$ by extending with zeroes. This congruence orbit is a differentiable manifold and a semi-algebraic set, in particular the dimension of the manifold agrees with the dimension of its Zariski-closure. The dimension of the manifold agrees with the dimension of the tangent space at $\phi(I_{d})$ (as it is a group orbit). Its codimension is the dimension of the kernel of the Jacobian $J(\phi)$ at $I_{d}$, in other words, the dimension of the solution space to the matrix equation
$SA^{\prime}_{m,n}+A^{\prime}_{m,n}S^{\top}=0,$ (24)
cf. [45, Lemma 2]. This dimension remains unchanged when $A^{\prime}_{m,n}$ is replaced by $A_{m,n}$. Thus in the following, we may assume $d=mn$.
[…]
As $A_{m,n}$ is indeed non-singular (by Example 3.17 we have $\det(A_{m,n})=\frac{1}{4^{mn}}$), our strategy is thus to determine the Jordan canonical form of $A_{m,n}^{-\top}A_{m,n}$. This is simplified significantly by the following observation: if we denote $A_{u}:=\sigma^{(2)}(\mathsf{Axis}^{u})$ then by Proposition 3.3 we have $A_{m,n}=A_{m}\otimes A_{n}$ where $\otimes$ denotes the Kronecker product. In particular
$A_{m,n}^{-\top}A_{m,n}=(A_{m}\otimes A_{n})^{-\top}(A_{m}\otimes A_{n})=A_{m}^{-\top}A_{m}\otimes A_{n}^{-\top}A_{n}.$
Moreover, as observed in the proof of [2, Theorem 3.4], the matrix $A_{u}$ is congruent to $N_{u}:=\begin{pmatrix}1&1\\ -1&0\end{pmatrix}\oplus\tau^{\frac{u-2}{2}}$ if $u$ is even, and congruent to $N_{u}:=\begin{pmatrix}1\end{pmatrix}\oplus\tau^{\frac{u-1}{2}}$ if $u$ is odd, where $\tau=\begin{pmatrix}1&-1\\ -1&0\end{pmatrix}$. Thus, we may replace $A_{m,n}$ by $N_{m}\otimes N_{n}$, $A_{m}$ by $N_{m}$ and $A_{n}$ by $N_{n}$. These observations reduce the problem to the following two questions:
  1. 1.
    What are the Jordan canonical forms of $N_{u}^{-\top}N_{u}$?
  2. 2.
    Given the Jordan canonical forms of two matrices, what is the Jordan canonical form of their Kronecker product?
For the first question, note that $\tau^{-\top}\tau=\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix}=J_{1}(-1)$ and
$\begin{pmatrix}1&1\\ -1&0\end{pmatrix}^{-\top}\begin{pmatrix}1&1\\ -1&0\end{pmatrix}=\begin{pmatrix}-1&0\\ -2&-1\end{pmatrix}$
whose Jordan canonical form is $\begin{pmatrix}-1&1\\ 0&-1\end{pmatrix}=J_{2}(-1)$. Thus, if $u$ is odd then the canonical form of $N_{u}^{-\top}N_{u}$ is $J_{1}(1)\oplus J_{1}(-1)^{\oplus u-1}$, and if $u$ is even then the canonical form of $N_{u}^{-\top}N_{u}$ is $J_{2}(-1)\oplus J_{1}(-1)^{\oplus u-2}$.
The answer to the second question is [7, Theorem 4.6] from which we obtain: the Jordan canonical form of $A_{m,n}^{-T}A_{m,n}$ is
  • $J_{3}(1)\oplus J_{1}(1)\oplus J_{2}(1)^{\oplus m-2+n-2}\oplus J_{1}(1)^{\oplus(m-2)(n-2)}$ if $m,n$ are even,
  • $J_{2}(-1)\oplus J_{2}(1)^{\oplus n-1}\oplus J_{1}(-1)^{\oplus m-2}\oplus J_{1}(1)^{\oplus(m-2)(n-1)}$ if $m$ is even, $n$ is odd,
  • $J_{1}(1)\oplus J_{1}(-1)^{\oplus m-1+n-1}\oplus J_{1}(1)^{\oplus(m-1)(n-1)}$ if $m,n$ are odd.
Thus the canonical form for congruence of $A_{m,n}$ is
  • $\Gamma_{3}\oplus H_{4}(1)^{\oplus\frac{m+n-4}{2}}\oplus\Gamma_{1}^{\oplus(m-2)(n-2)+1}$ if $m,n$ are even,
  • $\Gamma_{2}\oplus H_{4}(1)^{\oplus\frac{n-1}{2}}\oplus H_{2}(-1)^{\oplus\frac{m-2}{2}}\oplus\Gamma_{1}^{\oplus(m-2)(n-1)}$ if $m$ is even, $n$ is odd,
  • $H_{2}(-1)^{\oplus\frac{m+n-2}{2}}\oplus\Gamma_{1}^{\oplus(m-1)(n-1)+1}$ if $m,n$ are odd.
Finally, we conclude by calculating the codimensions using the recipe given in [45, Theorem 2]. ∎
$\left[\begin{array}[]{cccc}4&7&9&10\\ 7&14&\mathbf{16}&16\\ 9&\mathbf{16}&\mathbf{16}&16\\ 10&16&16&16\end{array}\right]$
$d=4$
$\left[\small\begin{array}[]{ccccc}5&9&12&14&15\\ 9&18&\mathbf{24}&\mathbf{25}&25\\ 12&\mathbf{24}&\mathbf{25}&\mathbf{25}&25\\ 14&\mathbf{25}&\mathbf{25}&\mathbf{25}&25\\ 15&25&25&25&25\end{array}\right]$
$d=5$
$\left[\small\begin{array}[]{cccccc}6&11&15&18&20&21\\ 11&22&31&\mathbf{34}&\mathbf{36}&36\\ 15&31&\mathbf{34}&\mathbf{36}&\mathbf{36}&36\\ 18&\mathbf{34}&\mathbf{36}&\mathbf{36}&\mathbf{36}&36\\ 20&\mathbf{36}&\mathbf{36}&\mathbf{36}&\mathbf{36}&36\\ 21&36&36&36&36&36\end{array}\right]$
$d=6$
$\left[\small\begin{array}[]{ccccccc}7&13&18&22&25&27&28\\ 13&26&37&\mathbf{43}&\mathbf{48}&\mathbf{49}&49\\ 18&37&\mathbf{43}&\mathbf{49}&\mathbf{49}&\mathbf{49}&49\\ 22&\mathbf{43}&\mathbf{49}&\mathbf{49}&\mathbf{49}&\mathbf{49}&49\\ 25&\mathbf{48}&\mathbf{49}&\mathbf{49}&\mathbf{49}&\mathbf{49}&49\\ 27&\mathbf{49}&\mathbf{49}&\mathbf{49}&\mathbf{49}&\mathbf{49}&49\\ 28&49&49&49&49&49&49\end{array}\right]$
$d=7$
$\left[\small\begin{array}[]{cccccccc}8&15&21&26&30&33&35&36\\ 15&30&43&52&\mathbf{60}&\mathbf{62}&\mathbf{64}&64\\ 21&43&\mathbf{52}&\mathbf{63}&\mathbf{63}&\mathbf{64}&\mathbf{64}&64\\ 26&52&\mathbf{63}&\mathbf{63}&\mathbf{64}&\mathbf{64}&\mathbf{64}&64\\ 30&\mathbf{60}&\mathbf{63}&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&64\\ 33&\mathbf{62}&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&64\\ 35&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&64\\ 36&64&64&64&64&64&64&64\\ \end{array}\right]$
$d=8$
Figure 5: The dimensions $\dim(\mathcal{P}_{d,m,n})$ stored in $d\times d$ matrices for varying $d$, where bold numbers are not covered by Theorem 4.7 or Theorem 4.9. We compute these remaining dimensions according to Remark 4.13. Remarkable is the fact that in all examples the full dimension $d^{2}$ is already attained on the antidiagonal $m+n=d+1$ for $m,n\geq 2$, in particular $\mathcal{P}_{d,d-1,2}$ has full dimension, which is not clear from Remark 4.8.
Remark 4.10.
In light of Lemma 4.5, one checks that for $n=1$ the formulas simplify to $md-\binom{m}{2}$, which is indeed the dimension of the affine cone of the path signature matrix varieties considered in [2], cf. Theorem 3.4 in loc. cit..
When rescaling the rows of $\sigma(\mathsf{Mom}^{m})$ we obtain a Cauchy matrix, hence with Corollary 3.6, $\sigma(\mathsf{Mom}^{m,n})$ has full rank for any $m$ and $n$. However, the rank of the symmetric and skew-symmetric part behaves differently as in the path case:
Corollary 4.11.
For $A_{m,n}:=\sigma^{(2)}(\mathsf{Mom}^{m,n})$ let $A_{m,n}^{\mathsf{sym}}$ and $A_{m,n}^{\mathsf{sk}}$ denote its symmetric and skew-symmetric part, respectively. Then
$\operatorname{rk}(A_{m,n}^{\mathsf{sym}})=\begin{cases}m\cdot n&m,n\text{ even}\\ 1+m(n-1)&m\text{ even},n\text{ odd}\\ 1+(m-1)(n-1)&m,n\text{ odd}\end{cases}$
(again we omit the case $m$ odd, $n$ even due to symmetry) and
$\operatorname{rk}(A_{m,n}^{\mathsf{sk}})=\begin{cases}m+n-1&m+n\text{ odd}\\ m+n-2&m+n\text{ even.}\end{cases}$
Proof.
In the proof of Theorem 4.9, add up the ranks of the symmetric/skew-symmetric parts of the blocks in the canonical form for congruence of $A_{m,n}$. ∎
For $m=1$ or $n=1$ we obtain a rank one symmetric part according to the path framework, [2, eq. (4)].
As the rank of symmetric and skew-symmetric part is preserved under congruence, Corollary 4.11 induces constraints on the congruence orbits. However not all relations on the orbit arise in this way:
Example 4.12.
Consider the case $d=4,m,n=2$. According to Corollary 4.11 the symmetric part of $A:=\sigma^{(2)}(\mathsf{Axis}^{m,n})$ has full rank and the skew-symmetric part has rank $2$. In particular, the Pfaffian $x_{12}x_{34}-x_{13}x_{24}+x_{14}x_{23}$ vanishes on the skew-symmetric part of matrices in the congruence orbit. However, according to Theorem 4.9 the dimension of $\mathcal{P}_{4,2,2}$ is 14, so this can not be the only relation. Indeed, another relation is given by $\det(A)\det(X^{\mathsf{sym}})=\det(X)\det(A^{\mathsf{sym}})$: the ratio $\det(X)/\det(X^{\mathsf{sym}})$ is well-defined and clearly constant on the congruence orbit. Using Macaulay 2 one checks that the variety $\mathcal{P}_{4,2,2}$ is a complete intersection, cut out by those two equations.
Remark 4.13.
Some choices of $m,n,d$ are neither covered by Theorem 4.7 nor by Theorem 4.9. Using Macaulay 2 with NumericalImplicitization, we computed the missing dimensions of $\mathcal{P}_{d,m,n}$ for $m,n\leq d$. We summarize our findings in Figure 5.
Example 4.14.
According to Figure 5, $\mathcal{P}_{5,3,2}$ is a hypersurface in $\mathbb{R}^{5\times 5}$. By numerical computations in Macaulay 2 using the packages NumericalImplicitization and MultigradedImplicitization one observes that this hypersurface is cut out by a single quintic $f$ with $999$ terms. One checks that $f\in\mathsf{minors}(3,X)\cap\mathsf{minors}(4,X^{\mathsf{sk}})\cap\mathsf{minors}(1,X^{\mathsf{sym}})$, where $\mathsf{minors}(i,X)$ denotes the ideal of $i$-minors of $X$.
5 Conclusions and future work
Our notion of signature was introduced by [26, 14] under the name of $\mathsf{id}$-signature, see Remark 2.3 for a detailed discussion. Both references observe that this $\mathsf{id}$-signature is not universal. We prove that in general the signature matrix of a membrane does not satisfy any algebraic relations, so in particular, no shuffle relations. This proof relies on matrix theory and is not generalizable for tensors in an obvious way. We conjecture that our result is true for all signature tensors.
Conjecture 5.1.
Theorem 4.7 holds for arbitrary level $k\geq 3$.
Furthermore, we propose novel closed formulas for signatures of piecewise bilinear and polynomial membranes. These membranes can be used for interpolating arbitrary image data, a necessary step in the iterated-integrals signature method. This should be useful for [48] which, instead of any interpolation, approximates two-parameter integrals via certain two-parameter sums.
Inspired by linear complexity of the two-parameter sums signature on diagonal matrix compositions [18, Theorem 4.5], we would like to improve our computational complexity for interpolated image data, Theorem 3.19.
Question 5.2.
Is there an algorithm with linear time complexity in $m$ and $n$ to compute a level $k$ signature entry of a piecewise bilinear membranes of order $(m,n)$?
Already in the path framework, a signature matrix is not enough to recover the path, see [2, Example 6.8]. In particular, we can not recover a membrane from its signature matrix in general. An important property of the ‘right’ notion of signature is the ability of recovering membranes from all of its signature tensors, as it is done in [8] for paths. It is an open question whether our signature tensors have this property.
Question 5.3.
Can we always recover membranes from their signature tensors?
One can also ask the following more special question:
Question 5.4.
Can piecewise bilinear membranes of low order already be recovered from a finite set of signature tensors?
Again, this is true for paths, see [41].
We also remark that recovery with the two-parameter sums signature is always possible. For this compare [18, Theorem 2.20] and its proof.
Finally, we would like to understand the missing dimensions for membranes of very low order. Computations as shown in Figure 5 inspire the following conjecture:
Conjecture 5.5.
For $m,n\geq 2$, $\mathcal{P}_{d,m,n}$ has full dimension whenever $m+n>d$.
Note that these cases are not covered by Remark 4.8.
Question 5.6.
What are the dimensions of $\mathcal{P}_{d,n,m}$ for $m+n\leq d<mn$?
Acknowledgments
We thank Joscha Diehl for detailed discussions on the various notions of two-parameter signatures, Sven Groß for introducing us to interpolation literature, Ben Hollering for providing Example 4.14, Rosa Preiß for pointing out [14, Example 4.2], and Bernd Sturmfels for his help with Example 4.12. The authors acknowledge seed support for DFG CRC/TRR 388 “Rough Analysis, Stochastic Dynamics and Related Topics”.

### v2

@@
 Abstract
-We prove that, unlike in the case of paths, the signature matrix of a membrane does not satisfy any algebraic relations.
-We derive novel closed-form expressions for the signatures of polynomial membranes and piecewise bilinear interpolations for arbitrary $2$-parameter data in $d$-dimensional space.
-We show that these two families of membranes admit the same set of signature matrices and scrutinize the corresponding affine variety.
+The signature of a membrane is a sequence of tensors whose entries are iterated integrals.
+We study algebraic properties of membrane signatures, with a focus on signature matrices of polynomial and piecewise bilinear membranes.
+Generalizing known results for path signatures, we show that the two families of membranes admit the same set of signature matrices and we examine the corresponding affine varieties.
+In particular, we prove that there are no algebraic relations on signature matrices of membranes, in contrast to the path case.
+We complement our results by a linear time algorithm for the computation of signature tensors for piecewise bilinear membranes.
 Keywords: two-parameter signatures, affine algebraic varieties, matrix congruence, geometry of paths and membranes, iterated integrals
@@
 1 Introduction
-The signature of a path.
 A path is a continuous map $X:[0,1]\to\mathbb{R}^{d}$ with piecewise continuously differentiable coordinate functions.
-Its $k$-th level signature $\sigma^{(k)}(X)$, as introduced by Chen in the seminal work [8], is a $k$-tensor of format $d\times\dots\times d$, defined by forming certain iterated integrals.
-In particular, its second level signature can be viewed as a signature matrix.
-It is defined as follows:
-$\sigma^{(2)}(X)_{i,j}$ $:=\int_{0}^{1}\int_{0}^{t_{2}}\partial X_{i}(t_{1})\,\partial X_{j}(t_{2})\;\mathrm{d}t_{1}\mathrm{d}t_{2}$ (1)
-These matrices were studied in [2]. They satisfy some interesting properties: for example, their symmetric part has always rank $1$, due to the shuffle relations satisfied by path signatures.
-Moreover, the set of signature matrices of polynomial paths with degree $\leq m$ and the set of signature matrices of piecewise linear paths with at most $m$ segments agree.
-The equations that vanish on these sets, that is, the varieties of signature matrices, are well-understood: they can be described as certain determinantal varieties [2, Theorem 3.4].
-The signature of a membrane.
-A membrane is a continuous map $X:[0,1]^{2}\rightarrow{\mathbb{R}}^{d}$ where the coordinate functions $X_{1},X_{2},\dots,X_{d}$ are (piecewise) continuously differentiable functions in the two arguments.
+The signature of a path was introduced by Chen in his seminal work [Chen1954IteratedIA] and has since become a central construction in the theory of rough paths.
+The second level signature $S(X)\in\mathbb{R}^{d\times d}$ of a path $X$ is a matrix whose entries are iterated integrals
+$S(X)_{i,j}$ $:=\int_{0}^{1}\int_{0}^{t_{2}}\dot{X}_{i}(t_{1})\,\dot{X}_{j}(t_{2})\;\mathrm{d}t_{1}\mathrm{d}t_{2}$
+corresponding to $1\leq i,j\leq d$, and where $\dot{X}$ denotes the derivative of $X$.
+These signature matrices were studied in [bib:AFS2019, Section 3] and satisfy several interesting properties: their symmetric parts always have rank $1$, due to the shuffle relations satisfied by path signatures.
+The sets of signature matrices of polynomial and piecewise linear paths coincide.
+This set is semi-algebraic and its Zariski closure can be described as a certain determinantal variety.
+The signature of a membrane
+Recently, the signature has been extended to membranes.
+A $d$-dimensional membrane is a continuous map $X:[0,1]^{2}\rightarrow{\mathbb{R}}^{d}$ where the coordinate functions $X_{1},X_{2},\dots,X_{d}$ are (piecewise) continuously differentiable functions in the two arguments.
 In particular the differentials $\partial_{12}X_{j}(s,t):=\frac{\partial^{2}}{\partial s\partial t}X_{j}(s,t)$ obey the usual rules of calculus.
 Generalizing the path setting, one can associate a collection of tensors to $X$ via iterated integrals (see Definition 2.2).
-In particular, the signature matrix of a membrane $X$ is defined as an $\mathbb{R}$-valued $d\times d$-matrix with entries
-$\sigma^{(2)}(X)_{i,j}$ $:=\int_{0}^{1}\int_{0}^{1}\int_{0}^{t_{2}}\int_{0}^{s_{2}}\partial_{12}X_{i}(s_{1},t_{1})\,\partial_{12}X_{j}(s_{2},t_{2})\;\mathrm{d}s_{1}\mathrm{d}t_{1}\mathrm{d}s_{2}\mathrm{d}t_{2}$ (2)
-for every $1\leq i,j\leq d$. This matrix has recently been introduced as the second level of the two-parameter $\mathsf{id}$-signature [26, 48, 14].
-While it is easy to see that, contrary to path signatures, this signature does not satisfy shuffle relations, little is known about the algebraic relations satisfied by signature tensors in general.
-In this paper, inspired by [2], we address this question for the case of matrices.
-We write $\sigma(X):=\sigma^{(2)}(X)$ in the following.
-We focus on polynomial membranes of order $(m,n)$, that is, membranes $X$ whose component functions are polynomials of multidegree $\leq(m,n)$, and piecewise bilinear membranes of order $(m,n)$, see Definition 3.14.
-Their signature matrices (2) form semialgebraic subsets of $\mathbb{R}^{d^{2}}$.
-To study the polynomials that vanish on these sets, we move to the algebraically closed field $\mathbb{C}$ and consider the Zariski closures
-$\mathcal{P}_{d,m,n}^{\mathsf{pol}}\quad\text{ and }\quad\mathcal{P}_{d,m,n}^{\mathsf{bil}}.$ (3)
-Main question.
-What can be said about the varieties of signature matrices of polynomial membranes and piecewise bilinear membranes?
+The associated object was termed $\mathsf{id}$-signature in [diehl2024signature, Definition 4.1], [GLNO2022, Section 4.7].
+We call the second level of the $\mathsf{id}$-signature the signature matrix of a membrane.
+It is defined as the $\mathbb{R}$-valued $d\times d$ matrix $S(X)$ with entries
+$S(X)_{i,j}$ $:=\int_{0}^{1}\int_{0}^{1}\int_{0}^{t_{2}}\int_{0}^{s_{2}}\partial_{12}X_{i}(s_{1},t_{1})\,\partial_{12}X_{j}(s_{2},t_{2})\;\mathrm{d}s_{1}\mathrm{d}t_{1}\mathrm{d}s_{2}\mathrm{d}t_{2}$ (1)
+for every $1\leq i,j\leq d$.
+Although it is easy to see that the $\mathsf{id}$-signature does not satisfy the same algebraic relations as the path signature [diehl2024signature, Section 4.3], little is known about its algebraic relations in general.
+In this paper, inspired by [bib:AFS2019], we address this question for the case of signature matrices.
+For this we consider the family of membranes whose component functions are polynomials of order $(m,n)$.
+Their signature matrices (1) form a semialgebraic subset of $\mathbb{R}^{d\times d}$.
+To study the polynomials that vanish on these sets, we move to the algebraically closed field $\mathbb{C}$ and consider the associated Zariski closure, in the following denoted by $\mathcal{M}_{d,m,n}$.
 Figure 1: A polynomial membrane in $\mathbb{R}^{2}$ and a piecewise bilinear membrane in $\mathbb{R}^{3}$.
@@
 See Example 3.7 and Figure 3 for a more detailed description.
-An implicitization problem.
-Our strategy is to view the varieties (3) as solutions to an implicitization problem.
-To this end, we show that both classes of membranes admit a dictionary membrane (see [41]), i.e.
-a membrane which yields all other membranes in the class via linear transforms, up to certain terms that vanish under the signature map.
-Our first candidate of a dictionary is the $mn$-dimensional moment membrane,
+An implicitization problem
+Our strategy is to view the variety $\mathcal{M}_{d,m,n}$ as the Zariski closure of a certain matrix congruence orbit.
+For this we use a dictionary membrane inspired by [bib:PSS2019], i.e., a special membrane that yields all polynomial membranes via a suitable linear transform, up to terms that vanish under the signature map.
+The moment membrane,
 $\mathsf{Mom}^{m,n}:[0,1]^{2}$ $\to\mathbb{R}^{mn},(s,t)\mapsto(s,s^{2},\dots,s^{m})\otimes(t,t^{2},\dots,t^{n}),$ (2)
-where $\otimes$ is the Kronecker product. If $m=n=2$, then $\mathsf{Mom}^{2,2}(s,t)=(st,st^{2},s^{2}t,s^{2}t^{2})$ has the signature matrix
-$\sigma(\mathsf{Mom}^{2,2})=\begin{pmatrix}\frac{1}{4}&\frac{1}{3}&\frac{1}{3}&\frac{4}{9}\\[4.0pt] \frac{1}{6}&\frac{1}{4}&\frac{2}{9}&\frac{1}{3}\\[4.0pt] \frac{1}{6}&\frac{2}{9}&\frac{1}{4}&\frac{1}{3}\\[4.0pt] \frac{1}{9}&\frac{1}{6}&\frac{1}{6}&\frac{1}{4}\end{pmatrix}=\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}\otimes\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}$ (5)
-with a Kronecker factorization into path signatures (Corollary 3.6).
-Note that indeed every $d$-dimensional polynomial membrane $X$ with terms of multidegree $\geq(1,1)$ is of the form $X=A\,\mathsf{Mom}^{m,n}$.
+is a dictionary for polynomial membranes of order $(m,n)$, where $\otimes$ denotes the Kronecker product.
+We call this membrane a product membrane since it factorizes into paths.
+We will see in Section 3 that the signature matrix of a product membrane is a Kronecker product of the signature matrices of its path factors.
+If for example $m=n=2$, then $\mathsf{Mom}^{2,2}(s,t)=(st,st^{2},s^{2}t,s^{2}t^{2})$ has the signature matrix
+$S(\mathsf{Mom}^{2,2})=\begin{pmatrix}\frac{1}{4}&\frac{1}{3}&\frac{1}{3}&\frac{4}{9}\\[4.0pt] \frac{1}{6}&\frac{1}{4}&\frac{2}{9}&\frac{1}{3}\\[4.0pt] \frac{1}{6}&\frac{2}{9}&\frac{1}{4}&\frac{1}{3}\\[4.0pt] \frac{1}{9}&\frac{1}{6}&\frac{1}{6}&\frac{1}{4}\end{pmatrix}=\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}\otimes\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}.$ (3)
+Since the moment membrane is a dictionary, we can express every polynomial membrane $X$ without linear terms of order $(m,n)$ as a linear transform $X=A\,\mathsf{Mom}^{m,n}$, where the matrix $A$ is chosen accordingly.
+Note that the membrane signature is invariant under addition of linear terms.
 With equivariance of the signature (Lemma 2.6) we obtain the closed formula,
-$\sigma(X)=A\sigma(\mathsf{Mom}^{m,n})A^{\top},$ (6)
-as a matrix congruence. Hence, continuing with our example (5) and $d=2$, we have
+$S(X)=AS(\mathsf{Mom}^{m,n})A^{\top},$ (4)
+through a matrix congruence transform. Continuing with our example (3) and $d=2$, we have
 $X=A\,\mathsf{Mom}^{2,2}=\begin{pmatrix}a_{1,1}st+a_{1,2}st^{2}+a_{1,3}s^{2}t+a_{1,4}s^{2}t^{2}\\ a_{2,1}st+a_{2,2}st^{2}+a_{2,3}s^{2}t+a_{2,4}s^{2}t^{2}\end{pmatrix}$ (5)
-Thus, we obtain the first entry of our signature matrix from (5) and (6): It is the homogeneous polynomial
-$\sigma(X)_{1,1}=\frac{1}{4}a_{1,1}^{2}$ $+\frac{1}{2}a_{1,1}a_{1,2}+\frac{1}{2}a_{1,1}a_{1,3}+\frac{5}{9}a_{1,1}a_{1,4}+\frac{1}{4}a_{1,2}^{2}$
+and thus with (3) and (4) the homogeneous polynomial
+$S(X)_{1,1}=\frac{1}{4}a_{1,1}^{2}$ $+\frac{1}{2}a_{1,1}a_{1,2}+\frac{1}{2}a_{1,1}a_{1,3}+\frac{5}{9}a_{1,1}a_{1,4}+\frac{1}{4}a_{1,2}^{2}$
 $+\frac{4}{9}a_{1,2}a_{1,3}+\frac{1}{2}a_{1,2}a_{1,4}+\frac{1}{4}a_{1,3}^{2}+\frac{1}{2}a_{1,3}a_{1,4}+\frac{1}{4}a_{1,4}^{2}$
-of degree two. The remaining three polynomial entries of $\sigma(X)$ are provided in Example 3.7. We illustrate the polynomial membrane (7) for explicit $A$ in Figure 3.
-The axis membrane.
-Interpolation is a necessary step to fit discrete data into the framework of iterated integrals.
-Our second important dictionary, the axis membrane $\mathsf{Axis}^{m,n}$, is inspired by the dictionary for piecewise linear paths [2, 41] (see Definition 3.15).
-This $mn$-dimensional membrane allows us to interpolate arbitrary $d$-dimensional discrete data on an $m\times n$ grid via a piecewise bilinear membrane.
-For an illustration of such an interpolation see $Y$ from Figure 3.
-We provide novel closed-form expressions and computational complexity statements for its signatures (Theorem 3.19).
-Main results.
-Using our description via dictionaries, we obtain an analogue of [2, Theorem 3.3] for membranes:
-Theorem (4.1).
-For all $d,m$ and $n$,
-$\mathcal{P}^{\mathsf{pol}}_{d,m,n}=\mathcal{P}^{\mathsf{bil}}_{d,m,n}=:\mathcal{P}_{d,m,n}.$
-In fact, already the defining semi-algebraic sets agree.
-These varieties are irreducible (Proposition 4.3) and we obtain a grid
+as the first entry of our signature matrix.
+The remaining $3$ polynomial entries of $S(X)$ are provided in Example 3.7.
+We illustrate the polynomial membrane (5) for an explicit matrix $A$ in Figure 3.
+Main results
+The varieties of polynomial signature matrices $\mathcal{M}_{d,m,n}$ can be arranged in the following grid of inclusions,
 […]
@@
 […]
-$\mathcal{P}_{d,1,n}\subseteq\dots\subseteq\mathcal{P}_{d,M-1,n}\subseteq\mathcal{P}_{d,M,n}=\mathcal{P}_{d,M+1,n}$ =$\dots$= …
+$\mathcal{M}_{d,1,n}\subseteq\dots\subseteq\mathcal{M}_{d,M-1,n}\subseteq\mathcal{M}_{d,M,n}=\mathcal{M}_{d,M+1,n}$ =$\dots$= …
 for all $m,n\in\mathbb{N}$.
-As $\mathcal{P}_{d,1,m}=\mathcal{P}_{d,m,1}$ agrees with the path varieties studied in [2] (Lemma 4.5), it follows that $\mathcal{P}_{d,1,N}=\mathcal{P}_{d,M,1}$ is the universal path variety.
-However, in strong contrast to the path case, in general there are no algebraic relations in our two-parameter signature matrices (2) at all:
+We show in Lemma 4.5 that $\mathcal{M}_{d,1,m}=\mathcal{M}_{d,m,1}$ agrees with the path varieties studied in [bib:AFS2019].
+In particular, it follows that $\mathcal{M}_{d,1,N}=\mathcal{M}_{d,M,1}$ is the variety parametrizing the signature matrices of all smooth paths, called universal path variety in [bib:AFS2019, Section 4.3].
+However, in sharp contrast to the path case, there are generally no algebraic relations in our two-parameter signature matrices (1) at all.
 Theorem (4.13).
-The limit variety $\mathcal{P}_{d,M,N}$ in (8) attains the full ambient dimension, i.e. $\mathcal{P}_{d,M,N}=\mathbb{C}^{d^{2}}$.
-The varieties become more interesting if $m$ and $n$ are small compared to $d$.
-Employing the theory of orbits under congruence [45, 30] we were able to find explicit formulas for the dimensions of $\mathcal{P}_{d,m,n}$ whenever $mn\leq d$, generalizing [2, Theorem 3.4].
+The variety $\mathcal{M}_{d,M,N}$ in (6) attains the full ambient dimension, i.e.
+$\mathcal{M}_{d,M,N}=\mathbb{C}^{d\times d}$.
+More precisely, $\mathcal{M}_{d,m,n}=\mathbb{C}^{d\times d}$ whenever $m+n>d$ with $m,n\not=1$.
+The proof relies on the theory of matrix congruence orbits [TERAN201144, HORN20061010]. Using this machinery, we can find lower and upper bounds for our varieties $\mathcal{M}_{d,m,n}$:
+Theorem (4.9).
+Let $\mathcal{S}_{a,b}$ be the variety of $d\times d$ matrices whose symmetric part has rank $\leq a$ and whose skew-symmetric part has rank $\leq b$. Then
+  • $\mathcal{S}_{(m-2)(n-2),m+n-2}\subseteq\mathcal{M}_{d,m,n}\subseteq\mathcal{S}_{mn,m+n-2}$ if $m,n$ are even,
+  • $\mathcal{S}_{2,n-1}\subseteq\mathcal{M}_{d,2,n}\subseteq\mathcal{S}_{2(n-1)+1,n+1}$ if $n$ is odd,
+  • $\mathcal{S}_{(m-2)(n-1)-1,m+n-1}\subseteq\mathcal{M}_{d,m,n}\subseteq\mathcal{S}_{m(n-1)+1,m+n-1}$ if $m\geq 4$ is even, $n$ is odd,
+  • $\mathcal{S}_{(m-1)(n-1)+1,m+n-2}=\mathcal{M}_{d,m,n}$ if $m,n$ are odd.
+In particular, it follows that our varieties are only interesting if $m$ and $n$ are small compared to $d$.
+From [TERAN201144] we obtain explicit formulas for the dimensions of $\mathcal{M}_{d,m,n}$ whenever $mn\leq d$, generalising [bib:AFS2019, Theorem 3.4].
 Note that there are $dmn\leq d^{2}$ parameters involved, so the dimension can be at most $dmn$.
 Theorem (4.7).
-For $mn\leq d$ the dimension of $\mathcal{P}_{d,m,n}$ is
-  • $dmn-\frac{1}{2}m^{2}n^{2}+m^{2}(n-1)+(m-1)n^{2}-\frac{7}{2}mn+4(m+n)-4$ for even $m,n$.
-  • $dmn-\frac{1}{2}m^{2}n^{2}+m^{2}(n-1)+(m-1)n^{2}-\frac{3}{2}mn+m+n$ for even $m$, odd $n$.
-  • $dmn-\frac{1}{2}m^{2}n^{2}+m^{2}(n-1)+(m-1)n^{2}-\frac{7}{2}mn+3(m+n)-2$ for odd $m,n$.
-These results are complemented by some explicit computations using the open-source computer algebra systems Macaulay 2 [28, 19], and OSCAR [40, 13], see e.g.
+For $mn\leq d$, the dimension of $\mathcal{M}_{d,m,n}$ is
+  • $dmn-\frac{1}{2}m^{2}n^{2}+m^{2}(n-1)+(m-1)n^{2}-\frac{7}{2}mn+4(m+n)-4$, if $m$ and $n$ even,
+  • $dmn-\frac{1}{2}m^{2}n^{2}+m^{2}(n-1)+(m-1)n^{2}-\frac{3}{2}mn+m+n$, if $m$ even and $n$ odd,
+  • $dmn-\frac{1}{2}m^{2}n^{2}+m^{2}(n-1)+(m-1)n^{2}-\frac{7}{2}mn+3(m+n)-2$, if $m$ and $n$ odd.
+The results are complemented by some explicit computations using the open-source computer algebra system Macaulay 2 [GraysonStillmanMacaulay2, eisenbud2001computations], see e.g.
 Examples 4.12 and 4.2, Remark 4.14 and Example 4.15.
+Piecewise bilinear membrances
+Bilinear interpolation is an easy way to fit discrete data into the framework of iterated integrals.
+It allows us to interpolate arbitrary $d$-dimensional discrete data on an $m\times n$ grid via a piecewise bilinear membrane of order $(m,n)$; see Definition 3.13.
+For an illustration of such an interpolation see $Y$ from Figure 3.
+In Definition 3.14 we define our second important dictionary, the axis membrane $\mathsf{Axis}^{m,n}$.
+It is inspired by the dictionary for piecewise linear paths from [bib:PSS2019], and it allows an analogue of [bib:AFS2019, Theorem 3.3] for membranes.
+Theorem (4.1).
+The sets of signature matrices of polynomial and piecewise bilinear membranes with order $(m,n)$ coincide. In particular, both sets define the same algebraic variety $\mathcal{M}_{d,m,n}$.
+A naive computation of signature matrices via congruence operation results only in a quadratic time algorithm with respect to the number of interpolation points; see Corollary 3.11.
+In particular, there is no analogue of Chen’s identity for the id-signature [GLNO2022, diehl2024signature].
+However, using a cumulated approach inspired by [DEFT22, DS22] we can drop this cost to linear complexity.
+Corollary (3.19).
+The signature matrix of a piecewise bilinear membrane in $\mathbb{R}^{d}$ of order $(m,n)$ is computable in $\mathcal{O}(d^{2}mn)$.
+We can generalize this idea for the $k$-th level of the $\mathsf{id}$-signature tensor; see Theorem 3.18. This yields a linear two-parameter integration algorithm for the signatures of images.
+The code accompanying this paper is available at
+https://github.com/felixlotter/signature-matrices-of-membranes.
 Related work
-The collection of iterated integrals of a path has been used as a tool for classification in various settings since Chen’s seminal 1954 work [8].
-After pioneering work by Lyons and collaborators on the use of signatures as features for data streams in the early 2010s [29, 32], there has been a surge in the number of applications, which range from motion recognition [47], finance [4], deep learning [33, 34, 46, 35], oceanography [44], and lately data science for images [48, 31].
-We refer the interested reader to the surveys [10, 38].
-Since [2, 41, 24] established the initial connection between stochastic analysis and algebraic geometry, the signature has also been applied in various related fields, including the theory of tensor decompositions [3, 22, 25], clustering tasks [11, 39], tropical time series analysis [15, 16] and the geometry of paths [17, 1, 42].
-Recently, the signature has been extended to two parameters, generalizing from paths to membranes: the mapping space signature [26] provides a principled two-parameter extension of Chen’s iterated integrals, based on its original use in topology, whereas the two-parameter sums signature [18] characterizes equivalence up to warping in two directions.
-These two approaches have recently been combined in [14] using mixed partial derivatives.
-Due to the lack of a two-parameter version of Chen’s identity in all the above references, recent works [9, 37, 36] use higher category theory to find an appropriate generalization of a group that is rich enough to represent the horizontal and vertical concatenation of membranes.
+The collection of iterated integrals of a path has been used as a tool for classification in various settings since Chen’s seminal 1954 work [Chen1954IteratedIA].
+After pioneering work by Lyons and collaborators on the use of signatures as features for data streams in the early 2010s [Gyu13, handw13], there has been a surge in the number of applications.
+Since [bib:AFS2019, bib:PSS2019, GALUPPI2019282] established the initial connection between stochastic analysis and algebraic geometry, the signature has also been applied in various related fields, including the theory of tensor decompositions [AGOSS2025, FLS24, galuppi2024ranksymmetriessignaturetensors], clustering tasks [clausel2024barycenterfreenilpotentlie, mignot:hal-04392568], tropical time series analysis [DEFT22, diehl2023fruitsfeatureextractionusing] and the geometry of paths [Diehl2019, 10.1007/978-3-031-38271-0_45, preiß2024algebraicgeometrypathsiteratedintegrals].
+Recently, the signature has been extended to two parameters, generalizing from paths to membranes: the mapping space signature [GLNO2022] provides a principled two-parameter extension of Chen’s iterated integrals, based on its original use in topology, whereas the two-parameter sums signature [DS22] characterizes equivalence up to warping in two directions.
+These two approaches have recently been combined in [diehl2024signature] using mixed partial derivatives.
+Due to the lack of a two-parameter version of Chen’s identity in all the above references, recent works [CDEFT2024, LO2023, lee2024] use higher category theory to find an appropriate generalization of a group that is rich enough to represent the horizontal and vertical concatenation of membranes.
+Outline.
+In section Section 2 we recall the notion of iterated-integrals signatures for paths and membranes.
+Section 3 introduces the notion of a product membrane.
+For arbitrary $k$, the $k$-th level signature of these membranes can be computed from path signatures, which will allow us to use theory from the path setting.
+We define the moment membrane and the canonical axis membrane and include algorithmic considerations of bilinear interpolations and linearly computable signature entries in Section 3.2.
+Finally, we return to the matrix case $k=2$ in Section 4 where we prove our main results on signature matrix varieties.
+For open problems and future work see Section 5.
 2 Iterated integrals
-The iterated-integrals signature $\sigma=\sigma(X)$ of a $d$-dimensional path $X:[0,1]\to\mathbb{R}^{d}$ is an element of the tensor series algebra $T((\mathbb{R}^{d}))\cong\operatorname{Hom}(T(\mathbb{R}^{d}),\mathbb{R})$, defined by
-$\sigma:T(\mathbb{R}^{d})$ $\to\mathbb{R},$
-$e_{i_{1}}\otimes\dots\otimes e_{i_{k}}$ $\mapsto\int_{\Delta}\mathrm{d}X_{i_{1}}(t_{1})\dots\mathrm{d}X_{i_{k}}(t_{k}),$ (9)
-where $\Delta$ is the simplex $10<t_{1}<\dots<t_{k}<$ 1.
-It is common to write the tensor $e_{i_{1}}\otimes\dots\otimes e_{i_{k}}$ as a word $i_{1}\dots i_{k}$, identifying the tensor algebra with the free associative algebra over the letters 1 to d.
+Let $V:=\mathbb{R}^{d}$ and let $T(V^{*})$ denote the tensor algebra $\bigoplus_{k\in\mathbb{N}_{0}}(V^{*})^{\otimes k}$.
+The iterated-integrals signature $\sigma(X)$ of a $d$-dimensional path $X:[0,1]\to V$ is an element of the tensor series algebra $T((V^{*}))\cong\operatorname{Hom}(T(V),\mathbb{R})$, defined by
+$\sigma(X):T(V^{*})$ $\to\mathbb{R},$
+$e_{i_{1}}^{*}\otimes\dots\otimes e_{i_{k}}^{*}$ $\mapsto\int_{\Delta_{k}}\mathrm{d}X_{i_{1}}(t_{1})\dots\mathrm{d}X_{i_{k}}(t_{k}),$ (7)
+where $\Delta_{k}$ is the simplex $0\leq t_{1}\leq\dots\leq t_{k}\leq 1$.
+The signature matrix of $X$ is $S_{ij}(X):=\sigma(X)(e_{i}^{*}\otimes e_{j}^{*})$.
+It is common to write the tensor $e^{*}_{i_{1}}\otimes\dots\otimes e^{*}_{i_{k}}$ as a word $\mathtt{i}_{1}\dots\mathtt{i}_{k}$, identifying the tensor algebra $T(V^{*})$ with the free associative algebra $\mathbb{R}\langle\mathtt{1},\ldots,\mathtt{d}\rangle$ over the letters 1 to d.
 However, the tensor algebra viewpoint allows us to consider the signature for paths (and later membranes) in any $\mathbb{R}$-vector space $V$, without choosing a basis, which will be useful in the following.
@@
   1. 1.
-    For fixed $u\in\mathbb{R}^{d}$, the $k$-th signature tensor of the linear path $[0,1]\rightarrow\mathbb{R}^{d},t\mapsto ut$ is given by
-    $\sigma^{(k)}_{i_{1},\dots,i_{k}}=\frac{u_{i_{1}}\dots u_{i_{k}}}{k!}$ (10)
+    For fixed $u\in\mathbb{R}^{d}$, the $k$-th signature tensor of the linear path $X:[0,1]\rightarrow\mathbb{R}^{d},t\mapsto ut$ is given by
+    $\sigma(X)(e_{i_{1}}^{*}\otimes\dots\otimes e_{i_{k}}^{*})=\frac{u_{i_{1}}\dots u_{i_{k}}}{k!}$ (8)
     where $1\leq i_{1},\dots,i_{k}\leq d$.
   2. 2.
-    Let $\mathsf{Axis}^{d}:[0,1]\to\mathbb{R}^{d}$ be the canonical axis path from [2, Example 2.1] or [41, eq.
+    Let $\mathsf{Axis}^{d}:[0,1]\to\mathbb{R}^{d}$ be the canonical axis path from [bib:AFS2019, Example 2.1] or [bib:PSS2019, eq.
 (8)] of order $d$, i.e., the unique piecewise linear path with $d$ segments whose graph interpolates the $d+1$ equidistant support points $(\frac{i}{d},\sum_{1\leq j\leq i}e_{j})$ with $0\leq i\leq d$.
@@
 For all $d$, its signature matrix is the upper triangular matrix
-    $\sigma^{(2)}_{i_{1},i_{2}}=\begin{cases}1&\text{if }i_{1}<i_{2}\\ \frac{1}{2}&\text{if }i_{1}=i_{2}\\ 0&\text{if }i_{1}>i_{2}.\end{cases}$ (11)
+    $S_{i_{1},i_{2}}(\mathsf{Axis}^{d})=\begin{cases}1&\text{if }i_{1}<i_{2}\\ \frac{1}{2}&\text{if }i_{1}=i_{2}\\ 0&\text{if }i_{1}>i_{2}.\end{cases}$ (9)
   3. 3.
     We denote by $\mathsf{Mom}^{d}:[0,1]\to\mathbb{R}^{d},t\mapsto(t,\dots,t^{d})$ the moment path of degree $d$. We have the closed formula
-    $\sigma^{(k)}_{i_{1},i_{2},\dots,i_{k}}=\frac{i_{2}i_{3}\dots i_{k}}{(i_{1}+i_{2})(i_{1}+i_{2}+i_{3})\dots(i_{1}+\dots+i_{k})}$ (12)
-    for all $i_{1},\dots,i_{k}$ due to [2, Example 2.2].
+    $\sigma(\mathsf{Mom}^{d})(e_{i_{1}}^{*}\otimes\dots\otimes e_{i_{k}}^{*})=\frac{i_{2}i_{3}\dots i_{k}}{(i_{1}+i_{2})(i_{1}+i_{2}+i_{3})\dots(i_{1}+\dots+i_{k})}$ (10)
+    for all $i_{1},\dots,i_{k}$ due to [bib:AFS2019, Example 2.2].
 $\mathsf{graph}(\mathsf{Axis}^{3}_{1})$
@@
 Compare also Figure 4 for the correspondending axis membrane.
-In the following we recall the ‘nearest’ extension of the path signature (2) for multiple parameters and comment on its various modifications in Remark 2.3.
-Definition 2.2.
-Let $V$ be a finite dimensional vector space and $X:[0,1]^{p}\to V$ a Lipschitz continuous map.
-Let $\partial_{1\dots p}:=\frac{\partial^{p}}{\partial x_{1}\dots\partial x_{p}}$ be the linear partial differential operator.
-Then, since both the integral and $\partial_{1\dots p}$ are linear operators, the following linear form $\sigma=\sigma(X)\in T(V^{*})^{*}$ is well-defined:
-$\sigma:T(V^{*})$ $\to\mathbb{R},$
-$\alpha_{1}\otimes\dots\otimes\alpha_{k}$ $\mapsto\int_{\Delta^{p}}\partial_{1\dots p}(\alpha_{1}\circ X)(t_{11},\dots,t_{1p})\dots\partial_{1\dots p}(\alpha_{k}\circ X)(t_{k1},\dots,t_{kp})\,\mathrm{d}\mathbf{t}$
-where $\mathrm{d}\mathbf{t}=\mathrm{d}t_{11}\dots\mathrm{d}t_{1p}\dots\mathrm{d}t_{k1}\dots\mathrm{d}t_{kp}$ and $\Delta^{p}=\Delta\times\dots\times\Delta$ is the Cartesian product of $p$ simplices.
-We call $\sigma$ the iterated-integrals signature of $X$.
-The $k$-th level signature $\sigma^{(k)}(X)$ is the restriction of $\sigma(X)$ along $(V^{*})^{\otimes k}\to T(V^{*})$, in particular it is naturally an element of $V^{\otimes k}$.
+The path signature (2) can be extended to two parameters in multiple ways.
+We will study the $\mathsf{id}$-signature as defined in [diehl2024signature, Definition 4.1].
+We comment on its various modifications in Remark 2.3.
+Let $V$ be a finite dimensional $\mathbb{R}$-vector space. A membrane in $V$ is a Lipschitz continuous map $X:[0,1]^{2}\to V$.
+Definition 2.2 ($\mathsf{id}$-signature).
+Let $V$ be a finite dimensional vector space and $X$ a membrane in $V$.
+Let $\partial_{12}:=\frac{\partial^{2}}{\partial x_{1}\partial x_{2}}$ be the linear partial differential operator.
+Then, since both the integral and $\partial_{12}$ are linear operators, the following linear form $\sigma(X)\in T(V^{*})^{*}$ is well-defined:
+$\sigma(X):T(V^{*})$ $\to\mathbb{R},$
+$\alpha_{1}\otimes\dots\otimes\alpha_{k}$ $\mapsto\int_{(\mathbf{s},\mathbf{t})\in\Delta^{2}_{k}}\partial_{12}(\alpha_{1}\circ X)(s_{1},t_{1})\dots\partial_{12}(\alpha_{k}\circ X)(s_{k},t_{k})\,\mathrm{d}\mathbf{s}\mathrm{d}\mathbf{t}$
+where $\mathrm{d}\mathbf{s}=\mathrm{d}s_{1}\dots\mathrm{d}s_{k}$, $\mathrm{d}\mathbf{t}=\mathrm{d}t_{1}\dots\mathrm{d}t_{k}$ and $\Delta_{k}^{2}=\Delta_{k}\times\Delta_{k}$.
+We call $\sigma(X)$ the $\mathsf{id}$-signature of $X$.
+The $k$-th level signature $\sigma^{(k)}(X)$ is the restriction of $\sigma(X)$ along $(V^{*})^{\otimes k}\to T(V^{*})$, in particular it is naturally an element of $V^{\otimes k}\cong((V^{*})^{\otimes k})^{*}$.
 Remark 2.3.
   1. 1.
-    For the one-parameter framework $p=1$ in $d$-dimensional space $V=\mathbb{R}^{d}$ equipped with standard basis $e_{1},\dots,e_{d}$ we obtain the classical formulation (2) for paths.
-We refer the interested reader to the text books [23, 21].
+    When ‘replacing’ our square increments $\partial_{12}$ with Jacobian minors we obtain the $\mathsf{id}$-signature introduced in [GLNO2022, Section 4.7].
   2. 2.
-    When ‘replacing’ our square increments $\partial_{1\dots p}$ with Jacobian minors we obtain the $\mathsf{id}$-signature introduced in [26].
+    For membranes and standard basis $e_{1}^{*},\dots,e_{d}^{*}$ of $V^{*}=(\mathbb{R}^{d})^{*}$, we obtain the $\mathsf{id}$-signature according to [diehl2024signature].
+Its level $k$ component can be stored in the $d\times\dots\times d$ array,
+    $\sigma^{(k)}_{i_{1},\dots,i_{k}}(X)$ $:=\sigma(e_{i_{1}}^{*}\otimes\dots\otimes e_{i_{k}}^{*})$
+    $=\int\limits_{\begin{subarray}{c}0\leq s_{1}\leq\dots\leq s_{k}\leq 1\\ 0\leq t_{1}\leq\dots\leq t_{k}\leq 1\end{subarray}}\partial_{12}X_{i_{1}}(s_{1},t_{1})\dots\partial_{12}X_{i_{k}}(s_{k},t_{k})\,\mathrm{d}s_{1}\mathrm{d}t_{1}\dots\mathrm{d}s_{k}\mathrm{d}t_{k}.$
+    In particular, for level $k=2$, we obtain our signature matrix $S(X)=\sigma^{(2)}(X)$ from the introduction, (1).
   3. 3.
-    For membranes, that is $p=2$ parameters and standard basis $e_{1},\dots,e_{d}$ of $V=\mathbb{R}^{d}$, we obtain the $\mathsf{id}$-signature according to [14].
-Its level $k$ component can be stored in the $d\times\dots\times d$ array,
-    $\sigma^{(k)}_{i_{1},\dots,i_{k}}$ $:=\sigma(e_{i_{1}}\otimes\dots\otimes e_{i_{k}})$
-    $=\int\limits_{\begin{subarray}{c}0<s_{1}<\dots<s_{k}<1\\ 0<t_{1}<\dots<t_{k}<1\end{subarray}}\partial_{12}X_{i_{1}}(s_{1},t_{1})\dots\partial_{12}X_{i_{k}}(s_{k},t_{k})\,\mathrm{d}s_{1}\mathrm{d}t_{1}\dots\mathrm{d}s_{k}\mathrm{d}t_{k}.$
-    In particular, for level $k=2$, we obtain our signature matrix $\sigma=\sigma^{(2)}$ from the introduction, (2).
-  4. 4.
-    When ‘replacing’ integrals by sums, and $\partial_{12}$ by its discrete counterpart, we obtain the two-parameter sums signature [18], evaluated at all diagonal matrix compositions.
-For this compare also [14, Example 4.16].
+    When ‘replacing’ integrals by sums, and $\partial_{12}$ by its discrete counterpart, we obtain the two-parameter sums signature [DS22], evaluated at all diagonal matrix compositions.
+For this compare also [diehl2024signature, Example 4.16].
 Remark 2.4.
-It is clear from the definition that the signatures of any two Lipschitz continuous maps $X,Y:[0,1]^{p}\to V$ agree if $\partial_{1...p}(\alpha\circ(X-Y))=0$ for all $\alpha\in V^{*}$.
-For $p=1$ this is just the translation invariance of the signature, for membranes, i.e.
-$p=2$, this is invariance under adding terms only depending on one of the variables.
+It is clear from the definition that the signatures of any two membranes $X,Y:[0,1]^{2}\to V$ agree if $\partial_{12}(\alpha\circ(X-Y))=0$ for all $\alpha\in V^{*}$.
 In particular, if $X:[0,1]^{2}\to V$ is a membrane, we can always find a membrane $X_{\mathrm{red}}$ which vanishes on $[0,1]\times\{0\}\cup\{0\}\times[0,1]$ and has the same signature as $X$, by letting $X_{\mathrm{red}}(s,t):=X(s,t)-X(0,t)-X(s,0)+X(0,0)$.
 See also Lemma 3.17 and Figure 4 for an illustration of piecewise bilinear membranes.
-We provide some first examples for membranes ($p=2$).
+We provide some first examples for membranes.
 Example 2.5.
   1. 1.
-    For fixed $u\in\mathbb{R}^{d}$, the $k$-th signature tensor of the $d$-dimensional bilinear membrane $[0,1]^{2}\rightarrow\mathbb{R}^{d},(s,t)\mapsto ust$ is given by
-    $\sigma^{(k)}_{i_{1},\dots,i_{k}}=\frac{u_{i_{1}}\dots u_{i_{k}}}{(k!)^{2}}$ (13)
+    For fixed $u\in\mathbb{R}^{d}$, the $k$-th signature tensor of the $d$-dimensional bilinear membrane $X:[0,1]^{2}\rightarrow\mathbb{R}^{d},(s,t)\mapsto ust$ is given by
+    $\sigma^{(k)}_{i_{1},\dots,i_{k}}(X)=\frac{u_{i_{1}}\dots u_{i_{k}}}{(k!)^{2}}$ (11)
     where $1\leq i_{1},\dots,i_{k}\leq d$. This formula follows from Corollary 3.5 or can be verified by an immediate computation.
   2. 2.
-    We recall the $2$-dimensional polynomial membrane $X=A\,\mathsf{Mom}^{2,2}$ from the introduction, (7). With integration over the differentials of the two-parameter moment coordinates,
+    We recall the $2$-dimensional polynomial membrane $X=A\,\mathsf{Mom}^{2,2}$ from the introduction, (5).
+    With integration over the differentials of the two-parameter moment coordinates,
     $\int_{0}^{1}\int_{0}^{1}\,\mathrm{d}s\mathrm{d}t=\int_{0}^{1}\int_{0}^{1}2t\,\mathrm{d}s\mathrm{d}t=\int_{0}^{1}\int_{0}^{1}2s\,\mathrm{d}s\mathrm{d}t=\int_{0}^{1}\int_{0}^{1}4st\,\mathrm{d}s\mathrm{d}t=1,$
@@
     $\sigma^{(1)}(X)=\begin{pmatrix}a_{1,1}+a_{1,2}+a_{1,3}+a_{1,4}\\ a_{2,1}+a_{2,2}+a_{2,3}+a_{2,4}\end{pmatrix}\in\mathbb{R}^{2}.$
-    In Example 3.7 and 3.11 we provide closed formulas for its level $2$ and $3$ signature tensors, respectively.
+    In Example 3.7 and 3.10 we provide closed formulas for its level $2$ and $3$ signature tensors, respectively.
 As in the previous example, we can use linearity of the integral to compute level $k$ signatures of transformed functions by corresponding actions on tensors.
 Lemma 2.6 (Equivariance).
-Let $V,W$ be finite dimensional $\mathbb{R}$-vector spaces, and $X:[0,1]^{p}\to V$ be a Lipschitz continuous map. If $A:V\to W$ is linear, then $\sigma(AX)=\sigma(X)\circ T(A^{*})$.
+Let $V,\ W$ be finite dimensional $\mathbb{R}$-vector spaces, $X:[0,1]^{2}\to V$ a membrane and $A:V\to W$ a linear map.
+Then $\sigma(AX)=\sigma(X)\circ T(A^{*})$ where $T(A^{*})$ denotes the map $T(W^{*})\to T(V^{*})$ induced by diagonal action of $A^{*}$ on tensors.
 Proof.
@@
 If $V=\mathbb{R}^{d}$ we can write this in the Tucker format $\sigma^{(k)}(AX)=[\![\sigma^{(k)}(X);A,\dots,A]\!]$.
-Explicitly, if $k=2$ we have
-$\sigma^{(k)}(AX)=A\sigma^{(k)}(X)A^{\top},$
+Explicitly, if $k=2$ and $S(X)$ denotes the signature matrix, we have
+$S(AX)=AS(X)A^{\top},$
 as was already mentioned in the introduction (4).
-The equivariance of signatures is a key property used in [41], allowing the computation of signature tensors for various families of paths from a small collection of dictionaries, also called core tensors.
-In the following section we introduce such dictionaries for membranes.
+The equivariance of signatures is a key property used in [bib:PSS2019], allowing the computation of signature tensors for various families of paths from a small collection of core tensors.
+These are given by signature tensors of special paths called dictionaries.
+In the following section we derive dictionaries for membranes from such dictionaries for paths.
 3 Product membranes
@@
   2. 2.
-    The moment membrane of multidegree $(m,n)$ in the introduction (4) is the product
+    The moment membrane of order $(m,n)$ in the introduction (2) is the product
     $\mathsf{Mom}^{m,n}=\mathsf{Mom}^{m}\boxtimes\mathsf{Mom}^{n}:[0,1]^{2}$ $\to\mathbb{R}^{m}\otimes\mathbb{R}^{n},$ (12)
     where the moment path $\mathsf{Mom}^{m}$ is according to Example 2.1.
-Note that we sometimes omit the last vectorization, i.e., we identity $\mathbb{R}^{m\times n}$ with $\mathbb{R}^{mn}$ whenever the context allows this.
+Note that we sometimes omit the last vectorization, i.e., we identify $\mathbb{R}^{m\times n}$ with $\mathbb{R}^{mn}$ whenever the context allows this.
 For this compare also Corollary 3.6.
-Note that every $d$-dimensional polynomial membrane $X$ with multidegree $(m,n)$ without terms of degree $\leq(1,1)$ is of the form $X=A\circ\mathsf{Mom}^{m,n}$, where $A$ is a linear map, e.g.
+Note that every $d$-dimensional polynomial membrane $X$ of order $(m,n)$ and with terms of order $\geq(1,1)$ is of the form $X=A\circ\mathsf{Mom}^{m,n}$, where $A$ is a linear map, e.g.
 our running Example 2.5 with $m=n=d=2$.
@@
 $\sigma(X\boxtimes Y),(\alpha_{1}\otimes\beta_{1})\otimes\dots\otimes(\alpha_{k}\otimes\beta_{k})\right>$ $=$
-$\sigma(X),\alpha_{1}\otimes\dots\otimes\alpha_{k}\right>$ $\sigma(Y),\beta_{1}\otimes\dots\otimes\beta_{k}\right>$
+$\sigma(X),\alpha_{1}\otimes\dots\otimes\alpha_{k}\right>$ $\sigma(Y),\beta_{1}\otimes\dots\otimes\beta_{k}\right>.$
 Caveat 3.4.
-Note that the tensor product of linear forms $\sigma(X):T(V^{*})\to\mathbb{R}$, $\sigma(Y):T(V^{*})\to\mathbb{R}$ is not the tensor product in the tensor series space $T(V^{*})$; the formula above should not be confused with Chen’s formula for the signature of a concatenation of paths.
+Note that the tensor product of linear forms $\sigma(X):T(V^{*})\to\mathbb{R}$, $\sigma(Y):T(V^{*})\to\mathbb{R}$ is not the tensor product in the tensor series space $T(V^{*})^{*}$; the formula above should not be confused with Chen’s formula for the signature of a concatenation of paths.
 Before proving Proposition 3.3, we provide examples and discuss several implications.
@@
 Unravel the factorization of $\sigma(X\boxtimes Y)$ in Proposition 3.3. ∎
-If $X:[0,1]^{2}\to V\otimes W$ is a product membrane, it can be convenient to use a vectorization $\nu:V\otimes W\cong\mathbb{R}^{\dim V}$, for example to obtain an actual signature matrix.
+If $X:[0,1]^{2}\to V\otimes W$ is a product membrane, it can be convenient to use a vectorization $\nu:V\otimes W\cong\mathbb{R}^{\dim(V)\cdot\dim(W)}$, for example to obtain an actual signature matrix.
 Using Lemma 2.6, we see that $\sigma(\nu X)$ then factors as
-$T(\mathbb{R}^{d})\to T((V\otimes W)^{*})\to T(V^{*})\otimes T(W^{*})\to\mathbb{R}$
+$T(\mathbb{R}^{\dim(V)\cdot\dim(W)})\to T((V\otimes W)^{*})\to T(V^{*})\otimes T(W^{*})\to\mathbb{R}$
 where the first map is $T(\nu^{*})$.
@@
 Let $X:[0,1]\to\mathbb{R}^{m}$ and $Y:[0,1]\to\mathbb{R}^{n}$ be two paths and let $\mathsf{kron}(X,Y):[0,1]^{2}\to\mathbb{R}^{mn}$ denote the membrane that maps $(s,t)$ to the Kronecker product of the column vectors $X(s)$ and $Y(t)$.
-Then
-$\sigma^{(2)}(\mathsf{kron}(X,Y))$ $=\sigma^{(2)}(X)\otimes\sigma^{(2)}(Y)$
+Then we have the following identity between signature matrices:
+$S(\mathsf{kron}(X,Y))$ $=S(X)\otimes S(Y)$
 where $\otimes$ denotes the Kronecker product of matrices.
@@
 We return to our running Example 2.5 with $X=A\,\mathsf{Mom}^{2,2}$ and matrix $A=(a_{i,j})_{1\leq i\leq 2,1\leq j\leq 4}$. With Corollary 3.6 and (10) we get the core matrix of the moment membrane,
-$C^{(2)}:=\sigma^{(2)}(\mathsf{vec}\circ\mathsf{Mom}^{2,2})=\sigma^{(2)}(\mathsf{Mom}^{2})\otimes\sigma^{(2)}(\mathsf{Mom}^{2})=\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}\otimes\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}$
-as it is presented in the introduction, (5). With $S^{(2)}:=\sigma^{(2)}(X)=AC^{(2)}A^{\top}$ we provide the three missing homogeneous entries,
-$S^{(2)}_{2,1}$ $=\frac{1}{4}a_{1,1}a_{2,1}+\frac{1}{6}a_{1,1}a_{2,2}+\frac{1}{6}a_{1,1}a_{2,3}+\frac{1}{9}a_{1,1}a_{2,4}+\frac{1}{3}a_{2,1}a_{1,2}+\frac{1}{3}a_{2,1}a_{1,3}$
+$C:=\sigma^{(2)}(\mathsf{Mom}^{2,2})=\sigma^{(2)}(\mathsf{Mom}^{2})\otimes\sigma^{(2)}(\mathsf{Mom}^{2})=\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}\otimes\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}$
+as it is presented in the introduction, (3). With $S(X)=\sigma^{(2)}(X)=ACA^{\top}$ we provide the three missing homogeneous entries,
+$S(X)_{2,1}$ $=\frac{1}{4}a_{1,1}a_{2,1}+\frac{1}{6}a_{1,1}a_{2,2}+\frac{1}{6}a_{1,1}a_{2,3}+\frac{1}{9}a_{1,1}a_{2,4}+\frac{1}{3}a_{2,1}a_{1,2}+\frac{1}{3}a_{2,1}a_{1,3}$
 $\quad+\frac{4}{9}a_{2,1}a_{1,4}+\frac{1}{4}a_{1,2}a_{2,2}+\frac{2}{9}a_{1,2}a_{2,3}+\frac{1}{6}a_{1,2}a_{2,4}+\frac{2}{9}a_{2,2}a_{1,3}+\frac{1}{3}a_{2,2}a_{1,4}$
 $\quad+\frac{1}{4}a_{1,3}a_{2,3}+\frac{1}{6}a_{1,3}a_{2,4}+\frac{1}{3}a_{2,3}a_{1,4}+\frac{1}{4}a_{1,4}a_{2,4}$
-$S^{(2)}_{1,2}$ $=\frac{1}{4}a_{1,1}a_{2,1}+\frac{1}{3}a_{1,1}a_{2,2}+\frac{1}{3}a_{1,1}a_{2,3}+\frac{4}{9}a_{1,1}a_{2,4}+\frac{1}{6}a_{2,1}a_{1,2}+\frac{1}{6}a_{2,1}a_{1,3}$
+$S(X)_{1,2}$ $=\frac{1}{4}a_{1,1}a_{2,1}+\frac{1}{3}a_{1,1}a_{2,2}+\frac{1}{3}a_{1,1}a_{2,3}+\frac{4}{9}a_{1,1}a_{2,4}+\frac{1}{6}a_{2,1}a_{1,2}+\frac{1}{6}a_{2,1}a_{1,3}$
 $\quad+\frac{1}{9}a_{2,1}a_{1,4}+\frac{1}{4}a_{1,2}a_{2,2}+\frac{2}{9}a_{1,2}a_{2,3}+\frac{1}{3}a_{1,2}a_{2,4}+\frac{2}{9}a_{2,2}a_{1,3}+\frac{1}{6}a_{2,2}a_{1,4}$
@@
 and
-$S^{(2)}_{2,2}$ $=\frac{1}{4}a_{2,1}^{2}+\frac{1}{2}a_{2,1}a_{2,2}+\frac{1}{2}a_{2,1}a_{2,3}+\frac{5}{9}a_{2,1}a_{2,4}+\frac{1}{4}a_{2,2}^{2}+\frac{4}{9}a_{2,2}a_{2,3}+\frac{1}{2}a_{2,2}a_{2,4}$
+$S(X)_{2,2}$ $=\frac{1}{4}a_{2,1}^{2}+\frac{1}{2}a_{2,1}a_{2,2}+\frac{1}{2}a_{2,1}a_{2,3}+\frac{5}{9}a_{2,1}a_{2,4}+\frac{1}{4}a_{2,2}^{2}+\frac{4}{9}a_{2,2}a_{2,3}+\frac{1}{2}a_{2,2}a_{2,4}$
 $\quad+\frac{1}{4}a_{2,3}^{2}+\frac{1}{2}a_{2,3}a_{2,4}+\frac{1}{4}a_{2,4}^{2}$
-using equivariance, Lemma 2.6. The polynomial $S_{1,1}^{(2)}$ is already provided in the introduction.
-We illustrate the membrane $X$ in the first row of Figure 3, for $A=\begin{pmatrix}1&-1&1&1\\ 1&1&0&-1\end{pmatrix}$.
+using equivariance, Lemma 2.6. The polynomial $S(X)_{1,1}$ is already provided in the introduction.
+We illustrate the membrane $X$ in the first row of Figure 3, for the explicit choice $A=\begin{pmatrix}1&-1&1&1\\ 1&1&0&-1\end{pmatrix}$.
 In particular, we provide the graphs of the coordinate functions $\mathsf{graph}(X_{1})$ and $\mathsf{graph}(X_{2})$, and a plot of the image $\mathsf{Im}(X)$ in $\mathbb{R}^{2}$.
@@
 Note that $X$ maps three sides of the square $[0,1]^{2}$ to the origin.
-The right-most picture shows an ”untangled” version of the membrane $X$ which has the same signature, cf.
+The right-most picture shows an “untangled” version of the membrane $X$ which has the same signature, cf.
 Remark 2.4.
@@
 Proof of Corollary 3.6.
-Choosing the vectorization $\nu:\mathbb{R}^{m}\otimes\mathbb{R}^{n}\to\mathbb{R}^{mn},e_{i}\otimes e_{j}\mapsto e_{n(i-1)+j}$, we see that $\mathsf{kron}=\nu\circ X\boxtimes Y$. Then we have
+Choosing the vectorization $\nu:\mathbb{R}^{m}\otimes\mathbb{R}^{n}\to\mathbb{R}^{mn},e_{i}\otimes e_{j}\mapsto e_{n(i-1)+j}$, we see that $\mathsf{kron}(X,Y)=\nu\circ X\boxtimes Y$. Then we have
 $\sigma(\nu(X\boxtimes Y)),e_{i}\otimes e_{j}\right>=\left<\sigma(X\boxtimes Y),\nu^{*}(e_{i})\otimes\nu^{*}(e_{j})\right>$
@@
 $\sigma^{(2)}(\nu(X\boxtimes Y))=\sigma^{(2)}(X)\otimes\sigma^{(2)}(Y).$
-As another consequence of Proposition 3.3 and equivariance we obtain the formula from [14, Example 4.2] for the Hadamard product of two paths:
+As another consequence of Proposition 3.3 and equivariance we obtain the formula from [diehl2024signature, Example 4.2] for the Hadamard product of two paths:
 Corollary 3.8.
-If $V=W=\mathbb{R}^{d}$ then
+Let $\mu:\mathbb{R}^{d}\otimes\mathbb{R}^{d}\to\mathbb{R}^{d}$ denote the componentwise multiplication map. If $V=W=\mathbb{R}^{d}$ then
 $\sigma(\mu(X\boxtimes Y))=\sigma(X)\cdot\sigma(Y)$ (13)
-as morphisms $T(\mathbb{R}^{d})\to\mathbb{R}$, and where $\mu:\mathbb{R}^{d}\otimes\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the componentwise multiplication map.
+as morphisms $T(\mathbb{R}^{d})\to\mathbb{R}$. In particular, $S(X\odot Y)=S(X)\odot S(Y)$ where $\odot$ denotes the Hadamard product of vectors resp. matrices.
 Proof.
@@
 and similarly for $Y$; thus reordering the integral reveals the claimed equality. ∎
-Remark 3.9.
-Proposition 3.3 and its proof immediately generalize to higher $p$-parameter patches $[0,1]^{p}\rightarrow\mathbb{R}^{d}$.
 In order to provide an explicit computation for higher tensor levels $k\geq 3$, we introduce some further notation.
-Remark and Notation 3.10.
+Remark and Notation 3.9.
 If $X:[0,1]\to\mathbb{R}^{m}$ and $Y:[0,1]\to\mathbb{R}^{n}$ are two paths, then by Proposition 3.3 the signature of $X\boxtimes Y$ is a linear form on $T((\mathbb{R}^{m}\otimes\mathbb{R}^{n})^{*})$ which can be identified with the free associative algebra $\mathbb{R}\langle(i,j)\rangle$ over tuples of letters $i\in[m]$ and $j\in[n]$.
@@
 Example 3.10.
-With Proposition 3.3, our identification in 3.10, and the closed formula for signatures of polynomial paths in Example 2.1, we obtain a closed formula for the level $k$ tensor of the moment membrane $\mathsf{Mom}^{m,n}$ from (14),
-$\langle\sigma(\mathsf{Mom}^{m,n}),(i_{1},j_{1})\dots(i_{k},j_{k})\rangle=\frac{i_{2}\dots i_{k}}{(i_{1}+i_{2})\dots(i_{1}+\dots+i_{k})}\frac{j_{2}\dots j_{k}}{(j_{1}+j_{2})\dots(j_{1}+\dots+j_{k})}$
-where $i_{1},\dots,i_{k}\in[m]$ and $j_{1},\dots,j_{k}\in[n]$. For instance for the $2$-dimensional polynomial membrane from our running example Example 2.5, we obtain the level $k=3$ core tensor
-$C^{(3)}:=\scriptsize\left(\begin{array}[]{cccc|cccc|cccc|cccc}\frac{1}{36}&\frac{1}{36}&\frac{1}{36}&\frac{1}{36}&\frac{1}{24}&\frac{2}{45}&\frac{1}{24}&\frac{2}{45}&\frac{1}{24}&\frac{1}{24}&\frac{2}{45}&\frac{2}{45}&\frac{1}{16}&\frac{1}{15}&\frac{1}{15}&\frac{16}{225}\\[4.0pt] \frac{1}{72}&\frac{1}{60}&\frac{1}{72}&\frac{1}{60}&\frac{1}{45}&\frac{1}{36}&\frac{1}{45}&\frac{1}{36}&\frac{1}{48}&\frac{1}{40}&\frac{1}{45}&\frac{2}{75}&\frac{1}{30}&\frac{1}{24}&\frac{8}{225}&\frac{2}{45}\\[4.0pt] \frac{1}{72}&\frac{1}{72}&\frac{1}{60}&\frac{1}{60}&\frac{1}{48}&\frac{1}{45}&\frac{1}{40}&\frac{2}{75}&\frac{1}{45}&\frac{1}{45}&\frac{1}{36}&\frac{1}{36}&\frac{1}{30}&\frac{8}{225}&\frac{1}{24}&\frac{2}{45}\\[4.0pt] \frac{1}{144}&\frac{1}{120}&\frac{1}{120}&\frac{1}{100}&\frac{1}{90}&\frac{1}{72}&\frac{1}{75}&\frac{1}{60}&\frac{1}{90}&\frac{1}{75}&\frac{1}{72}&\frac{1}{60}&\frac{4}{225}&\frac{1}{45}&\frac{1}{45}&\frac{1}{36}\end{array}\right)$
-of shape $4\times 4\times 4$, and with equivariance (Lemma 2.6), the level $k=3$ signature tensor $\sigma^{(3)}(X)=[\![C^{(3)};A,A,A]\!]$ of shape $2\times 2\times 2$.
+With Proposition 3.3, our identification in 3.9, and the closed formula for signatures of polynomial paths in Example 2.1, we obtain a closed formula for the level $k$ tensor of the moment membrane $\mathsf{Mom}^{m,n}$ from (12),
+$\langle\sigma(\mathsf{Mom}^{m,n})$ $,(i_{1},j_{1})\dots(i_{k},j_{k})\rangle$
+$=\frac{i_{2}\dots i_{k}}{(i_{1}+i_{2})\dots(i_{1}+\dots+i_{k})}\frac{j_{2}\dots j_{k}}{(j_{1}+j_{2})\dots(j_{1}+\dots+j_{k})}$
+where $i_{1},\dots,i_{k}\in[m]$ and $j_{1},\dots,j_{k}\in[n]$.
+For instance for the $2$-dimensional polynomial membrane from our running example Example 2.5, we obtain the level $k=3$ core tensor, here denoted by $C^{(3)}$,
+$\scriptsize\left(\begin{array}[]{cccc|cccc|cccc|cccc}\frac{1}{36}&\frac{1}{36}&\frac{1}{36}&\frac{1}{36}&\frac{1}{24}&\frac{2}{45}&\frac{1}{24}&\frac{2}{45}&\frac{1}{24}&\frac{1}{24}&\frac{2}{45}&\frac{2}{45}&\frac{1}{16}&\frac{1}{15}&\frac{1}{15}&\frac{16}{225}\\[4.0pt] \frac{1}{72}&\frac{1}{60}&\frac{1}{72}&\frac{1}{60}&\frac{1}{45}&\frac{1}{36}&\frac{1}{45}&\frac{1}{36}&\frac{1}{48}&\frac{1}{40}&\frac{1}{45}&\frac{2}{75}&\frac{1}{30}&\frac{1}{24}&\frac{8}{225}&\frac{2}{45}\\[4.0pt] \frac{1}{72}&\frac{1}{72}&\frac{1}{60}&\frac{1}{60}&\frac{1}{48}&\frac{1}{45}&\frac{1}{40}&\frac{2}{75}&\frac{1}{45}&\frac{1}{45}&\frac{1}{36}&\frac{1}{36}&\frac{1}{30}&\frac{8}{225}&\frac{1}{24}&\frac{2}{45}\\[4.0pt] \frac{1}{144}&\frac{1}{120}&\frac{1}{120}&\frac{1}{100}&\frac{1}{90}&\frac{1}{72}&\frac{1}{75}&\frac{1}{60}&\frac{1}{90}&\frac{1}{75}&\frac{1}{72}&\frac{1}{60}&\frac{4}{225}&\frac{1}{45}&\frac{1}{45}&\frac{1}{36}\end{array}\right)$
+which is of shape $4\times 4\times 4$.
+With equivariance (Lemma 2.6), this results in the level $k=3$ signature tensor $\sigma^{(3)}(X)=[\![C^{(3)};A,A,A]\!]$ of shape $2\times 2\times 2$.
 We provide its first homogeneous entry,
@@
 The other (polynomial) entries of $\sigma^{(3)}(X)$ are similar and are therefore omitted here.
-Inspired by [18, Theorem 4.5], we obtain the following computational implication.
 Corollary 3.11.
-For every $d$-dimensional polynomial membrane $X$ of multi-degree $(m,n)$ we can compute $\sigma^{(k)}(X)$ in $\mathcal{O}(m^{k}n^{k}+kd^{k+1}mn)$ elementary operations.
-Proof.
-We use the closed formula in Example 3.11 and $\mathsf{vec}^{\otimes k}$ to obtain the core tensor $C^{(k)}$ in $\mathcal{O}(m^{k}n^{k})$. The remaining follows with equivariance. ∎
+For every $d$-dimensional polynomial membrane $X$ of order $(m,n)$ we can compute $\sigma^{(k)}(X)$ in $\mathcal{O}(m^{k}n^{k}+kd^{k+1}mn)$ elementary operations.
 Remark 3.12.
@@
 Definition 3.13.
-A piecewise bilinear membrane is a membrane $X$ such that there are some $m,n\in\mathbb{N}$ with the property that $X$ is biaffine on all squares $[\frac{i-1}{m},\frac{i}{m}]\times[\frac{j-1}{n},\frac{j}{n}]$ with $i\in[m]$ and $j\in[n]$.
+A piecewise bilinear membrane is a membrane $X$ such that there are some $m,n\in\mathbb{N}$ with the property that $X$ is biaffine on all rectangles $[\frac{i-1}{m},\frac{i}{m}]\times[\frac{j-1}{n},\frac{j}{n}]$ with $i\in[m]$ and $j\in[n]$.
 We say that $X$ has order $(m,n)$.
-Our interest in this class of membranes is threefold: from the path case, examined in [2], we expect a certain relation of this class of membranes to the polynomial membranes in terms of their signatures, e.g.
+Our interest in this class of membranes is threefold: from the path case, examined in [bib:AFS2019], we expect a certain relation of this class of membranes to the polynomial membranes in terms of their signatures, e.g.
 Theorem 4.1.
-We also know that the recovery problem is solvable for piecewise linear paths of low order from [41, Corollary 6.3], and we would like to obtain a similar result for membranes (5.4).
-Finally, they form a rather natural class that can be used for the interpolation of discrete data to fit into the frame of iterated integrals, e.g.
-image data used in [48].
+We also know that the recovery problem is solvable for piecewise linear paths of low order from [bib:PSS2019, Corollary 6.3], and we would like to obtain a similar result for membranes (5.3).
+Finally, they form a rather natural class that can be used for the interpolation of discrete data, e.g.
+image data used in [ZLT22].
 Bilinear interpolation of $d$-dimensional discrete data, arranged on an $(m+1)\times(n+1)$ grid, is a well-known procedure, e.g.
-in image processing [6, Chapter 3.7.2] or in the field of finite elements [49, Chapters 5.1.3.2 and 6.2.2.1], [20, Chapter 1.3.2].
+in image processing [bovik2009essential, Chapter 3.7.2] or in the field of finite elements [zienkiewicz_finite_2005, Chapters 5.1.3.2 and 6.2.2.1], [elman2014finite, Chapter 1.3.2].
 The approach involves considering the four (unique) nodal bilinear basis functions on the reference rectangle, where each function is $1$ at one corner and 00 at all others.
@@
 $\psi_{i,j}\left(\frac{i-1}{m},\frac{j}{n}\right)$ $=0,\quad\psi_{i,j}\left(\frac{i}{m},\frac{j}{n}\right)=1.$ (15)
-By solving (3.2) in $c$ we obtain $c=\left(mn,m(1-j),n(1-i),(i-1)(j-1)\right)$ and hence $\psi=\mathsf{Axis}^{m,n}$ when restricted on the square $\left[\frac{i-1}{m},\frac{i}{m}\right]\times\left[\frac{j-1}{n},\frac{j}{n}\right]$.
-The claim follows with a similar (linear) system for the remaining squares.
+By solving (3.2) in $c$ we obtain $c=\left(mn,m(1-j),n(1-i),(i-1)(j-1)\right)$ and hence $\psi=\mathsf{Axis}^{m,n}$ when restricted on the rectangle $\left[\frac{i-1}{m},\frac{i}{m}\right]\times\left[\frac{j-1}{n},\frac{j}{n}\right]$.
+The claim follows with a similar (linear) system for the remaining rectangle.
 ∎
@@
 Example 3.16.
-From [2, Example 2.1] we know that $\sigma(\mathsf{Axis}^{m})_{i_{1},\dots,i_{k}}$ is zero unless $i_{1}\leq\dots\leq i_{k}$, and $\frac{1}{k!}$ times the number of distinct permutations of the string $i_{1}\dots i_{k}$ otherwise.
+From [bib:AFS2019, Example 2.1] we know that $\sigma(\mathsf{Axis}^{m})_{i_{1},\dots,i_{k}}$ is zero unless $i_{1}\leq\dots\leq i_{k}$, and $\frac{1}{k!}$ times the number of distinct permutations of the string $i_{1}\dots i_{k}$ otherwise.
 Explicitly, with (9) the signature matrix is
@@
 Proof.
-Clearly, $\{\mathsf{Axis}^{m,n}_{i,j}\mid i,j\}$ is an $\mathbb{R}$-basis of the space of $1$-dimensional piecewise bilinear membranes which are 00 on $[0,1]\times 0\cup 0\times[0,1]$.
+The coordinate functions $\{\mathsf{Axis}^{m,n}_{i,j}\mid i\in\{1,\dots,m\},j\in\{1,\dots,n\}\}$ of the canonical axis membrane form an $\mathbb{R}$-basis of the space of $1$-dimensional piecewise bilinear membranes.
 We can extend this basis via $m+n-1$ functions which are piecewise linear in $s$ and constant in $t$, or vice versa.
@@
 ∎
-The decomposition is illustrated in Figure 4. See also [18, Lemma 2.22] for a similar result applied on discrete data, using two-parameter sums signatures.
+The decomposition is illustrated in Figure 4.
+See also [DS22, Lemma 2.22] for a similar result applied on discrete data, using two-parameter sums signatures.
+Inspired by [DEFT22, Algorithm 1] and [DS22, Theorem 4.5], we obtain the following computational result.
 Theorem 3.18.
 Let $X$ be a piecewise bilinear membrane in $\mathbb{R}^{d}$ of order $(m,n)$ and $w\in T(\mathbb{R}^{d})$ a word of length $k$.
-Then we can compute $\langle\sigma(X),w\rangle$ in $\mathcal{O}(k^{2}mn^{k})$.
+Then we can compute $\langle\sigma(X),w\rangle$ in $\mathcal{O}(k^{3}mn)$.
+In particular, the $k$-th signature tensor $\sigma^{(k)}(X)$ is computable in $\mathcal{O}(d^{k}k^{3}mn)$.
 Proof.
-With Lemma 3.18 we assume that $X$ is 00 on $\{0\}\times[0,1]\cup[0,1]\times\{0\}$.
-Let $\Lambda$ denote the piecewise linear path in $\mathbb{R}^{n}$ of order $n$ with control points $(0,e_{1},\dots,e_{n})$.
-Then we can write
-$X(s,t)=\sum_{i=0}^{n}X(s,\frac{i}{n})\Lambda_{i}(t)$
-Indeed, the right hand side is clearly piecewise bilinear of order $(m,n)$ and agrees with $X$ on all control points.
-We can write this as a linear transformation $A$ applied to the product membrane $Y\boxtimes\Lambda$ in $\mathbb{R}^{nd}\otimes\mathbb{R}^{n}$ where
-$Y(s):=\begin{pmatrix}X(s,\frac{1}{n})\\ \vdots\\ X(s,1)\end{pmatrix}$
-is a piecewise linear path of order $m$. The dual of $A$ is given via
-$e_{i}^{*}\mapsto\sum_{j=1}^{n}e_{i+(j-1)d}^{*}\otimes e_{j}^{*}.$
-In particular, we obtain a decomposition of $(A^{*})^{\otimes k}w$ into $n^{k}$ biwords of length $k$.
-By Proposition 3.3 and Chen’s identity (for paths) we can compute each of the corresponding signatures in $\mathcal{O}(k^{2}m)$.
-Thus, by Lemma 2.6, we obtain $\langle\sigma(X),w\rangle$ in $\mathcal{O}(k^{2}mn^{k})$.
+For $a,b\in\mathbb{N}$, write $I_{a,b}$ for the rectangle $[\frac{a-1}{m},\frac{a}{m}]\times[\frac{b-1}{n},\frac{b}{n}]$. For a word $w=i_{1}\ldots i_{j}$ and $(u,v)\in[0,1]^{2}$ define
+$f_{w}(u,v):=\int_{\Delta_{j}(u,v)}\partial_{12}X_{i_{1}}(s_{1},t_{1})\ldots\partial_{12}X_{i_{j}}(s_{j},t_{j})\ \mathrm{d}t_{1}\ldots\mathrm{d}t_{j}\mathrm{d}s_{1}\ldots\mathrm{d}s_{j}$
+where $\Delta_{j}(u,v)$ denotes the set
+$\{0\leq s_{1}\leq\dots\leq s_{j}\leq u\}\times\{0\leq t_{1}\leq\dots\leq t_{j}\leq v\}.$
+Set $f_{w}\equiv 1$ for $w$ the empty word. Then by definition, $f_{w}(1,1)=\langle\sigma(X),w\rangle$ and we have the recursion
+$f_{wi}(u,v)=\int_{0}^{v}\int_{0}^{u}f_{w}(s,t)\cdot\partial_{12}X_{i}(s,t)\ \mathrm{d}s\mathrm{d}t$
+for any letter $i$. Thus it suffices to show that we can compute $f_{wi}$ from $f_{w}$ in time $\mathcal{O}(k^{2}mn)$. We can rewrite the integral above as the sum of the four terms
+$\mathrm{Bl}_{wi}(u,v):=\sum_{a=1}^{\lfloor vm\rfloor}\sum_{b=1}^{\lfloor un\rfloor}\int_{\frac{a-1}{m}}^{\frac{a}{m}}\int_{\frac{b-1}{n}}^{\frac{b}{n}}f_{w}(s,t)\cdot\partial_{12}X_{i}(s,t)\ \mathrm{d}s\mathrm{d}t,$
+$\mathrm{U}_{wi}(u,v):=\sum_{a=1}^{\lfloor vm\rfloor}\int_{\frac{a-1}{m}}^{\frac{a}{m}}\int_{\frac{\lfloor un\rfloor}{n}}^{u}f_{w}(s,t)\cdot\partial_{12}X_{i}(s,t)\ \mathrm{d}s\mathrm{d}t,$
+$\mathrm{R}_{wi}(u,v):=\sum_{b=1}^{\lfloor un\rfloor}\int_{\frac{b-1}{n}}^{\frac{b}{n}}\int_{\frac{\lfloor vm\rfloor}{m}}^{v}f_{w}(s,t)\cdot\partial_{12}X_{i}(s,t)\ \mathrm{d}s\mathrm{d}t,$
+and
+$\mathrm{Ur}_{wi}(u,v):=\int_{\frac{\lfloor vm\rfloor}{m}}^{v}\int_{\frac{\lfloor un\rfloor}{n}}^{u}f_{w}(s,t)\cdot\partial_{12}X_{i}(s,t)\ \mathrm{d}s\mathrm{d}t,$
+see Figure 5.
+Figure 5: The four terms in the proof of Theorem 3.18 correspond to the division of a rectangle in the $m\times n$ grid into four regions. In this example, $m=n=3$ and $(u,v)$ lies in $I_{3,2}$.
+Note that if $f_{w}$ is given by a polynomial in $u$ and $v$ on all $I_{a,b}$, then again for all $a,b$ all four terms are polynomials in $u$ and $v$ for $(u,v)\in I_{a,b}$ and thus $f_{wi}$ will be a polynomial on all $I_{a,b}$ as well, since $\partial_{12}X_{i}$ is piecewise constant.
+As for the empty word $w=\emptyset$ we have $f_{w}\equiv 1$ it follows inductively that $f_{w}$ will always be polynomial of bidegree $(j,j)$ (where $j$ denotes the length of $w$) on each $I_{a,b}$.
+Thus, forming the at most $mn$ integrals in the definition of $\mathrm{Bl}_{wi},\mathrm{U}_{wi},\mathrm{R}_{wi}$ and $\mathrm{Ur}_{wi}$ takes time $\mathcal{O}(mnk^{2})$.
+Then $f_{wi}$ can be computed via cumulative sums in $\mathcal{O}(mn)$ so that we get total runtime $\mathcal{O}(k^{2}mn)$, as desired.
 ∎
-4 Varieties of two-parameter signature matrices
-Our primary interest lies in the image of polynomial resp.
-piecewise bilinear membranes under the signature map at a specific level $k$.
-Here, we focus on the special case $k=2$.
-Following the approach in [2] we restrict to families of $d$-dimensional membranes,
-$\left\{\sigma^{(2)}(X)\;\begin{array}[]{|l}X:[0,1]^{2}\rightarrow{\mathbb{R}}^{d}\text{ polynomial}\\ \text{with multi-degree}\leq(m,n)\end{array}\right\}$ (19)
+Corollary 3.19.
+For every piecewise bilinear membrane $X$ in $\mathbb{R}^{d}$ of order $(m,n)$ we can compute its signature matrix $S(X)=\sigma^{(2)}(X)$ in $\mathcal{O}(d^{2}mn)$.
+4 Signature matrix varieties
+For the remainder of this article, we focus on the second level signature, that is, on signature matrices of membranes $S=\sigma^{(2)}$.
+Following the approach in [bib:AFS2019] we restrict to families of $d$-dimensional membranes,
+$\left\{S(X)\;\begin{array}[]{|l}X:[0,1]^{2}\rightarrow{\mathbb{R}}^{d}\text{ polynomial}\\ \text{with order}\leq(m,n)\end{array}\right\}$ (17)
 and
-$\left\{\sigma^{(2)}(X)\;\begin{array}[]{|l}X:[0,1]^{2}\rightarrow{\mathbb{R}}^{d}\text{ piecewise bilinear}\\ \text{of order }\leq(m,n)\end{array}\right\}$ (20)
+$\left\{S(X)\;\begin{array}[]{|l}X:[0,1]^{2}\rightarrow{\mathbb{R}}^{d}\text{ piecewise bilinear}\\ \text{of order }\leq(m,n)\end{array}\right\}$ (18)
 with fixed $m,n\in\mathbb{N}$. By Lemma 2.6 and Lemma 3.17, both (17) and (18) can be considered as the image of a polynomial map of degree two,
-$\sigma^{(2)}:\mathbb{R}^{d\times mn}\rightarrow\mathbb{R}^{d^{2}},A\mapsto ACA^{\top},$ (21)
-where $C\in\mathbb{R}^{mn\times mn}$ is either the signature core matrix $\sigma^{(2)}(\nu\mathsf{Mom}^{m,n})$ (Example 3.11) for polynomial membranes or the axis core matrix $\sigma^{(2)}(\nu\mathsf{Axis}^{m,n})$ for the piecewise bilinear family (Example 3.17) where we choose $\nu:\mathbb{R}^{m}\otimes\mathbb{R}^{n}\to\mathbb{R}^{mn}$ as in the proof of Corollary 3.6.
-In particular, these signature images are semialgebraic subsets of $\mathbb{R}^{d\times d}$ by the Tarski-Seidenberg theorem ([5, Theorem 2.2.1]).
+$S:\mathbb{R}^{d\times mn}\rightarrow\mathbb{R}^{d^{2}},A\mapsto ACA^{\top}.$ (19)
+Here $C\in\mathbb{R}^{mn\times mn}$ is either the signature core matrix $S(\nu\mathsf{Mom}^{m,n})$ according to Example 3.10 for polynomial membranes, or the axis core matrix $S(\nu\mathsf{Axis}^{m,n})$ for the piecewise bilinear family from Example 3.16.
+We choose the vectorization $\nu:\mathbb{R}^{m}\otimes\mathbb{R}^{n}\to\mathbb{R}^{mn}$ as in the proof of Corollary 3.6.
+In particular, these signature images are semialgebraic subsets of $\mathbb{R}^{d\times d}$ by the Tarski-Seidenberg theorem ([bochnak2013real, Theorem 2.2.1]).
 For convenience, we omit $\nu$ from the notation in the following.
-Using our result on the signature of product membranes, we see that [2, Theorem 3.3] generalizes immediately:
+Using our result on the signature of product membranes, we see that [bib:AFS2019, Theorem 3.3] generalizes immediately:
 Theorem 4.1.
@@
 Proof.
-By definition of the two sets, it suffices to find an invertible matrix $A$ such that $\sigma^{(2)}(\mathsf{Mom}^{m,n})=A\sigma^{(2)}(\mathsf{Axis}^{m,n})A^{\top}$.
-Now by Corollary 3.6, we have $\sigma^{(2)}(\mathsf{Mom}^{m,n})=\sigma^{(2)}(\mathsf{Mom}^{m})\otimes\sigma^{(2)}(\mathsf{Mom}^{n})$ and $\sigma^{(2)}(\mathsf{Axis}^{m,n})=\sigma^{(2)}(\mathsf{Axis}^{m})\otimes\sigma^{(2)}(\mathsf{Axis}^{n})$.
-Moreover, in the proof of [2, Theorem 3.3] it is shown that for arbitrary $\ell$ there is an invertible $\ell\times\ell$ matrix $H_{\ell}$ such that $\sigma^{(2)}(\mathsf{Mom}_{\ell})(w)=H_{\ell}\sigma^{(2)}(\mathsf{Axis}_{\ell})H_{\ell}^{\top}$.
+By definition of the two sets, it suffices to find an invertible matrix $A$ such that $S(\mathsf{Mom}^{m,n})=AS(\mathsf{Axis}^{m,n})A^{\top}$.
+Now by Corollary 3.6, we obtain that $S(\mathsf{Mom}^{m,n})=S(\mathsf{Mom}^{m})\otimes S(\mathsf{Mom}^{n})$ and $S(\mathsf{Axis}^{m,n})=S(\mathsf{Axis}^{m})\otimes S(\mathsf{Axis}^{n})$.
+Moreover, in the proof of [bib:AFS2019, Theorem 3.3] it is shown that for arbitrary $\ell$ there is an invertible $\ell\times\ell$ matrix $H_{\ell}$ such that $S(\mathsf{Mom}^{\ell})(w)=H_{\ell}S(\mathsf{Axis}^{\ell})H_{\ell}^{\top}$.
 Thus, we can choose $A:=H_{m}\otimes H_{n}$.
@@
 We want to study the polynomials that vanish on this set, recorded in the homogeneous prime ideals $P_{d,m,n}$.
-Equivalently, we examine the tightest outer approximation by an algebraic variety, i.e., we move to $\mathbb{C}$ and consider the Zariski closure $\mathcal{P}_{d,m,n}$ of (19) $=$ (20) with arbitrary $d,m,n\in\mathbb{N}$.
-For its vanishing ideal we consider the graph of (21), which is an algebraic variety with vanishing ideal
+Equivalently, we examine the tightest outer approximation by an algebraic variety.
+As it is common in applied algebraic geometry, we move to $\mathbb{C}$ and consider the Zariski closure $\mathcal{M}_{d,m,n}$ of the image of $S_{\mathbb{C}}:=\mathbb{C}\otimes S$ for arbitrary $d,m,n\in\mathbb{N}$.
+The graph of $S_{\mathbb{C}}$ is an algebraic variety with vanishing ideal
 $I_{d,m,n}:=\langle(ACA^{\top})_{i,j}-s_{i,j}\mid 1\leq i,j\leq d\rangle.$
+Note that $I_{d,m,n}$ is an ideal in the polynomial ring of $mnd+d^{2}$ variables: the entries of $A$ and $s_{i,j}$ for $1\leq i,j\leq d$.
 The projection to the target of (19) is then cut out by the prime ideal
 $P_{d,m,n}:=I_{d,m,n}\cap\mathbb{C}[s_{i,j}\mid 1\leq i,j\leq d],$
-and agrees with $\mathcal{P}_{d,m,n}$ by construction.
+and agrees with $\mathcal{M}_{d,m,n}$ by construction.
 Compare Example 4.2 for an step-by-step construction with explicit choices of $d,m$ and $n$.
@@
 Example 4.2.
-For explicit $d,m$ and $n$ we compute $P_{d,m,n}$ with the moment core matrix $C:=\sigma^{(2)}(\mathsf{Mom}^{m})\otimes\sigma^{(2)}(\mathsf{Mom}^{n})$ and with Gröbner bases, applying an elimination of variables procedure.
-We refer to [12] for an introduction into the latter.
-If $m=d=2$ and $n=1$, then $ACA^{\top}$ is
+For explicit $d,m$ and $n$ we compute $P_{d,m,n}$ with the moment core matrix $C_{m,n}=S(\mathsf{Mom}^{m})\otimes S(\mathsf{Mom}^{n})$ and with Gröbner bases, applying an elimination of variables procedure.
+We refer to [CLOS07] for an introduction to the latter.
+  1. 1.
+    If $m=d=2$ and $n=1$, then $AC_{m,n}A^{\top}$ is
     $\begin{bmatrix}\frac{1}{4}a_{1,1}^{2}+\frac{1}{2}a_{1,1}a_{1,2}+\frac{1}{4}a_{1,2}^{2}&\!\!\!\!\frac{1}{4}a_{1,1}a_{2,1}+\frac{1}{3}a_{1,1}a_{2,2}+\frac{1}{6}a_{2,1}a_{1,2}+\frac{1}{4}a_{1,2}a_{2,2}\\ \frac{1}{4}a_{1,1}a_{2,1}+\frac{1}{6}a_{1,1}a_{2,2}+\frac{1}{3}a_{2,1}a_{1,2}+\frac{1}{4}a_{1,2}a_{2,2}&\frac{1}{4}a_{2,1}^{2}+\frac{1}{2}a_{2,1}a_{2,2}+\frac{1}{4}a_{2,2}^{2}\end{bmatrix}$
@@
     $P_{2,2,1}=\langle 4s_{1,1}s_{2,2}-s_{2,1}^{2}-2s_{2,1}s_{1,2}-s_{1,2}^{2}\rangle$ (20)
-is generated by the relation [2, eq.
-(5)] in $\mathbb{C}^{2\times 2}$ (the determinant of the symmetric part of the $2\times 2$-matrix $S$), according to Corollary 3.5.
-If, on the other hand $m=d=n=2$, then
-$I_{2,2,1}=\langle S^{(2)}_{1,1}-s_{1,1},S^{(2)}_{2,1}-s_{2,1},S^{(2)}_{1,2}-s_{1,2},S^{(2)}_{2,2}-s_{2,2}\rangle$ (23)
-with polynomials $S^{(2)}_{i,j}$ according to Example 3.7. With Gröbner bases we can show that $P_{d,m,n}$ is the zero ideal, thus $\dim(\mathcal{P}_{2,2,2})=4$.
-Even for $m=n=2$, the last example illustrates that membranes generate larger varieties than in the path framework [2].
-In fact, the dimension of $\mathcal{P}_{2,2,2}$ is maximal, providing an initial illustration of our main result, Theorem 4.7.
-We now proceed with some technical preparations.
+    is generated by the relation [bib:AFS2019, eq. (5)] in $\mathbb{C}^{2\times 2}$, i.e., the determinant of the symmetric part of the $2\times 2$ matrix $S$.
+  2. 2.
+    If $m=d=n=2$, then
+    $I_{2,2,1}=\langle S_{1,1}-s_{1,1},S_{2,1}-s_{2,1},S_{1,2}-s_{1,2},S_{2,2}-s_{2,2}\rangle$ (21)
+    with polynomials $S_{i,j}=S(X)_{i,j}$ according to Example 3.7. With Gröbner bases we can show that $P_{d,m,n}$ is the zero ideal, thus $\dim(\mathcal{M}_{2,2,2})=4$.
+Even for $m=n=2$, the last example illustrates that membranes generate larger varieties than in the path framework [bib:AFS2019].
+In fact, the dimension of $\mathcal{M}_{2,2,2}$ is maximal, providing an initial illustration of our main result, Theorem 4.13.
+Let us make some easy observations about the varieties $\mathcal{M}_{d,m,n}$.
 Proposition 4.3.
-The variety $\mathcal{P}_{d,m,n}$ is irreducible.
+The variety $\mathcal{M}_{d,m,n}$ is irreducible.
 Proof.
-By Lemma 2.6, $\mathcal{L}_{d,m,n}$ is the schematic image of $A\mapsto A^{\otimes 2}\sigma^{({2})}(\mathsf{Axis}^{m,n})$ where
-$A\in\mathsf{Lin}(\mathbb{C}^{m}\otimes\mathbb{C}^{n},\mathbb{C}^{d})\cong\mathbb{C}^{mnd}$
-is the variety of linear maps $\mathbb{C}^{m}\otimes\mathbb{C}^{n}\to\mathbb{C}^{d}$.
-Thus, as $\mathsf{Lin}(\mathbb{C}^{m}\otimes\mathbb{C}^{n},\mathbb{C}^{d})$ is irreducible, so is $\mathcal{P}_{d,m,n}$.
+By Lemma 2.6, $\mathcal{M}_{d,m,n}$ is the schematic image of $A\mapsto AS(\mathsf{Axis}^{m,n})A^{\top}$ where $A\in\mathbb{C}^{mn\times d}$.
+Thus, as $\mathbb{C}^{mn\times d}$ is irreducible, so is $\mathcal{M}_{d,m,n}$.
 ∎
-Analogously to the stabilizing chain of varieties from [2, Theorem 5.7], we obtain a stabilizing grid:
 Corollary 4.4.
-For every $d\in\mathbb{N}$ there is a grid (8) where we can choose $M,N\in\mathbb{N}$ such that all row and column chains stabilize at (or before) $M$ resp. $N$, that is, for all $m,n\in\mathbb{N}$:
-[…]
-$\mathcal{P}_{d,1,n}\subseteq\dots\subseteq\mathcal{P}_{d,M-1,n}\subseteq\mathcal{P}_{d,M,n}=\mathcal{P}_{d,M+1,n}$ $=\dots$
-Here $M$ and $N$ are constants and depend only on $k$ and $d$. We denote the limit in (8) by $\mathcal{U}_{d}:=\mathcal{P}_{d,M,N}\subseteq\mathbb{C}^{d^{2}}$.
-Proof.
-Follows with Proposition 4.3 and $\dim(\mathcal{P}_{d,m,n})\leq d^{2}$. ∎
-Note that $\mathcal{P}_{d,m,n}=\mathcal{P}_{d,n,m}$ for all $m$ and $n$, since $\mathsf{Axis}^{m,n}$ is transformed into $\mathsf{Axis}^{n,m}$ via the natural isomorphism $\mathbb{R}^{m}\otimes\mathbb{R}^{n}\cong\mathbb{R}^{n}\otimes\mathbb{R}^{m}$.
+For every $d\in\mathbb{N}$ there exist $M,N\in\mathbb{N}$ such that $\mathcal{M}_{d,m,N+i}=\mathcal{M}_{d,m,N}$ and $\mathcal{M}_{d,M+i,n}=\mathcal{M}_{d,M,n}$ for all $i\in\mathbb{N}$.
+In other words, the chains in (6) stabilize.
 Taking a product of an $m$-dimensional path with a $1$-dimensional path results in (scaled) path signature tensors. This immediately implies the following:
 Lemma 4.5.
-$\mathcal{P}_{d,m,1}=\mathcal{P}_{d,m}^{\mathsf{path}}$ where $\mathcal{P}^{\mathsf{path}}_{d,m}$ denotes (the affine cone of) the signature matrix variety considered in [2, Section 3.1].
+$\mathcal{M}_{d,m,1}=\mathcal{M}_{d,m}^{\mathsf{path}}$ where $\mathcal{M}^{\mathsf{path}}_{d,m}$ denotes (the affine cone of) the signature matrix variety considered in [bib:AFS2019, Section 3.1].
 Proof.
@@
 ∎
+The next lemma is technical, but holds the key to understanding the varieties $\mathcal{M}_{d,m,n}$.
+Lemma 4.6.
+Let $C_{m,n}$ denote the signature matrix $S(\mathsf{Axis}_{m,n})$. Then the canonical form for congruence of $C_{m,n}$, as defined in [HORN20061010], is
+  • $\Gamma_{3}\oplus H_{4}(1)^{\oplus\frac{m+n-4}{2}}\oplus\Gamma_{1}^{\oplus(m-2)(n-2)+1}$ if $m,n$ are even,
+  • $\Gamma_{2}\oplus H_{4}(1)^{\oplus\frac{n-1}{2}}\oplus H_{2}(-1)^{\oplus\frac{m-2}{2}}\oplus\Gamma_{1}^{\oplus(m-2)(n-1)}$ if $m$ is even, $n$ is odd,
+  • $H_{2}(-1)^{\oplus\frac{m+n-2}{2}}\oplus\Gamma_{1}^{\oplus(m-1)(n-1)+1}$ if $m,n$ are odd.
+We refer to [TERAN201144] for the definition of the matrices $J_{k}(0)$, $\Gamma_{k}$ and $H_{2k}(\mu)$.
+Proof.
+We use the following fact which is proven in [HORN20061010, page 1016]: let $A$ be a non-singular complex square matrix.
+Then the blocks in the canonical form for congruence of $A$ are in one-to-one correspondence with the blocks in the Jordan canonical form of the cosquare $A^{-\top}A$ in the following way: the blocks $\Gamma_{k}$ in the canonical form for congruence correspond to the blocks $J_{k}((-1)^{k+1})$ in the Jordan canonical form of $A^{-\top}A$ and the blocks $H_{2k}(\mu)$ correspond to pairs of blocks $J_{k}(\mu)\oplus J_{k}(\mu^{-1})$ (the blocks $J_{k}(0)$ do not appear in the non-singular case).
+As $C_{m,n}$ is indeed non-singular (by Example 3.16 we have $\det(C_{m,n})=\frac{1}{4^{mn}}$), our strategy is thus to determine the Jordan canonical form of $C_{m,n}^{-\top}C_{m,n}$.
+This is simplified significantly by the following observation: writing $C_{u}:=S(\mathsf{Axis}^{u})$, we have $C_{m,n}=C_{m}\otimes C_{n}$ by Proposition 3.3, where $\otimes$ denotes the Kronecker product.
+In particular
+$C_{m,n}^{-\top}C_{m,n}=(C_{m}\otimes C_{n})^{-\top}(C_{m}\otimes C_{n})=C_{m}^{-\top}C_{m}\otimes C_{n}^{-\top}C_{n}.$
+Moreover, as observed in the proof of [bib:AFS2019, Theorem 3.4], the matrix $C_{u}$ is congruent to $N_{u}:=\begin{pmatrix}1&1\\ -1&0\end{pmatrix}\oplus\tau^{\frac{u-2}{2}}$ if $u$ is even, and congruent to $N_{u}:=\begin{pmatrix}1\end{pmatrix}\oplus\tau^{\frac{u-1}{2}}$ if $u$ is odd, where $\tau=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$.
+Thus, we may replace $C_{m,n}$ by $N_{m}\otimes N_{n}$, $C_{m}$ by $N_{m}$ and $C_{n}$ by $N_{n}$.
+These observations reduce the problem to the following two questions:
+  1. 1.
+    What are the Jordan canonical forms of $N_{u}^{-\top}N_{u}$?
+  2. 2.
+    Given the Jordan canonical forms of two matrices, what is the Jordan canonical form of their Kronecker product?
+For the first question, note that $\tau^{-\top}\tau=\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix}=J_{1}(-1)^{\oplus 2}$ and
+$\begin{pmatrix}1&1\\ -1&0\end{pmatrix}^{-\top}\begin{pmatrix}1&1\\ -1&0\end{pmatrix}=\begin{pmatrix}-1&0\\ -2&-1\end{pmatrix}$
+whose Jordan canonical form is $\begin{pmatrix}-1&1\\ 0&-1\end{pmatrix}=J_{2}(-1)$.
+Thus, if $u$ is odd then the canonical form of $N_{u}^{-\top}N_{u}$ is $J_{1}(1)\oplus J_{1}(-1)^{\oplus u-1}$, and if $u$ is even then the canonical form of $N_{u}^{-\top}N_{u}$ is $J_{2}(-1)\oplus J_{1}(-1)^{\oplus u-2}$.
+The answer to the second question is [BRUALDI198531, Theorem 4.6] from which we obtain: the Jordan canonical form of $C_{m,n}^{-T}C_{m,n}$ is
+  • $J_{3}(1)\oplus J_{1}(1)\oplus J_{2}(1)^{\oplus m-2+n-2}\oplus J_{1}(1)^{\oplus(m-2)(n-2)}$ if $m,n$ are even,
+  • $J_{2}(-1)\oplus J_{2}(1)^{\oplus n-1}\oplus J_{1}(-1)^{\oplus m-2}\oplus J_{1}(1)^{\oplus(m-2)(n-1)}$ if $m$ is even, $n$ is odd,
+  • $J_{1}(1)\oplus J_{1}(-1)^{\oplus m-1+n-1}\oplus J_{1}(1)^{\oplus(m-1)(n-1)}$ if $m,n$ are odd.
+Thus the canonical form for congruence of $C_{m,n}$ is
+  • $\Gamma_{3}\oplus H_{4}(1)^{\oplus\frac{m+n-4}{2}}\oplus\Gamma_{1}^{\oplus(m-2)(n-2)+1}$ if $m,n$ are even,
+  • $\Gamma_{2}\oplus H_{4}(1)^{\oplus\frac{n-1}{2}}\oplus H_{2}(-1)^{\oplus\frac{m-2}{2}}\oplus\Gamma_{1}^{\oplus(m-2)(n-1)}$ if $m$ is even, $n$ is odd,
+  • $H_{2}(-1)^{\oplus\frac{m+n-2}{2}}\oplus\Gamma_{1}^{\oplus(m-1)(n-1)+1}$ if $m,n$ are odd.
 Corollary 4.7.
-For each $d$ there exists an $M\in\mathbb{N}$ such that
-$\mathcal{P}_{d,m,1}=\mathcal{P}_{d,1,m}=\mathcal{U}^{\mathsf{path}}_{d,2}$
-for all $m\geq M$, and where $\mathcal{U}^{\mathsf{path}}_{d,2}$ is the universal path variety at level $2$.
-Proof.
-This is now immediate from Lemma 4.5 and [2, Theorem 5.6]. ∎
-This gives a lower bound on the dimension of $\mathcal{U}_{d}$ because of the grid structure (8); in the next section we show that $\mathcal{U}_{d}$ is actually the whole space.
-4.1 The case $m,n\gg d$
-Theorem 4.7.
-The limit variety $\mathcal{U}_{d}$ attains the full ambient dimension, that is, $\mathcal{U}_{d}=\mathbb{C}^{d^{2}}$.
-Proof.
-We claim that already $\mathcal{P}_{d,d,2}=\mathbb{C}^{d^{2}}$.
-For this, write $A=\sigma^{(2)}(\mathsf{Axis}^{d})$ and note that by Lemma 2.6, Proposition 3.3 and Example 3.17, $\mathcal{P}_{d,d,2}$ is the Zariski closure of the image of the map $\phi:\mathbb{C}^{d\times d}\times\mathbb{C}^{d\times d}\to\mathbb{C}^{d\times d}$ sending $(S,T)$ to
-$\begin{pmatrix}S^{\top}&T^{\top}\end{pmatrix}\begin{pmatrix}A&A\\ -A&0\end{pmatrix}\begin{pmatrix}S\\ T\end{pmatrix}=S^{\top}AS+S^{\top}AT-T^{\top}A.$
-Here we used that $\begin{pmatrix}1&1\\ -1&0\end{pmatrix}$ is congruent to $\sigma^{(2)}(\mathsf{Axis}^{2})$. The Jacobian of $\phi$ at $\begin{pmatrix}I_{d}\\ I_{d}\end{pmatrix}$ is the linear map
-$\mathbb{C}^{d\times d}\times\mathbb{C}^{d\times d}$ $\to\mathbb{C}^{d\times d},$
-$(S,T)$ $\mapsto 2S^{\top}A+AT-T^{\top}A.$
-As $A$ is invertible ($\det(A)=\frac{1}{2^{d}}$ by (11)), we see that the kernel of this map has dimension $d^{2}$ since, if $(S,T)$ lies in the kernel, we can solve for $S$.
-Thus the Jacobian has rank $d^{2}$ at this point, proving that $\phi$ is smooth in a neighbourhood $U$ of $\begin{pmatrix}I_{d}\\ I_{d}\end{pmatrix}$, [27, Definition 6.14, Proposition 6.15], and thus $\phi|_{U}$ is open by [43, Tag 056G].
-This proves the claim since $\mathbb{C}^{d\times d}$ is irreducible.
-∎
-Remark 4.8.
-The same proof idea yields a more general bound: for $e\leq d$, we have $\dim(\mathcal{P}_{d,e,2})\geq 2de-e^{2}$.
-To see this, one only needs to adapt the matrix formats in the proof above and replace $\begin{pmatrix}I_{d}\\ I_{d}\end{pmatrix}$ by the matrix $\begin{pmatrix}X\\ X\end{pmatrix}$ where $X=\begin{pmatrix}I_{e}&0\end{pmatrix}$.
-Unfortunately, this bound is not tight in any of our examples, cf.
-Figure 5 below.
-4.2 The case $mn\leq d$
-Theorem 4.9.
-For $mn\leq d$ the dimension of $\mathcal{P}_{d,m,n}$ is
+For $mn\leq d$ the dimension of $\mathcal{M}_{d,m,n}$ is
   • $dmn-\frac{1}{2}m^{2}n^{2}+m^{2}(n-1)+(m-1)n^{2}-\frac{7}{2}mn+4(m+n)-4$ for even $m,n$.
@@
 Proof.
-The variety $\mathcal{P}_{d,m,n}$ is the Zariski closure of the image of the map $\phi:\mathbb{C}^{d\times mn}\to\mathbb{C}^{d\times d},\ S\mapsto SA_{m,n}S^{\top}$ where $A_{m,n}=\sigma^{(2)}(\text{Axis}^{m,n})$.
-Equivalently, it is the Zariski closure of the $\operatorname{GL}_{d}$-congruence orbit of the $d\times d$-matrix $A^{\prime}_{m,n}$ which is obtained from $A_{m,n}$ by extending with zeroes.
-This congruence orbit is a differentiable manifold and a semi-algebraic set, in particular the dimension of the manifold agrees with the dimension of its Zariski-closure.
-The dimension of the manifold agrees with the dimension of the tangent space at $\phi(I_{d})$ (as it is a group orbit).
-Its codimension is the dimension of the kernel of the Jacobian $J(\phi)$ at $I_{d}$, in other words, the dimension of the solution space to the matrix equation
-$SA^{\prime}_{m,n}+A^{\prime}_{m,n}S^{\top}=0,$ (24)
-cf. [45, Lemma 2]. This dimension remains unchanged when $A^{\prime}_{m,n}$ is replaced by $A_{m,n}$. Thus in the following, we may assume $d=mn$.
-[…]
-As $A_{m,n}$ is indeed non-singular (by Example 3.17 we have $\det(A_{m,n})=\frac{1}{4^{mn}}$), our strategy is thus to determine the Jordan canonical form of $A_{m,n}^{-\top}A_{m,n}$.
-This is simplified significantly by the following observation: if we denote $A_{u}:=\sigma^{(2)}(\mathsf{Axis}^{u})$ then by Proposition 3.3 we have $A_{m,n}=A_{m}\otimes A_{n}$ where $\otimes$ denotes the Kronecker product.
-In particular
-$A_{m,n}^{-\top}A_{m,n}=(A_{m}\otimes A_{n})^{-\top}(A_{m}\otimes A_{n})=A_{m}^{-\top}A_{m}\otimes A_{n}^{-\top}A_{n}.$
-Moreover, as observed in the proof of [2, Theorem 3.4], the matrix $A_{u}$ is congruent to $N_{u}:=\begin{pmatrix}1&1\\ -1&0\end{pmatrix}\oplus\tau^{\frac{u-2}{2}}$ if $u$ is even, and congruent to $N_{u}:=\begin{pmatrix}1\end{pmatrix}\oplus\tau^{\frac{u-1}{2}}$ if $u$ is odd, where $\tau=\begin{pmatrix}1&-1\\ -1&0\end{pmatrix}$.
-Thus, we may replace $A_{m,n}$ by $N_{m}\otimes N_{n}$, $A_{m}$ by $N_{m}$ and $A_{n}$ by $N_{n}$.
-These observations reduce the problem to the following two questions:
+This follows from Lemma 4.6 by using the codimension count in [TERAN201144, Theorem 2]. ∎
+Remark 4.8.
+In light of Lemma 4.5, one checks that for $n=1$ the polynomials in Corollary 4.7 simplify to $md-\binom{m}{2}$, which is indeed the dimension of the affine cone of the path signature matrix varieties considered in [bib:AFS2019, Theorem 3.4].
+Corollary 4.9.
+Let $\mathcal{S}_{a,b}$ be the variety of $d\times d$ matrices whose symmetric part has rank $\leq a$ and whose skew-symmetric part has rank $\leq b$. Then
   1. 1.
-    What are the Jordan canonical forms of $N_{u}^{-\top}N_{u}$?
+    $\mathcal{S}_{(m-2)(n-2),m+n-2}\subseteq\mathcal{M}_{d,m,n}\subseteq\mathcal{S}_{mn,m+n-2}$ if $m,n$ are even,
   2. 2.
-    Given the Jordan canonical forms of two matrices, what is the Jordan canonical form of their Kronecker product?
-For the first question, note that $\tau^{-\top}\tau=\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix}=J_{1}(-1)$ and
-$\begin{pmatrix}1&1\\ -1&0\end{pmatrix}^{-\top}\begin{pmatrix}1&1\\ -1&0\end{pmatrix}=\begin{pmatrix}-1&0\\ -2&-1\end{pmatrix}$
-whose Jordan canonical form is $\begin{pmatrix}-1&1\\ 0&-1\end{pmatrix}=J_{2}(-1)$.
-Thus, if $u$ is odd then the canonical form of $N_{u}^{-\top}N_{u}$ is $J_{1}(1)\oplus J_{1}(-1)^{\oplus u-1}$, and if $u$ is even then the canonical form of $N_{u}^{-\top}N_{u}$ is $J_{2}(-1)\oplus J_{1}(-1)^{\oplus u-2}$.
-The answer to the second question is [7, Theorem 4.6] from which we obtain: the Jordan canonical form of $A_{m,n}^{-T}A_{m,n}$ is
-  • $J_{3}(1)\oplus J_{1}(1)\oplus J_{2}(1)^{\oplus m-2+n-2}\oplus J_{1}(1)^{\oplus(m-2)(n-2)}$ if $m,n$ are even,
-  • $J_{2}(-1)\oplus J_{2}(1)^{\oplus n-1}\oplus J_{1}(-1)^{\oplus m-2}\oplus J_{1}(1)^{\oplus(m-2)(n-1)}$ if $m$ is even, $n$ is odd,
-  • $J_{1}(1)\oplus J_{1}(-1)^{\oplus m-1+n-1}\oplus J_{1}(1)^{\oplus(m-1)(n-1)}$ if $m,n$ are odd.
-Thus the canonical form for congruence of $A_{m,n}$ is
-  • $\Gamma_{3}\oplus H_{4}(1)^{\oplus\frac{m+n-4}{2}}\oplus\Gamma_{1}^{\oplus(m-2)(n-2)+1}$ if $m,n$ are even,
-  • $\Gamma_{2}\oplus H_{4}(1)^{\oplus\frac{n-1}{2}}\oplus H_{2}(-1)^{\oplus\frac{m-2}{2}}\oplus\Gamma_{1}^{\oplus(m-2)(n-1)}$ if $m$ is even, $n$ is odd,
-  • $H_{2}(-1)^{\oplus\frac{m+n-2}{2}}\oplus\Gamma_{1}^{\oplus(m-1)(n-1)+1}$ if $m,n$ are odd.
-Finally, we conclude by calculating the codimensions using the recipe given in [45, Theorem 2]. ∎
-$\left[\begin{array}[]{cccc}4&7&9&10\\ 7&14&\mathbf{16}&16\\ 9&\mathbf{16}&\mathbf{16}&16\\ 10&16&16&16\end{array}\right]$
+    $\mathcal{S}_{2,n-1}\subseteq\mathcal{M}_{d,2,n}\subseteq\mathcal{S}_{2(n-1)+1,n+1}$ if $n$ is odd,
+  3. 3.
+    $\mathcal{S}_{(m-2)(n-1)-1,m+n-1}\subseteq\mathcal{M}_{d,m,n}\subseteq\mathcal{S}_{m(n-1)+1,m+n-1}$ if $m\geq 4$ is even, $n$ is odd,
+  4. 4.
+    $\mathcal{S}_{(m-1)(n-1)+1,m+n-2}=\mathcal{M}_{d,m,n}$ if $m,n$ are odd.
+Proof.
+For the upper bounds, we just need to compute the ranks of the symmetric and skew-symmetric parts of the normal forms in Lemma 4.6.
+  • For $m,n$ even the rank of $C_{m,n}^{\mathsf{sym}}$ is $(m-2)(n-2)+1+2m+2n-8+3=mn$. The rank of $C_{m,n}^{\mathsf{sk}}$ is $22+m+n-4=m+n-$ 2.
+  • For $m$ even, $n$ odd the rank of $C_{m,n}^{\mathsf{sym}}$ is $11+2(n-1)+(m-2)(n-1)=m(n-1)+$ 1. The rank of $C_{m,n}^{\mathsf{sk}}$ is $12+n-1+m-2=m+n-$ 1.
+  • For $m,n$ odd the rank of $C_{m,n}^{\mathsf{sym}}$ is $(m-1)(n-1)+1$. The rank of $C_{m,n}^{\mathsf{sk}}$ is $m+n-2$.
+For the lower bound, note that for a general matrix $X\in\mathcal{S}_{a,b}$ there are $S,T\in\operatorname{GL}_{d}(\mathbb{C})$ with $S^{\top}X^{\mathsf{sym}}S=\Gamma_{1}^{\oplus a^{\prime}}$ and $T^{\top}X^{\mathsf{sk}}T=H_{2}(-1)^{\oplus b^{\prime}}$ where $a^{\prime}\leq a$ and $b^{\prime}\leq\frac{b}{2}$.
+Thus, to give a lower bound, one can count blocks in the normal form of $C_{m,n}$ that are congruent to $H_{2}(-1)$ and can thus contribute to the skew-symmetric rank, such that all other blocks are congruent to $\Gamma_{1}$ and can contribute to the symmetric rank.
+For $m,n$ even, we have that $\Gamma_{3}\oplus\Gamma_{1}$ is congruent to $H_{2}(-1)$, as is $H_{4}(1)$.
+Thus, every general skew-symmetric matrix of rank $42+m+n-$ 4 is congruent to $\Gamma_{3}\oplus H_{4}(1)^{\oplus\frac{m+n-4}{2}}\oplus\Gamma_{1}$.
+As every rank $(m-2)(n-2)$ symmetric matrix is congruent to $\Gamma_{1}^{\oplus(m-2)(n-2)}$ we conclude.
+For $m=2$, $n$ odd, we have that $\Gamma_{2}\oplus H_{4}(1)$ is congruent to $H_{2}(-1)\oplus\Gamma_{1}^{\oplus 2}$.
+Every general skew-symmetric matrix of rank $12+n-3=n-$ 1 is congruent to $H_{2}(-1)\oplus H_{4}(1)^{\oplus\frac{n-1}{2}}$ and every symmetric matrix of rank $2$ is congruent to $\Gamma_{1}^{\oplus 2}$ and we conclude.
+For $m$ even, $n$ odd, and $m\geq 4$ we know that $\Gamma_{2}\oplus\Gamma_{1}$ is congruent to $H_{2}(-1)$; thus every general skew-symmetric matrix of rank $12+n-1+m-2=m+n-$ 1 is congruent to $\Gamma_{2}\oplus H_{4}(1)^{\oplus\frac{n-1}{2}}\oplus H_{2}(-1)^{\oplus\frac{m-2}{2}}\oplus\Gamma_{1}$; and we conclude again.
+By the same reasoning, we immediately see that for $m,n$ odd we just obtain the variety $\mathcal{S}_{(m-1)(n-1)+1,m+n-2}$ as a lower bound as well and deduce equality with $\mathcal{M}_{d,m,n}$. ∎
+Note that for $m=1$ or $n=1$ we obtain a rank one symmetric part in accordance with the path framework, [bib:AFS2019, eq.
+(4)].
+Moreover, the corollary shows for all cases except $m=2$, $n$ odd, that the projection of matrices in $\mathcal{M}_{d,m,n}$ to the skew-symmetric part gives precisely the variety of skew-symmetric matrices of rank $\leq m+n-1$ (if $m+n$ is odd) or $\leq m+n-2$ (if $m+n$ is even).
+In the case where $m,n$ are odd, we can describe the variety precisely:
+Corollary 4.10.
+The degree of $\mathcal{M}_{d,m,n}$ is
+$\frac{1}{2^{d-(m+n)+2}}\prod_{\alpha=0}^{d-(m-1)(n-1)}\frac{\binom{d+\alpha}{d-(m-1)(n-1)+1-\alpha}}{\binom{2\alpha+1}{\alpha}}\prod_{\alpha=0}^{d-(m+n)}\frac{\binom{d+\alpha}{d-(m+n)+1-\alpha}}{\binom{2\alpha+1}{\alpha}}.$
+Proof.
+This follows from Corollary 4.9 since $\mathcal{M}_{d,m,n}=\mathcal{S}_{(m-1)(n-1)+1,m+n-2}$.
+Indeed, for $m,n$ odd, $m+n\leq d$, $\mathcal{S}_{(m-1)(n-1)+1,m+n-2}$ is the transverse intersection of the vanishing locus of the $((m-1)(n-1)+2)$-minors of the symmetric part with the vanishing locus of the $(m+n)$-Pfaffians of the skew-symmetric part.
+The degree of these two vanishing loci is calculated in [HARRIS198471, Proposition 12].
+∎
+Example 4.11.
+According to Corollary 4.7 and Corollary 4.10 the variety $\mathcal{M}_{6,3,3}$ has dimension $34$ and degree $18$.
+It is a complete intersection of the vanishing locus of the Pfaffian of the skew-symmetric part and the vanishing locus of the determinant of the symmetric part.
+Example 4.12.
+Consider the case $d=4$ and $m=n=2$.
+By Corollary 4.9, we have $\mathcal{S}_{0,2}\subseteq\mathcal{M}_{d,m,n}\subseteq\mathcal{S}_{4,2}$.
+In particular, the Pfaffian $x_{12}x_{34}-x_{13}x_{24}+x_{14}x_{23}$ vanishes on the skew-symmetric part of matrices in the congruence orbit.
+However, according to Corollary 4.7 the dimension of $\mathcal{M}_{4,2,2}$ is 14, so there must be additional relations on the symmetric part.
+Indeed, another relation is given by $\det(A)\det(X^{\mathsf{sym}})=\det(X)\det(A^{\mathsf{sym}})$: the ratio $\det(X)/\det(X^{\mathsf{sym}})$ is well-defined and clearly constant on the congruence orbit.
+Using Macaulay 2 one checks that the variety $\mathcal{M}_{4,2,2}$ is a complete intersection, cut out by those two equations.
+It is immediate from Corollary 4.9 that $\mathcal{M}_{d,m,n}$ attains the full ambient dimension $d^{2}$ when $m,n\to\infty$. Our main theorem gives a precise bound:
+Theorem 4.13.
+For $m,n\geq 2$, we have $\mathcal{M}_{d,m,n}=\mathbb{C}^{d\times d}$ whenever $m+n\geq d+1$.
+Proof.
+Clearly, it suffices to prove the statement for $d=m+n-1$.
+We make a case distinction for the different parities of $m$ and $n$.
+If $m$ and $n$ are odd, then we have $(m-1)(n-1)+1\geq m+n-1$.
+Moreover, $d=m+n-1$ is odd, so that every $d\times d$ matrix has skew-symmetric rank $\leq m+n-2$.
+Thus, we immediately conclude by Corollary 4.9.
+We prove the statement for the remaining cases by induction on $m+n$. We settle the base cases by explicit linear algebra calculations.
+The base case. Consider the map $\phi:\mathbb{C}^{mn\times d}\to\mathbb{C}^{d\times d}$ given by $A\mapsto A^{\top}C_{m,n}A$. Let us write $\phi_{B}^{\prime}$ for its derivative
+$\mathbb{C}^{mn\times d}\to\mathbb{C}^{d\times d},A\mapsto B^{\top}C_{m,n}A+A^{\top}C_{m,n}B$
+at a matrix $B\in\mathbb{C}^{mn\times d}$. It suffices to give a matrix $B$ where $\phi_{B}^{\prime}$ has full rank.
+The base case for $m,n$ even is $m=n=2$, $d=3$.
+Then the canonical form of $A_{2,2}$ is $\Gamma_{3}\oplus J_{1}(1)$.
+We compute that for $B=\begin{pmatrix}\mathrm{I}_{d}&\mathbbm{1}\end{pmatrix}$, $\phi_{B}^{\prime}$ has full rank.
+The base case for $m$ even, $n$ odd is $m=2$, $n=3$, $d=4$. We compute that for $B=\begin{pmatrix}\mathrm{I}_{d}&\mathbbm{1}&\mathbbm{1}\end{pmatrix}$, $\phi_{B}^{\prime}$ has full rank.
+The induction step.
+Assume the statement is true for $m+n=d+1$ for some $d$.
+We claim that it then also holds for $m+n=d+3=(d+2)+1$, i.e.
+$\mathcal{M}_{d+2,m,n}=\mathbb{C}^{d+2,d+2}$.
+By assumption, we have $\mathcal{M}_{d,m,n-2}=\mathbb{C}^{d\times d}$ or $\mathcal{M}_{d,m-2,n}=\mathbb{C}^{d\times d}$.
+A general $\mathbb{(}d+2)\times(d+2)$-matrix has normal form $\oplus_{i\leq k}H_{2}(\mu_{i})\oplus\Gamma_{1}^{\oplus l}$ where $22k+l=d+$ 2 with $l\in\{0,1\}$ and $\mu_{i}\in\mathbb{C}$.
+In particular, wlog.
+we can replace a general $(d+2)\times(d+2)$ matrix by a matrix $D\oplus E$ for $D$ a general $d\times d$-matrix and $E$ a general $2\times 2$-matrix.
+If $m$ and $n$ are even and wlog.
+$m\geq 4$, then we have $C_{m,n}=C_{m-2,n}\oplus H_{4}(1)\oplus\Gamma_{1}^{\oplus 2(n-2)+1}$.
+A calculation in Macaulay 2 shows that generically the fibers of the map $\mathbb{C}^{4\times 2}\to\mathbb{C}^{2\times 2},A\mapsto A^{\top}H_{4}(1)A$ are $5$-dimensional.
+If $m$ is even and $n$ is odd, then if $m\geq 4$ we have $C_{m,n}=C_{m-2,n}\oplus H_{2}(-1)\oplus\Gamma_{1}^{\oplus 2(n-1)}$.
+Again, a calculation in Macaulay 2 shows that generically the fibers of the map $\mathbb{C}^{4\times 2}\to\mathbb{C}^{2\times 2},A\mapsto A^{\top}(H_{2}(-1)\oplus\Gamma_{1}^{\oplus 2})A$ are $5$-dimensional.
+If $n\geq 5$, then $C_{m,n}=A_{m,n-2}\oplus H_{4}(1)\oplus\Gamma_{1}^{\oplus 2(m-2)}$ and we conclude as in the case $m,n$ even.
+Finally, note that for matrices $A,B,X,Y$ we have $(A\oplus B)^{\top}(X\oplus Y)(A\oplus B)=A^{\top}XA\oplus B^{\top}YB$.
+Thus by the induction hypothesis we conclude in all cases that $\mathcal{M}_{d+2,m,n}=\mathbb{C}^{d+2,d+2}$.
+∎
+For $m=1$ or $n=1$ we obtain a rank one symmetric part according to the path framework, [bib:AFS2019, eq. (4)].
+$\left[\begin{array}[]{cccc}4&7&9&10\\ 7&14&16&\\ 9&16&&\\ 10&&&\end{array}\right]$
 $d=4$
-$\left[\small\begin{array}[]{ccccc}5&9&12&14&15\\ 9&18&\mathbf{24}&\mathbf{25}&25\\ 12&\mathbf{24}&\mathbf{25}&\mathbf{25}&25\\ 14&\mathbf{25}&\mathbf{25}&\mathbf{25}&25\\ 15&25&25&25&25\end{array}\right]$
+$\left[\small\begin{array}[]{ccccc}5&9&12&14&15\\ 9&18&\mathbf{24}&25&\\ 12&\mathbf{24}&25&&\\ 14&25&&&\\ 15&&&&\end{array}\right]$
 $d=5$
-$\left[\small\begin{array}[]{cccccc}6&11&15&18&20&21\\ 11&22&31&\mathbf{34}&\mathbf{36}&36\\ 15&31&\mathbf{34}&\mathbf{36}&\mathbf{36}&36\\ 18&\mathbf{34}&\mathbf{36}&\mathbf{36}&\mathbf{36}&36\\ 20&\mathbf{36}&\mathbf{36}&\mathbf{36}&\mathbf{36}&36\\ 21&36&36&36&36&36\end{array}\right]$
+$\left[\small\begin{array}[]{cccccc}6&11&15&18&20&21\\ 11&22&31&\mathbf{34}&36&\\ 15&31&\mathbf{34}&36&&\\ 18&\mathbf{34}&36&&&\\ 20&36&&&&\\ 21&&&&&\end{array}\right]$
 $d=6$
-$\left[\small\begin{array}[]{ccccccc}7&13&18&22&25&27&28\\ 13&26&37&\mathbf{43}&\mathbf{48}&\mathbf{49}&49\\ 18&37&\mathbf{43}&\mathbf{49}&\mathbf{49}&\mathbf{49}&49\\ 22&\mathbf{43}&\mathbf{49}&\mathbf{49}&\mathbf{49}&\mathbf{49}&49\\ 25&\mathbf{48}&\mathbf{49}&\mathbf{49}&\mathbf{49}&\mathbf{49}&49\\ 27&\mathbf{49}&\mathbf{49}&\mathbf{49}&\mathbf{49}&\mathbf{49}&49\\ 28&49&49&49&49&49&49\end{array}\right]$
+$\left[\small\begin{array}[]{ccccccc}7&13&18&22&25&27&28\\ 13&26&37&\mathbf{43}&\mathbf{48}&49&\\ 18&37&\mathbf{43}&\mathbf{49}&&&\\ 22&\mathbf{43}&\mathbf{49}&&&&\\ 25&\mathbf{48}&&&&&\\ 27&49&&&&&\\ 28&&&&&&\end{array}\right]$
 $d=7$
-$\left[\small\begin{array}[]{cccccccc}8&15&21&26&30&33&35&36\\ 15&30&43&52&\mathbf{60}&\mathbf{62}&\mathbf{64}&64\\ 21&43&\mathbf{52}&\mathbf{63}&\mathbf{63}&\mathbf{64}&\mathbf{64}&64\\ 26&52&\mathbf{63}&\mathbf{63}&\mathbf{64}&\mathbf{64}&\mathbf{64}&64\\ 30&\mathbf{60}&\mathbf{63}&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&64\\ 33&\mathbf{62}&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&64\\ 35&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&64\\ 36&64&64&64&64&64&64&64\\ \end{array}\right]$
+$\left[\small\begin{array}[]{cccccccc}8&15&21&26&30&33&35&36\\ 15&30&43&52&\mathbf{60}&\mathbf{62}&64&\\ 21&43&\mathbf{52}&\mathbf{63}&\mathbf{63}&64&&\\ 26&52&\mathbf{63}&\mathbf{63}&64&&&\\ 30&\mathbf{60}&\mathbf{63}&64&&&&\\ 33&\mathbf{62}&64&&&&&\\ 35&64&&&&&&\\ 36&&&&&&&\\ \end{array}\right]$
 $d=8$
-Figure 5: The dimensions $\dim(\mathcal{P}_{d,m,n})$ stored in $d\times d$ matrices for varying $d$, where bold numbers are not covered by Theorem 4.7 or Theorem 4.9.
+Figure 6: The dimensions $\dim(\mathcal{M}_{d,m,n})$ stored in $d\times d$ matrices for varying $d$, where bold numbers are not covered by Theorem 4.13 or Corollary 4.7.
+We omit the values once the full dimension has been reached.
 We compute these remaining dimensions according to Remark 4.14.
-Remarkable is the fact that in all examples the full dimension $d^{2}$ is already attained on the antidiagonal $m+n=d+1$ for $m,n\geq 2$, in particular $\mathcal{P}_{d,d-1,2}$ has full dimension, which is not clear from Remark 4.8.
 Remark 4.14.
-In light of Lemma 4.5, one checks that for $n=1$ the formulas simplify to $md-\binom{m}{2}$, which is indeed the dimension of the affine cone of the path signature matrix varieties considered in [2], cf.
-Theorem 3.4 in loc.
-cit..
-When rescaling the rows of $\sigma(\mathsf{Mom}^{m})$ we obtain a Cauchy matrix, hence with Corollary 3.6, $\sigma(\mathsf{Mom}^{m,n})$ has full rank for any $m$ and $n$.
-However, the rank of the symmetric and skew-symmetric part behaves differently as in the path case:
-Corollary 4.11.
-For $A_{m,n}:=\sigma^{(2)}(\mathsf{Mom}^{m,n})$ let $A_{m,n}^{\mathsf{sym}}$ and $A_{m,n}^{\mathsf{sk}}$ denote its symmetric and skew-symmetric part, respectively. Then
-$\operatorname{rk}(A_{m,n}^{\mathsf{sym}})=\begin{cases}m\cdot n&m,n\text{ even}\\ 1+m(n-1)&m\text{ even},n\text{ odd}\\ 1+(m-1)(n-1)&m,n\text{ odd}\end{cases}$
-(again we omit the case $m$ odd, $n$ even due to symmetry) and
-$\operatorname{rk}(A_{m,n}^{\mathsf{sk}})=\begin{cases}m+n-1&m+n\text{ odd}\\ m+n-2&m+n\text{ even.}\end{cases}$
-Proof.
-In the proof of Theorem 4.9, add up the ranks of the symmetric/skew-symmetric parts of the blocks in the canonical form for congruence of $A_{m,n}$. ∎
-For $m=1$ or $n=1$ we obtain a rank one symmetric part according to the path framework, [2, eq. (4)].
-As the rank of symmetric and skew-symmetric part is preserved under congruence, Corollary 4.11 induces constraints on the congruence orbits. However not all relations on the orbit arise in this way:
-Example 4.12.
-Consider the case $d=4,m,n=2$.
-According to Corollary 4.11 the symmetric part of $A:=\sigma^{(2)}(\mathsf{Axis}^{m,n})$ has full rank and the skew-symmetric part has rank $2$.
-In particular, the Pfaffian $x_{12}x_{34}-x_{13}x_{24}+x_{14}x_{23}$ vanishes on the skew-symmetric part of matrices in the congruence orbit.
-However, according to Theorem 4.9 the dimension of $\mathcal{P}_{4,2,2}$ is 14, so this can not be the only relation.
-Indeed, another relation is given by $\det(A)\det(X^{\mathsf{sym}})=\det(X)\det(A^{\mathsf{sym}})$: the ratio $\det(X)/\det(X^{\mathsf{sym}})$ is well-defined and clearly constant on the congruence orbit.
-Using Macaulay 2 one checks that the variety $\mathcal{P}_{4,2,2}$ is a complete intersection, cut out by those two equations.
-Remark 4.13.
-Some choices of $m,n,d$ are neither covered by Theorem 4.7 nor by Theorem 4.9.
-Using Macaulay 2 with NumericalImplicitization, we computed the missing dimensions of $\mathcal{P}_{d,m,n}$ for $m,n\leq d$.
+Some choices of $m,n,d$ are neither covered by Theorem 4.13 nor by Corollary 4.7.
+Using Macaulay 2 with NumericalImplicitization [chen2019numerical], we computed the missing dimensions of $\mathcal{M}_{d,m,n}$ for $m,n\leq d$.
 We summarize our findings in Figure 6.
 Example 4.15.
-According to Figure 5, $\mathcal{P}_{5,3,2}$ is a hypersurface in $\mathbb{R}^{5\times 5}$.
-By numerical computations in Macaulay 2 using the packages NumericalImplicitization and MultigradedImplicitization one observes that this hypersurface is cut out by a single quintic $f$ with $999$ terms.
-One checks that $f\in\mathsf{minors}(3,X)\cap\mathsf{minors}(4,X^{\mathsf{sk}})\cap\mathsf{minors}(1,X^{\mathsf{sym}})$, where $\mathsf{minors}(i,X)$ denotes the ideal of $i$-minors of $X$.
-5 Conclusions and future work
-Our notion of signature was introduced by [26, 14] under the name of $\mathsf{id}$-signature, see Remark 2.3 for a detailed discussion.
+Due to Figure 6, $\mathcal{M}_{5,3,2}$ is a hypersurface in $\mathbb{R}^{5\times 5}$.
+Using the packages NumericalImplicitization and MultigradedImplicitization [cummings2025computing], one observes that this hypersurface is cut out by a single quintic $f$ with $999$ terms.
+One checks that $f\in\mathsf{minors}(3,X)\cap\mathsf{minors}(4,X^{\mathsf{sk}})\cap\mathsf{minors}(1,X^{\mathsf{sym}})$, where $\mathsf{minors}(i,X)$ denotes the ideal of $i$-minors of $X$ from the congruence orbit.
+5 Outlook and future work
+Our notion of signature was introduced by [GLNO2022, diehl2024signature] under the name of $\mathsf{id}$-signature, see Remark 2.3 for a detailed discussion.
 Both references observe that this $\mathsf{id}$-signature is not universal.
@@
 Theorem 4.13 holds for arbitrary level $k\geq 3$.
-Furthermore, we propose novel closed formulas for signatures of piecewise bilinear and polynomial membranes.
-These membranes can be used for interpolating arbitrary image data, a necessary step in the iterated-integrals signature method.
-This should be useful for [48] which, instead of any interpolation, approximates two-parameter integrals via certain two-parameter sums.
-Inspired by linear complexity of the two-parameter sums signature on diagonal matrix compositions [18, Theorem 4.5], we would like to improve our computational complexity for interpolated image data, Theorem 3.19.
+We provide computational evidence for $d=k=3$ in Figure 7.
+$m\!\setminus\!n$ 1 2 3 4 5
+1 3 6 9 12 14
+2 6 12 18 24 27
+3 9 18 27
+4 12 24
+5 14 27
+Figure 7: Dimensions of the Zariski closure of third level signature tensors of piecewise bilinear membranes of order $(m,n)$ in $\mathbb{R}^{3}$, where we omit the values once the full dimension has been reached.
+Already in the path framework, a signature matrix is not enough to recover the path, see [bib:AFS2019, Example 6.8].
+In particular, we can not recover a membrane from its signature matrix in general.
+An important property of the ‘right’ notion of signature is the ability of recovering membranes from all of its signature tensors, as it is done in [Chen1954IteratedIA] for paths.
+It is an open question whether our signature tensors have this property.
 Question 5.2.
-Is there an algorithm with linear time complexity in $m$ and $n$ to compute a level $k$ signature entry of a piecewise bilinear membranes of order $(m,n)$?
-Already in the path framework, a signature matrix is not enough to recover the path, see [2, Example 6.8].
-In particular, we can not recover a membrane from its signature matrix in general.
-An important property of the ‘right’ notion of signature is the ability of recovering membranes from all of its signature tensors, as it is done in [8] for paths.
-It is an open question whether our signature tensors have this property.
+Can we always recover membranes from its signature tensors?
+One can also ask the following more special question:
 Question 5.3.
-Can we always recover membranes from their signature tensors?
-One can also ask the following more special question:
-Question 5.4.
 Can piecewise bilinear membranes of low order already be recovered from a finite set of signature tensors?
-Again, this is true for paths, see [41].
-We also remark that recovery with the two-parameter sums signature is always possible. For this compare [18, Theorem 2.20] and its proof.
-Finally, we would like to understand the missing dimensions for membranes of very low order. Computations as shown in Figure 5 inspire the following conjecture:
-Conjecture 5.5.
-For $m,n\geq 2$, $\mathcal{P}_{d,m,n}$ has full dimension whenever $m+n>d$.
-Note that these cases are not covered by Remark 4.8.
-Question 5.6.
-What are the dimensions of $\mathcal{P}_{d,n,m}$ for $m+n\leq d<mn$?
-Acknowledgments
-We thank Joscha Diehl for detailed discussions on the various notions of two-parameter signatures, Sven Groß for introducing us to interpolation literature, Ben Hollering for providing Example 4.14, Rosa Preiß for pointing out [14, Example 4.2], and Bernd Sturmfels for his help with Example 4.12.
-The authors acknowledge seed support for DFG CRC/TRR 388 “Rough Analysis, Stochastic Dynamics and Related Topics”.
+Again, this is true for paths, see [bib:PSS2019].
+We also remark that recovery with the two-parameter sums signature is always possible. For this compare [DS22, Theorem 2.20] and its proof.
+Acknowledgements
+We thank Joscha Diehl for detailed discussions on various notions of two-parameter signatures, Sven Groß for introducing us to the interpolation literature, Rosa Preiß for pointing out [diehl2024signature, Example 4.2], and Ben Hollering and Bernd Sturmfels for their help with Examples 4.15 and 4.12.
+Funding sources
+The authors acknowledge support from DFG CRC/TRR 388 “Rough Analysis, Stochastic Dynamics and Related Fields”, Project A04.
+[…]